### v2

# CMSO-transducing tree-like graph decompositions

Abstract
We show that given a graph $G$ we can $\operatorname{C}{\!\!}\operatorname{MSO}$-transduce its modular decomposition, its split decomposition and its bi-join decomposition. This improves results by Courcelle [Logical Methods in Computer Science, 2006] who gave such transductions using order-invariant $\operatorname{MSO}$, a strictly more expressive logic than $\operatorname{C}{\!\!}\operatorname{MSO}$. Our methods more generally yield $\operatorname{C}_{2}{\!\!}\operatorname{MSO}$-transductions of the canonical decomposition of weakly-partitive set systems and weakly-bipartitive systems of bipartitions.
keywords:
MSO-transduction, MSO-definability, graph decomposisions
category:
[…]
1 Introduction
A decomposition of a graph, especially a tree-like decomposition, is a result of recursive separations of a graph and is extremely useful for investigating combinatorial properties such as colourability, and for algorithm design. Such a decomposition also plays a fundamental role when one wants to understand the relationship between logic and a graph class. Different notions of the complexity of a separation motivate different ways to decompose, such as tree-decomposition, branch-decomposition, rank-decomposition and carving-decomposition. Furthermore, some important graph classes can be defined through the tree-like decomposition they admit; cographs with cotrees and distance-hereditary graphs with split decompositions being prominent examples.
For a logic $\mathsf{L}$, an $\mathsf{L}$-transduction is a non-deterministic map from a class of relational structures to a new class of relational structures using $\mathsf{L}$-formulas. Transducing a tree-like decomposition is of particular interest. Notably, transducing a decomposition of a graph implies that any property that is definable using a decomposition, is also definable on graphs that admit such a decomposition. Moreover, tree-like decompositions can be often represented by labelled trees, for which the equivalence of recognisability by a tree automaton and definability in CMSO is well known [8]. Hence, it is an interesting question to consider what kind of graph decompositions can be transduced using $\mathsf{L}$-transductions for some extension $\mathsf{L}$ of $\operatorname{MSO}$.
In a series of papers [8, 9, 10, 12], Courcelle investigated the relationship between the graph properties that can be defined in an extension of $\operatorname{MSO}$ and the graph properties that can be recognized by a tree automaton. In particular, for graphs of bounded treewidth, Courcelle’s theorem states that any property that is definable in the logic $\operatorname{MSO}$ with modulo counting predicate, denoted $\operatorname{C}{\!\!}\operatorname{MSO}$, can be recognized by a tree automaton [8]. Combining this result with the linear time algorithm for computing tree-decompositions [2], yields that $\operatorname{C}{\!\!}\operatorname{MSO}$ model-checking can be done in linear time on graphs of bounded treewidth. The converse statement – whether recognisability by a tree automaton implies definability in $\operatorname{C}{\!\!}\operatorname{MSO}$ on graphs of bounded treewidth – was conjectured by Courcelle in [8] and finally settled by Bojańczyk and Pilipczuk [5]. The key step to obtain this result is obtaining a tree-decomposition of a graph via an $\operatorname{MSO}$-transduction, a strategy which was initially proposed in [9] and is now standard.
The obvious next question is whether an analogous result can be proved for graphs of bounded clique-width and for more general combinatorial objects, most notably, matroids representable over a fixed finite field and of bounded branch-width. Due to the above-mentioned strategy, the key challenge is to produce corresponding tree-like decompositions by $\operatorname{MSO}$-transduction. It is known that clique-width decompositions can be $\operatorname{MSO}$-transduced for graphs of bounded linear clique-width [4]. However, it is unknown if clique-width decompositions can be $\operatorname{MSO}$-transduced in general. In fact, this question remains open even for distance-hereditary graphs, which are precisely graphs of rank-width 1 (thus, of constant clique-width).
Besides tree-decompositions, the problem of transducing cotrees, and in general hierarchical decompositions such as modular decompositions and split decompositions were considered in the literature [10, 11, 12, 14]. In [10], Courcelle provides transductions using order-invariant $\operatorname{MSO}$ for cographs and modular decompositions of graphs of bounded modular width. Order-invariant $\operatorname{MSO}$ allows the use of a linear order of the set of vertices and is more expressive than $\operatorname{C}{\!\!}\operatorname{MSO}$ [21]. The applicability of these transductions was later generalized using the framework of weakly-partitive set systems [^1] Weakly-partitive set systems are set systems enjoying some nice closure properties, which were then used to show that some set systems allow canonical tree representations, see for instance the thesis by Montgolfier and Rao [17, 25]. to obtain order-invariant $\operatorname{MSO}$-transductions of modular and split decompositions [12]. It was left as an open question whether one can get rid of the order (see for instance [13] where an overview of the result on hierarchical decompositions was given).
1.1 Our results
In this paper, we consider decompositions given by separations that do not overlap with any other separations of the same type. We view separations of a given kind as a ‘set system’. A set system consists of a set $U$, the universe, and a set $\mathcal{S}$ of subsets of $U$. Two sets overlap if they have non-empty intersection but neither of them is contained in the other. If no two elements in a set system $(U,\mathcal{S})$ overlap, i.e. the set system is laminar, then the subset relation in $(U,\mathcal{S})$ can be described by a rooted tree $T$, called the laminar tree of $(U,\mathcal{S})$. For any set system $(U,\mathcal{S})$ we can look at the subset of strong sets, i.e., sets that do not overlap with any other set, and the laminar tree $T$ they induce. Additionally, we consider systems of bipartitions $\mathcal{B}$ of a universe $U$. Two bipartitions of $U$ overlap if neither side of one of the bipartition is contained in either side of the other bipartition. In case $(U,\mathcal{B})$ has no overlapping bipartitions, then $(U,\mathcal{B})$ can be described by an undirected tree, also called laminar tree, in which bipartitions correspond to edge cuts. We can define the concept of strong bipartitions equivalently and consider the laminar tree induced by the strong bipartitions in $(U,\mathcal{B})$.
Given a graph $G$ we can consider the set system $(V(G),\mathcal{M})$ where $\mathcal{M}$ is the set of all modules in $G$, or the system of bipartition $(V(G),\mathcal{S})$ where $\mathcal{S}$ contains all splits in $G$, or the system of bipartitions $(V(G),\mathcal{B})$ where $\mathcal{B}$ contains all bi-joins in $G$. We obtain the modular decomposition/cotree/split decomposition/bi-join decomposition by equipping the laminar tree of the suitable set system/system of bipartitions with some additional structure. The additional structure allows us to recover the graph from the respective decomposition.
Abstractly, the set systems/systems of bipartitions mentioned are instances of weakly-partitive set systems/weakly-bipartitive systems of bipartitions. Roughly speaking, if a set system $(U,\mathcal{S})$ is well behaved, i.e. $(U,\mathcal{S})$ is weakly-partitive (definition in Section 2), then there is a labelling $\lambda$ of $T$ and a partial order $<$ of its nodes such that $(U,\mathcal{S})$ is completely described by $(T,\lambda,<)$ [7, 25]. A similar statement holds for systems of bipartitions [17].
We show the following, wherein each item $\lambda$ is a suitable labelling of the nodes of the laminar tree $T$, $<$ is a partial ordering of its nodes, and $F$ is an additional edge relation defined only on pairs of siblings in $T$. A visualization how results depend on each other is given in Figure 1.
Theorem 1.1.
There are non-deterministic $\operatorname{C}_{2}{\!\!}\operatorname{MSO}$-transductions $\tau_{1},\ldots,\tau_{7}$ such that:
  1. 1.
    For any laminar set system $(U,\mathcal{S})$, $\tau_{1}(U,\mathcal{S})$ is non-empty and every output in $\tau_{1}(U,\mathcal{S})$ is a laminar tree $T$ of $(U,\mathcal{S})$ (Theorem 3.1);
  2. 2.
    For any graph $G$, $\tau_{2}(G)$ is non-empty and every output in $\tau_{2}(G)$ is a modular decomposition $(T,F)$ of $G$ (Theorem 4.8);
  3. 3.
    For any cograph $G$, $\tau_{3}(G)$ is non-empty and every output in $\tau_{3}(G)$ is a cotree $(T,\lambda)$ of $G$ (Corollary 4.11);
  4. 4.
    For any graph $G$, $\tau_{4}(G)$ is non-empty and every output in $\tau_{4}(G)$ is a split decomposition $(T,F)$ of $G$ (Theorem 5.11);
  5. 5.
    For any graph $G$, $\tau_{5}(G)$ is non-empty and every output in $\tau_{5}(G)$ is a bi-join decomposition $(T,F)$ of $G$ (Theorem 5.15 and Theorem 5.21);
  6. 6.
    For any weakly-partitive set system $(U,\mathcal{S})$, $\tau_{6}(U,\mathcal{S})$ is non-empty and every output in $\tau_{6}(U,\mathcal{S})$ is a weakly-partitive tree $(T,\lambda,<)$ of $(U,\mathcal{S})$ (Theorem 4.4);
  7. 7.
    For any weakly-bipartitive set system $(U,\mathcal{B})$, $\tau_{7}(U,\mathcal{B})$ is non-empty and every output in $\tau_{7}(U,\mathcal{B})$ is a weakly-bipartitive tree $(T,\lambda,<)$ of $(U,\mathcal{B})$ (Theorem 5.5).
The key step in obtaining these transductions is to transduce the laminar tree $T$ of a set system $(U,\mathcal{S})$. The crux here is to find a suitable representative of each node of $T$ amongst the elements of $U$ and a non-deterministic coloring which allows the assignment of representatives to nodes by means of a $\operatorname{C}_{2}{\!\!}\operatorname{MSO}$-formula. It should be mentioned that a similar result is claimed in the preprint [3], where a proof sketch designing a $\operatorname{C}_{3}{\!\!}\operatorname{MSO}$-transduction is described. Once the laminar tree is obtained, the additional relations for each decomposition can be obtained using a deterministic $\operatorname{MSO}$-transduction. Notice that for each of these transductions, there exists an inverse deterministic $\operatorname{MSO}$-transduction, namely a transduction that from the tree-like decomposition outputs the original structure.
Section 3Section 4Section 5 Theorem 3.1blablabla laminar set system $\rightarrow$ laminar tree Lemma 4.1blablabla weakly-partitive set system $\rightarrow$ laminar tree Lemma 5.2blablabla weakly-bipartitive set system $\rightarrow$ laminar tree Theorem 4.4blablabla weakly-partitive set system $\rightarrow$ weakly-partitive tree Theorem 4.8blablabla graph $\rightarrow$ modular decomposition Theorem 5.5blablabla weakly-bipartitive set system $\rightarrow$ weakly-bipartitive tree Theorem 5.11blablabla graph $\rightarrow$ split decomposition Corollary 4.11blablabla cograph $\rightarrow$ cotree Theorem 5.15 & 5.21 graph $\rightarrow$ bi-join decomposition
Figure 1: Overview of the various transductions in the paper. An arrow from $x$ to $y$ indicates that result $x$ is used in the proof of result $y$.
1.2 Organization
The paper is organized as follows. In Section 2 we introduce terminology and notation needed. In Section 3 we prove Theorem 3.1. In Section 4 we provide the proof of Theorem 4.4 and Theorem 4.8 and obtain Corollary 4.11. Both formulation of the theorem as well as the proofs of Theorem 5.5, Theorem 5.11, Theorem 5.15 and Theorem 5.21 are postponed to the appendix.
2 Preliminaries
2.1 Graphs, trees, set systems
Graphs
We use standard terminology of graph theory, and we fix some notations. Given a directed graph $G$, its sets of vertices and edges are denoted by $V(G)$ and $E(G)$, respectively. We denote by uv an edge $(u,v)\in E(G)$. An undirected graph is no more than a directed graph for which $E(G)$ is symmetric (i.e., $uv\in E(G)\iff vu\in E(G)$). The notions of paths, connected components, etc… are defined as usual. Given a subset $Z$ of $V(G)$, we denote by $G[Z]$ the sub-graph of $G$ induced by $Z$.
Trees
A tree is a connected undirected graph without cycles. In the context of trees, we use a slightly different terminology than for graphs. In particular, vertices are called nodes, nodes of degree at most $1$ are called leaves, and nodes of degree greater than $1$ are called inner. The set of leaves is denoted $L(T)$; thus the set of inner nodes is $V(T)\setminus L(T)$. For a node $t$ of a tree $T$ and a neighbor $s$ of $t$, we denote by $T^{t}_{s}$ the connected component of $T-t$ containing $s$. We sometimes consider rooted trees, namely trees with a distinguished node, called the root. Rooted trees enjoy a natural orientation of their edges toward the root, which induces the usual notions of parent, child, sibling, ancestor and descendant. Hence, we represent a rooted tree by a set of nodes with an ancestor/descendant relationship (instead of specifying the root). We use the convention that every node is one of its own ancestors and descendants. We refer to ancestors (resp. descendants) of a node that are not the node itself as proper ancestors (resp. proper descendants). For a node $t$ of a rooted tree $T$, we denote by $T_{t}$ the subtree of $T$ rooted in $t$ (i.e., the restriction of $T$ to the set of descendants of $t$).
Set systems and laminar trees
A set system is a pair $(U,\mathcal{S})$ where $U$ is a finite set, called the universe, and $\mathcal{S}$ is a family of subsets of $U$ where $\emptyset\notin\mathcal{S}$, $U\in\mathcal{S}$, and $\{a\}\in\mathcal{S}$ for every $a\in U$ [^2] Though these restrictions on $\mathcal{S}$ are not usual for set systems, it is convenient for our contribution and it does not significantly impact the generality of set systems: every family $\mathcal{F}$ of subsets of $U$ can be associated with a set system $(U,\mathcal{S})$ where $\mathcal{S}=\big{(}\mathcal{F}\setminus\{\emptyset\}\big{)}\cup\big{\{}U\big{\}}\cup\big{\{}\{a\}\mid a\in U\big{\}}$. Two sets $X$ and $Y$ overlap if they are neither disjoint nor related by containment. A set system $(U,\mathcal{S})$ is said to be laminar (aka overlap-free) when no two sets from $\mathcal{S}$ overlap. By extension, a set family $\mathcal{S}$ is laminar if $\left(\bigcup\mathcal{S},\mathcal{S}\right)$ is a laminar set system (note this also requires that $\emptyset\notin\mathcal{S}$, $\bigcup\mathcal{S}\in\mathcal{S}$, and $\{a\}\in\mathcal{S}$ for every $a\in\bigcup\mathcal{S}$).
A laminar family $\mathcal{S}$ of subsets of $U$ naturally defines a rooted tree where the nodes are the sets from $\mathcal{S}$, the root is $U$, and the ancestor relation corresponds to set inclusion. We call this rooted tree the laminar tree of $U$ induced by $\mathcal{S}$ (or laminar tree of $(U,\mathcal{S})$). In this rooted tree, the leaves are the singletons $\{x\}$ for $x\in U$, which we identify with the elements themselves. That is to say, $L(T)=U$. Laminar trees have the property that each inner node has at least two children. Observe also that the size of a laminar tree is linearly bounded in the size of the universe.
2.2 Logic and transductions
We use relational structures to model both graphs and the various tree-like decompositions used in this paper. In order to concisely model set systems we use the more general notion of extended relational structures, namely relational structure extended with set predicate names. Such structures also naturally arise as outputs of $\operatorname{MSO}$-transductions defined below.
Define an (extended) vocabulary to be a set of symbols, each being either a relation name $R$ with associated arity $\operatorname{ar}(R)\in\mathbb{N}$, or a set predicate name $P$ with associated arity $\operatorname{ar}(P)\in\mathbb{N}$. Set predicate names are aimed to describe relations between sets, e.g., one may have a unary set predicate for selecting finite sets of even size, or a binary set predicate for selecting pairs of disjoint sets. We use capital $R$ or names starting with a lowercase letter (e.g., $\mathsf{edge}$, $\mathsf{ancestor}$, $\mathsf{t}\text{-}\mathsf{edge}$) for relation names, and capital $P$ or uppercase names (e.g., $\mathsf{SET}$, $\operatorname{C}_{2}$) for set predicate names. To refer to an arbitrary symbol of undetermined kind, we use capital $Q$. A relational vocabulary is an extended vocabulary in which every symbol is a relation name.
Let $\Sigma$ be a vocabulary. An extended relational structure over $\Sigma$ ($\Sigma$-structure) is a structure $\mathbb{A}=\langle U_{\mathbb{A}},(Q_{\mathbb{A}})_{Q\in\Sigma}\rangle$ consisting on the one hand of a set $U_{\mathbb{A}}$ called universe, and on the other hand, for each symbol $Q$ in $\Sigma$, an interpretation $Q_{\mathbb{A}}$ of $Q$, which is a relation of arity $\operatorname{ar}(Q)$ either over the universe if $Q$ is a relation name, or over the family of subsets of the universe if $Q$ is a set predicate name. When $\Sigma$ is not extended, $\mathbb{A}$ is simply a relational structure.
Given a $\Sigma$-structure $\mathbb{A}$ and, for some vocabulary $\Gamma$, a $\Gamma$-structure $\mathbb{B}$, we write $\mathbb{A}\sqsubseteq\mathbb{B}$ if $\Sigma\subseteq\Gamma$, $U_{\mathbb{A}}\subseteq U_{\mathbb{B}}$ and for each symbol $Q$ in $\Sigma$, $Q_{\mathbb{A}}=Q_{\mathbb{B}}$. [^3] We require equality here (in particular only elements or subsets of $U_{\mathbb{A}}$ are related in $Q_{\mathbb{B}}$). This differs from classical notions of inclusions of relational structures which typically require equality only on the restriction of the universe to $U_{\mathbb{A}}$, i.e., $Q_{\mathbb{A}}=Q_{\mathbb{B}_{/U_{\mathbb{A}}}}$, e.g., in order to correspond to induced graphs. We write $\mathbb{A}\sqcup\mathbb{B}$ to denote the $(\Sigma\cup\Gamma)$-structure consisting of the universe $U_{\mathbb{A}}\cup U_{\mathbb{B}}$ and, for each symbol $Q\in\Sigma\cup\Gamma$, the interpretation $Q_{\mathbb{A}\sqcup\mathbb{B}}$ which is $Q_{\mathbb{A}}$, $Q_{\mathbb{B}}$, or $Q_{\mathbb{A}}\cup Q_{\mathbb{B}}$ according to whether $Q$ belongs to $\Sigma\setminus\Gamma$, to $\Gamma\setminus\Sigma$, or to $\Sigma\cap\Gamma$ [^4] We do not require $\mathbb{A}$ and $\mathbb{B}$ to be disjoint structures whence we may have $\mathbb{A}\not\sqsubseteq\mathbb{A}\sqcup\mathbb{B}$.
To describe properties of (extended) relational structures, we use monadic second order logic ($\operatorname{MSO}$) and refer for instance to [15, 20, 23, 27] for the definition of $\operatorname{MSO}$ on extended relational structures such as matroids or set systems in general. This logic allows quantification both over single elements of the universe and over subsets of the universe. We also use counting $\operatorname{MSO}$ ($\operatorname{C}{\!\!}\operatorname{MSO}$), which is the extension of $\operatorname{MSO}$ with, for every positive integer $p$, a unary set predicate $\operatorname{C}_{p}$ that checks whether the size of a subset is divisible by $p$ or not. We only use $\operatorname{C}_{2}$. As usual, lowercase variables indicates first-order-quantified variables, while uppercase variables indicates monadic-quantified variables. For a formula $\phi$, we write, e.g.$\phi(x,y,X)$ to indicate that the variables $x$, $y$, and $X$ belong to the set of free variables of $\phi$, namely, the set of variables occurring in $\phi$ that are not bound to a quantifier within $\phi$. A sentence is a formula without free variables.
We now fix some (extended) vocabularies that we will use.
Graphs.
To model both graphs, unrooted trees, and directed graphs, we use the relational vocabulary $\{\mathsf{edge}\}$ where $\mathsf{edge}$ is a relation name of arity $2$. A (directed) graph $G=(V,E)$ is modeled as the $\{\mathsf{edge}\}$-structure $\mathbb{G}$ with universe $U_{\mathbb{G}}=V$ and interpretation $\mathsf{edge}_{\mathbb{G}}=E$. In particular if $G$ is undirected, then $\mathsf{edge}_{\mathbb{G}}$ is symmetric.
Rooted trees.
We use the relational vocabulary $\{\mathsf{ancestor}\}$ to model rooted trees where $\mathsf{ancestor}$ is a relation name of arity $2$. A rooted tree $T$ is modeled as the $\{\mathsf{ancestor}\}$-structure $\mathbb{T}$ with universe $U_{\mathbb{T}}=V(T)$ and the interpretation $\mathsf{ancestor}_{\mathbb{T}}$ being the set of pairs $(u,v)$ such that $u$ is an ancestor of $v$ in $T$. It is routine to define $\operatorname{FO}$-formulas over this vocabulary to express the binary relations parent, child, proper ancestor, (proper) descendant, as well as the unary relations leaf and root.
Set systems.
To model set systems, we use the extended vocabulary $\{\mathsf{SET}\}$ where $\mathsf{SET}$ is a set predicate name of arity $1$. A set system $S=(U,\mathcal{S})$ is thus naturally modeled as the $\{\mathsf{SET}\}$-structure $\mathbb{S}$ with universe $U_{\mathbb{S}}=U$ and interpretation $\mathsf{SET}_{\mathbb{S}}=\mathcal{S}$.
Transductions
Let $\Sigma$ and $\Gamma$ be two extended vocabularies. A $\Sigma$-to-$\Gamma$ transduction is a set $\tau$ of pairs formed by a $\Sigma$-structure, call the input, and a $\Gamma$-structure, called the output. We write $\mathbb{B}\in\tau(\mathbb{A})$ when $(\mathbb{A},\mathbb{B})\in\tau$. When for every pair $(\mathbb{A},\mathbb{B})\in\tau$ we have $\mathbb{A}\sqsubseteq\mathbb{B}$, we call $\tau$ an overlay transduction. Some transductions can be defined by means of $\operatorname{MSO}$- or $\operatorname{C}_{2}{\!\!}\operatorname{MSO}$-formulas. This leads to the notion of $\operatorname{MSO}$- and $\operatorname{C}_{2}{\!\!}\operatorname{MSO}$-transductions. Following the presentation of [4], for $\mathsf{L}$ denoting $\operatorname{MSO}$ or $\operatorname{C}_{2}{\!\!}\operatorname{MSO}$, define an $\mathsf{L}$-transduction to be a transduction obtained by composing a finite number of atomic $\mathsf{L}$-transductions of the following kinds.
Filtering.
An overlay transduction specified by an $\mathsf{L}$-sentence $\phi$ over the input vocabulary $\Sigma$, which discards the inputs that do not satisfy $\phi$ and keeps the other unchanged. Hence, it defines a partial function (actually, a partial identity) from $\Sigma$-structures to $\Sigma$-structures.
Universe restriction.
A transduction specified by a $\mathsf{L}$-formula $\phi$ over the input vocabulary $\Sigma$, with one free first-order variable, which restricts the universe to those elements that satisfy $\phi$. The output vocabulary is $\Sigma$ and the interpretation of every relation (resp. every predicate) in the output structure is defined as the restriction of its interpretation in the input structure to those tuples of elements satisfying $\phi$ (resp. tuples of sets of elements that satisfy $\phi$). This defines a total function from $\Sigma$-structures to $\Sigma$-structures.
Interpretation.
A transduction specified by a family $(\phi_{Q})_{Q\in\Gamma}$ over the input vocabulary $\Sigma$ where $\Gamma$ is the output vocabulary and each $\phi_{Q}$ has $\operatorname{ar}(Q)$ free variables which are first-order if $Q$ is a relation name and monadic if it is a set predicate name. The transduction outputs the $\Gamma$-structure that has the same universe as the input structure and in which each relation or predicate $Q$ is interpreted as those set of tuples that satisfy $\phi_{Q}$. This defines a total function from $\Sigma$-structures to $\Gamma$-structures.
Copying.
An overlay transduction parametrized by a positive integer $k$ that adds $k$ copies of each element to the universe. The output vocabulary consists in the input vocabulary $\Sigma$ extended with $k$ binary relational symbols $(\mathsf{copy}_{i})_{i\in[k]}$ interpreted as pairs of elements $(x,y)$ saying that “$y$ is the $i$-th copy of $x$”. The interpretation of the relations (resp. predicates) of the input structure are preserved, on original elements. This defines a total function from $\Sigma$-structures to $\Gamma$-structures, where $\Gamma=\Sigma\cup\{\mathsf{copy}_{i}\mid i\in[k]\}$.
Colouring.
An overlay transduction that adds a new unary relation $\mathsf{color}\notin\Sigma$ to the structure. Any possible interpretation yields an output; indeed the interpretation is chosen non-deterministically. The interpretation of the relations (resp. predicates) of the input structure are preserved. Hence, it defines a total (non-functional) relation from $\Sigma$-structures to $\Gamma$-structures where $\Gamma=\Sigma\cup\{\mathsf{color}\}$ (providing $\mathsf{color}\notin\Sigma$).
We say an $\mathsf{L}$-transduction is deterministic if it does not use colouring, it is non-deterministic otherwise. By definition, deterministic $\mathsf{L}$-transductions define functions.
3 Transducing the laminar-tree
In this section, we present an overlay $\operatorname{C}_{2}{\!\!}\operatorname{MSO}$-transduction that takes as input a laminar set system and outputs the laminar tree it induces.
Theorem 3.1.
Let $\Sigma$ be an extended vocabulary, including a unary set predicate name $\mathsf{SET}$ and not including the binary relational symbol $\mathsf{ancestor}$. There exists a non-deterministic $\operatorname{C}_{2}{\!\!}\operatorname{MSO}$-transduction $\tau$ such that, for each laminar set system $(U,\mathcal{S})$ represented as the $\{\mathsf{SET}\}$-structure $\mathbb{S}$ and inducing the laminar tree $T$ with ${L(T)=U}$, and for each $\Sigma$-structure $\mathbb{A}$ with $\mathbb{S}\sqsubseteq\mathbb{A}$, $\tau(\mathbb{A})$ is non-empty and every output in $\tau(\mathbb{A})$ is equal to some $\{\mathsf{ancestor}\}$-structure $\mathbb{T}$ representing $T$.
Since the sets from $\mathcal{S}$ are precisely the sets of leaves of the subtrees of $T$, there is an $\operatorname{MSO}$-transduction which is the inverse of the above $\operatorname{C}_{2}{\!\!}\operatorname{MSO}$-transduction; that is to say, given an $\{\mathsf{ancestor}\}$-structure $\mathbb{T}$ representing the laminar tree, it outputs the original set system $\mathbb{S}$. Namely,
  1. 1.
    An interpretation $\Phi_{\mathsf{SET}}(S)$ that is true for a set $S$ when there exists an element $a$ such that an element $x$ is in $S$ if and only if $a$ is an ancestor of $x$.
  2. 2.
    The filtering $\phi(x)$ keeping leaves only.
We prove Theorem 3.1 in Section 3.2, using the key tools developed in Section 3.1, that allow us to represent each inner node of $T$ with a pair of leaves from its subtree, while keeping the number of inner nodes represented by a given leaf bounded.
3.1 Inner node representatives
In this section, the root of any rooted tree is always an inner node (unless the tree is a unique node). We fix a rooted tree $T$, in which every inner node has at least two children (a necessary assumption that is satisfied by laminar trees), and we let $V$ denote the set of its nodes ($V=V(T)$) and $L\subseteq V$ denote the subset of its leaves ($L=L(T)$).
Let $S\subseteq V\setminus L$, and let $(\pi,\sigma)$ be a pair of injective mappings from $S$ to $L$. We say that the pair $(\pi,\sigma)$ identifies $S$ if for each $s\in S$, $s$ is the least common ancestor of $\pi(s)$ and $\sigma(s)$. For $s\in S$ and $x\in V$, we say that $x$ is $s$-requested in $(\pi,\sigma)$ if $x$ lies on the path from $\pi(s)$ to $\sigma(s)$ (namely, on either of the paths from $\pi(s)$ to $s$ and from $\sigma(s)$ to $s$). We say that $x$ is requested in $(\pi,\sigma)$ if it is $s$-requested for some $s$. The pair $(\pi,\sigma)$ has unique request if every node $x$ of $T$ is requested at most once in $(\pi,\sigma)$, i.e., there exists at most one $s\in S$ such that $x$ is $s$-requested in $(\pi,\sigma)$. Note that if $(\pi,\sigma)$ has unique request then the paths from $\pi(s)$ to $\sigma(s)$ are pairwise disjoint for all the $s\in S$. We now state a few basic observations.
Remark 3.2.
Let $(\pi,\sigma)$ identifying some subset $S$ of $V\setminus L$.
  1. 1.
    The reversed pair $(\sigma,\pi)$ also identifies $S$ and has unique request whenever $(\pi,\sigma)$ does.
  2. 2.
    If $S^{\prime}\subset S$, then $(\pi_{|S^{\prime}},\sigma_{|S^{\prime}})$ identifies $S^{\prime}$, and has unique request if $(\pi,\sigma)$ does.
  3. 3.
    For each $s\in S$, $s$ is $s$-requested in $(\pi,\sigma)$.
  4. 4.
    If $(\sigma,\pi)$ has unique request, then $\pi(S)$ and $\sigma(S)$ are disjoint subsets of $L$.
Lemma 3.3.
Let $(\pi,\sigma)$ identifying some subset $S$ of $V\setminus L$ with unique request. Then for each $a\in\pi(S)$ the node $\pi^{-1}(a)$ is the least ancestor of $a$ which belongs to $S$.
Proof 3.4.
Let $a\in\pi(S)$ and let $s=\pi^{-1}(a)$. By definition, $s$ is an ancestor of $a$ which belongs to $S$. Let $y$ be the least ancestor of $a$ that is contained in $S$. As $s$ is an ancestor of $a$ belonging to $S$, $y$ must be a descendant of $s$. Hence, $y$ is $s$-requested in $(\pi,\sigma)$. Additionally, by Item 3 of Remark 3.2, $y$ is $y$-requested in $(\pi,\sigma)$. Since $(\pi,\sigma)$ has unique request, $y=s$. Thus, $s$ is the least ancestor of $a$ which belongs to $S$.
Notice that, by Item 1 of Remark 3.2, a similar result holds for each $b\in\sigma(S)$. It follows that the sets $\pi(S)$ and $\sigma(S)$ characterize $(\pi,\sigma)$.
Lemma 3.5.
Let $S\subseteq V\setminus L$, and $(\pi,\sigma)$ and $(\pi^{\prime},\sigma^{\prime})$ be two pairs of injections from $S$ to $L$ identifying $S$ with unique request. If $\pi(S)=\pi^{\prime}(S)$ and $\sigma(S)=\sigma^{\prime}(S)$ then $\pi=\pi^{\prime}$ and $\sigma=\sigma^{\prime}$.
Proof 3.6.
Let $s\in S$, $a=\pi(s)$, and $s^{\prime}=\pi^{\prime-1}(a)$. Both $s$ and $s^{\prime}$ are the least ancestor of $a$ which belongs to $S$, hence $s=s^{\prime}$ and $\pi^{\prime}(s)=a$. Thus $\pi=\pi^{\prime}$. By Item 1 of Remark 3.2, we also obtain that $\sigma=\sigma^{\prime}$.
Let $A$ and $B$ be two disjoint subsets of $L$. We call such a pair $(A,B)$ a bi-colouring. We say that $(A,B)$ identifies $S$ if there exists a pair $(\pi,\sigma)$ identifying $S$ with unique request such that $\pi(S)=A$ and $\sigma(S)=B$. By the previous lemma, for a fixed set $S\subseteq V\setminus L$ the pair $(\pi,\sigma)$ identifying $S$ is unique when it exists. We will also prove that $S$ is actually uniquely determined from $(A,B)$. Before that, we state the following technical lemma.
Lemma 3.7.
Let $(A,B)$ identify some subset $S$ of $V\setminus L$ through a pair $(\pi,\sigma)$ of injections having unique request. Then, for each inner node $x$, exactly one of the three following cases holds:
  1. 1.
    $x\notin S$, $x$ is not requested in $(\pi,\sigma)$, and for each child $c$ of $x$, $\left|A\cap V(T_{c})\right|=\left|B\cap V(T_{c})\right|$;
  2. 2.
    $x\notin S$, $x$ is requested in $(\pi,\sigma)$, and there exists one leaf $z\in(A\cup B)\cap V(T_{x})$ such that, for each child $c$ of $x$, $\left|(A\setminus\{z\})\cap V(T_{c})\right|=\left|(B\setminus\{z\})\cap V(T_{c})\right|$;
  3. 3.
    $x\in S$, $x$ is requested in $(\pi,\sigma)$, and there exists two leaves $a\in A\cap V(T_{x})$ and $b\in B\cap V(T_{x})$ such that, for each child $c$ of $x$, $\left|(A\setminus\{a\})\cap V(T_{c})\right|=\left|(B\setminus\{b\}\cap V(T_{c}))\right|$ and $\{a,b\}\nsubseteq V(T_{c})$.
Proof 3.8.
Let $x$ be an inner node. We consider the set $S^{\prime}=S\setminus(V(T_{x})\setminus\{x\})$ of all nodes which are not proper descendants of $x$ and the restrictions $\pi^{\prime}$ and $\sigma^{\prime}$ of, respectively, $\pi$ and $\sigma$ to $S^{\prime}$. By Item 2 of Remark 3.2 $(\pi^{\prime},\sigma^{\prime})$ identify $S^{\prime}$ with unique request. We denote $A^{\prime}=\pi^{\prime}(S^{\prime})$ and $B^{\prime}=\sigma^{\prime}(S^{\prime})$, thus $(A^{\prime},B^{\prime})$ identify $S^{\prime}$. Let $A^{\prime}_{x}=A^{\prime}\cap V(T_{x})$ and $B^{\prime}_{x}=B^{\prime}\cap V(T_{x})$ be the sets of representatives contained in $T_{x}$ of nodes in $S^{\prime}$. Let $a$ be an element of $A^{\prime}_{x}$. The element $s_{a}=\pi^{-1}(a)$ is an ancestor of $x$ because it is an ancestor of $a\in V(T_{x})$ and belongs to $S^{\prime}$ whence not to $V(T_{x})\setminus\{x\}$. Therefore, $x$ is $s_{a}$-requested. As $(\pi^{\prime},\sigma^{\prime})$ has unique request, there exists at most one element in $A^{\prime}_{x}$. Similarly, $B^{\prime}_{x}$ has size at most $1$. Moreover, $A^{\prime}_{x}\cap B^{\prime}_{x}=\emptyset$ by Item 4 of Remark 3.2. We thus we have three cases:
Case $A^{\prime}_{x}\cup B^{\prime}_{x}=\emptyset$.
Then there is no $s\in S^{\prime}$ such that $\pi(s)\in V(T_{x})$ or $\sigma(s)\in V(T_{x})$. Hence, $x$ is not requested in $(\pi,\sigma)$, in particular, $x\notin S$.
Let $c$ be a child of $x$. If $\left|A\cap V(T_{c})\right|\neq\left|B\cap V(T_{c})\right|$, then there exists $a\in(A\cup B)\cap V(T_{c})$ such that, assuming without loss of generality that $a\in A$ and denoting $s_{a}=\pi^{-1}(a)$, $\sigma(s_{a})\notin V(T_{c})$. Hence, $s_{a}$ is a proper ancestor of $c$, thus, equivalently, an ancestor of $x$, implying that $x$ is $s_{a}$-requested in $(\pi,\sigma)$. This contradicts the above argument. So, $\left|A\cap V(T_{c})\right|=\left|B\cap V(T_{c})\right|$ for each child $c$ of $x$.
Case $A^{\prime}_{x}\cup B^{\prime}_{x}=\{a\}$.
Assume, without loss of generality, that $a\in A$, and denote $s_{a}=\pi^{-1}(a)$ and $b=\sigma(s_{a})$. We have that $s_{a}$ is a proper ancestor of $x$, since it is the least common ancestor of $a\in V(T_{x})$ and $b\notin V(T_{x})$. Thus, $x$ is $s_{a}$-requested and $s_{a}\neq x$ so $x\notin S$.
Let $c$ be a child of $x$. If $\left|A\cap V(T_{c})\right|\neq\left|B\cap V(T_{c})\right|$, then it means that there exists $z\in(A\cup B)\cap V(T_{c})$, such that either $z\in A$ and $s_{z}=\pi^{-1}(z)\notin V(T_{c})$, or $z\in B$ and $s_{z}=\sigma^{-1}(z)\notin V(T_{c})$. In both cases, $s_{z}$ is a proper ancestor of $c$, whence an ancestor of $x$. Thus, $x$ is $s_{z}$-requested in $(\pi,\sigma)$. However, $x$ is $s_{a}$-requested in $(\pi,\sigma)$ which has unique request, hence $s_{z}=s_{a}$. If $z\in B$, it follows that $z=b$, which contradicts the fact $b\notin V(T_{x})$. Hence, $z\in A$ and thus, $z=\pi(s_{a})=a$. Therefore, $\left|(A\setminus\{a\})\cap V(T_{c})\right|=\left|B\cap V(T_{c})\right|$ for each child $c$ of $x$.
Case $A^{\prime}_{x}=\{a\}$ and $B^{\prime}_{x}=\{b\}$.
Let $s_{a}=\pi^{-1}(a)$ and $s_{b}=\sigma^{-1}(b)$. The node $x$ is $s_{a}$-requested and $s_{b}$-requested in $(\pi,\sigma)$, so, by the unique request property, $s_{a}=s_{b}$. Since $s_{a}$ is the least common ancestor of $a$ and $b$, it belongs to $V(T_{x})$ whence $s_{a}=x$ implying $x\in X$.
Let $c$ be a child of $x$. If $\left|A\cap V(T_{c})\right|\neq\left|B\cap V(T_{c})\right|$, then there exists $z\in(A\cup B)\cap V(T_{c})$ such that either $z\in A$ and $s_{z}=\pi^{-1}(z)\notin V(T_{c})$, or $z\in B$ and $s_{z}=\sigma^{-1}(z)\notin V(T_{c})$. In both cases, $s_{z}$ is a proper ancestor of $c$, whence an ancestor of $x$, and thus $x$ is $s_{z}$-requested in $(\pi,\sigma)$. However, $x$ is $x$-requested in $(\pi,\sigma)$ which has unique request, hence $s_{z}=x$ and thus $z\in\{a,b\}$. Therefore, $\left|(A\setminus\{a\})\cap V(T_{c})\right|=\left|(B\setminus\{b\})\cap V(T_{c})\right|$ for each child $c$ of $x$. Because the least common ancestor of $a$ and $b$ is $x$, there is no child $c$ of $x$ containing both $a$ and $b$ as leaves, i.e., $\{a,b\}\nsubseteq V(T_{c})$.
This concludes the proof of the statement.
It follows that for each bi-colouring $(A,B)$, there exists at most one set $S$ of inner nodes identified by $(A,B)$.
Lemma 3.9.
Let $(A,B)$ be two disjoint subsets of $L$ and let $S$ and $S^{\prime}$ be two subsets of $V\setminus L$. If $(A,B)$ identify both $S$ and $S^{\prime}$, then $S=S^{\prime}$.
Proof 3.10.
We proceed by contradiction and thus assume $S\neq S^{\prime}$. Let $s\in S\setminus S^{\prime}$. By Lemma 3.7, the number of children $c$ of $s$ for which the set $(A\cup B)\cap V(T_{c})$ has odd size is $2$ since $s\in S$, while it should also be less than $2$ since $s\notin S^{\prime}$. A contradiction.
When $(A,B)$ identifies a set $S$, we call $A$-representative (resp. $B$-representative) of $s\in S$ the leaf $\pi(s)\in A$ (resp. $\sigma(s)\in B$), where $(\pi,\sigma)$ witnessing that $(A,B)$ identifies $S$. An example of bi-colouring identifying a subset of inner nodes is given in Figure 2.
Figure 2: Illustration of a bi-colouring $(A,B)$ identifying some set $S\subseteq V\setminus L$ in a binary tree $T$. Leaves from $A$ are coloured blue, leaves from $B$ are coloured red, and inner nodes from $S$ are marked purple. Furthermore, the two paths connecting an inner node $s\in S$ to its $A$- and $B$-representatives are coloured blue and red, respectively.
Not every set of inner nodes has a bi-colouring identifying it. To ensure that such a pair exists, we consider thin sets of inner nodes. While thin sets always have bi-colourings identifying them, it is also guaranteed that the set of inner nodes can be partitioned into only 4 thin sets. A subset $X\subseteq V\setminus L$ is thin when, for each $x\in X$ not being the root, on the one hand, the parent $p_{x}$ of $x$ does not belong to $X$, and on the other hand, $x$ admits at least one sibling (including possible leaves) that does not belong to $X$. Having a thin set $X$ allows to find branches avoiding it.
Lemma 3.11.
Let $X\subseteq V\setminus L$ and $s\in V\setminus X$. If $X$ is thin, then there exists a leaf $t\in V(T_{s})$ such that the path from $t$ to $s$ avoids $X$ (i.e., none of the nodes along this path belong to $X$).
Proof 3.12.
If $T_{s}$ has height 00, then $s$ is a leaf and taking $t=s$ trivially gives the expected path. Otherwise, $s$ is an inner node and, because $X$ is thin, $s$ has at least one child $c_{s}$ not belonging to $X$. By induction, there is a path from some leaf $t\in V(T_{c_{s}})$ to $c_{s}$ avoiding $X$ and, since $s\notin X$, this path could be extended into a path from $t$ to $s$ avoiding $X$.
The previous lemma allows to identify every thin set.
Lemma 3.13.
If $X$ is a thin set, then there exists a pair $(\pi,\sigma)$ of injections from $X$ to $L$ that has unique request and that identifies $X$.
Proof 3.14.
We proceed by induction on the size of $X$. If $X=\emptyset$, the result is trivial. Let $n\in\mathbb{N}$ and suppose that for every thin set of size $n$ there exists a pair of injections identifying it with unique request. Let $X$ be a thin set of size $n+1$, and let $s\in X$ be of minimal depth. Clearly, $X\setminus\{s\}$ is thin and thus there exists, by induction, a pair $(\pi,\sigma)$ of injections from $X\setminus\{s\}$ to $L$ identifying $X\setminus\{s\}$ with unique request. Since $X$ is thin and $s\in X$, we can find two distinct children $c_{a}$ and $c_{b}$ of $s$ not belonging to $X$ [^5] Remember that every inner node of $T$ has at least two children (including possible leaves). Then, by Lemma 3.11, there exists a leaf $a\in V(T_{c_{a}})$ (resp. a leaf $b\in V(T_{c_{b}})$) such that the path from $a$ to $c_{a}$ (resp. from $b$ to $c_{b}$) avoids $X$. In particular, for each node $y$ along these paths, since $y$ has no ancestor that belongs to $X$ but $s$, $y$ is not requested in $(\pi,\sigma)$. Hence, extending $\pi$ (resp. $\sigma$) in such a way that, besides mapping each $x\in X\setminus\{s\}$ to $\pi(x)$ (resp. to $\sigma(x)$), it maps $s$ to $a$ (resp. to $b$), we obtain a pair $(\hat{\pi},\hat{\sigma})$ of injections from $X$ to $L$ that identifies $X$ with unique request.
A family $F=(A_{1},B_{1}),\ldots,(A_{n},B_{n})$ of bi-colourings identifies a set $S\subseteq V\setminus L$, if there exists a partition $(S_{1},\ldots,S_{n})$ of $S$ such that, for each $i\in[n]$, $(A_{i},B_{i})$ identifies $S_{i}$. Whenever $S=V\setminus L$ we say that $F$ identifies $T$. A collection of subsets of $V\setminus L$ is thin if each of its subsets is thin. We now show that there exists a thin $4$-partition.
Lemma 3.15.
There exists a thin $4$-partition of $V\setminus L$.
Proof 3.16.
We build such a thin $4$-partition as follows. First, consider the partition $(D_{e},D_{o})$ of $V\setminus L$ in which $D_{e}$ (resp. $D_{o}=V\setminus(D_{e}\cup L)$) is the set of all inner nodes of even (resp. odd) depth. Second, arbitrarily fix one child $c_{x}$ of $x$ for each inner node $x$, and consider the set $C=\{c_{x}\mid x\in V\setminus L\}\setminus L$, inducing a partition $(C,\overline{C})$ of $V\setminus L$ where $\overline{C}=V\setminus(L\cup C)$. By refining these two bi-partitions, we obtain a $4$-partition which is thin by construction.5
Hence, four bi-colourings are enough to identify $T$. For an example see Figure 3.
Corollary 3.17.
There exists a family of four bi-colourings identifying $T$.
Proof 3.18.
The result immediately follows from Lemmas 3.13 and 3.15.
Figure 3: Illustration of a partition of the inner nodes into 4 thin sets indicated by different shapes/fillings. The four colourings identifying each thin set are indicated in the table below the leaves.
3.2 The transduction
The goal of this section is to prove Theorem 3.1, that is, to design a $\operatorname{C}_{2}{\!\!}\operatorname{MSO}$-transduction that produces the laminar tree induced by an input laminar set system. We fix a laminar set system $(U,\mathcal{S})$, represented by a $\{\mathsf{SET}\}$-structure $\mathbb{S}$. Before proving the theorem, we make the following basic observation. On $\mathbb{S}$, we can define two $\operatorname{MSO}$-formulas $\mathsf{desc}(X,Y)$ and $\mathsf{child}(X,Y)$ expressing that in the laminar tree induced by $\mathcal{S}$, $X$ and $Y$ are nodes and $X$ is a descendant or a child of $Y$, respectively:
$\mathsf{desc}(X,Y)$ $:=\mathsf{SET}(X)\land\mathsf{SET}(Y)\land X\subseteq Y;$
$\mathsf{child}(X,Y)$ $:=\mathsf{desc}(X,Y)\land X\neq Y\land\forall Z\>\Big{(}\big{(}\mathsf{desc}(Z,Y)\land Z\neq Y\big{)}\rightarrow\mathsf{desc}(Z,X)\Big{)}.$
[…]
Lemma 3.19.
[…]
Proof 3.20.
First, we define a formula $\phi_{A,B}(X)$ that, under the above assumption, is satisfied exactly when $X$ belongs to $S$. According to Lemma 3.7, this happens if and only if $X$ is a node of $T$ (i.e.$\mathsf{SET}(X)$ is satisfied) and there exists $a\in X\cap A$ and $b\in X\cap B$ such that for each child $Z$ of $X$, $\{a,b\}\nsubseteq Z$ and the set $(Z\setminus\{a,b\})\cap(A\cup B)$ has even size. This property is easily expressed in $\operatorname{C}_{2}{\!\!}\operatorname{MSO}$, using the $\operatorname{MSO}$-formula $\mathsf{child}(X,Y)$ defined previously, as well as the predicate $\mathsf{SET}$:
$\phi_{A,B}(X):=\>$ $\mathsf{SET}(X)\land\exists a\exists b\Big{[}a\in(X\cap A)\land b\in(X\cap B)\land$
$\qquad\forall Z\>\Big{(}\mathsf{child}(Z,X)\rightarrow\Big{(}\{a,b\}\nsubseteq Z\land\operatorname{C}_{2}\big{(}(Z\setminus\{a,b\})\cap(A\cup B)\big{)}\Big{)}\Big{)}\Big{]}.$
[…]
This concludes the proof.
We are now ready to prove the theorem.
Proof 3.21 (Proof of Theorem 3.1).
[…]
  1. 1.
    Guess a family of four bi-colourings $(A_{i},B_{i})_{i\in[4]}$ identifying $T$ (which exists by Corollary 3.17).
  2. 2.
    Copy the input graph four times, thus introducing four binary relations $(\mathsf{copy}_{i})_{i\in[4]}$ where $\mathsf{copy}_{i}(x,y)$ indicates that $x$ is the $i$-th copy of the original element $y$.
  3. 3.
[…]
  4. 4.
[…]
4 Transducing modular decompositions
The set of modules of a directed graph is a specific example of a particular type of set system, a “weakly-partitive set system” (and a “partitive set system” in the case of an undirected graph). In this section, we first give a general $\operatorname{C}_{2}{\!\!}\operatorname{MSO}$-transduction to obtain the canonical tree-like decomposition of a weakly-partitive set system from the set system itself. We then show how to obtain the modular decomposition of a graph via a $\operatorname{C}_{2}{\!\!}\operatorname{MSO}$-transduction as an application.
4.1 Transducing weakly-partitive trees
A set system $(U,\mathcal{S})$ is weakly-partitive if for every two overlapping sets $X,Y\in\mathcal{S}$, the sets $X\cup Y$, $X\cap Y$, $X\setminus Y$, and $Y\setminus X$ belong to $\mathcal{S}$. It is partitive if, moreover, for every two overlapping sets $X,Y\in\mathcal{S}$, their symmetric difference, denoted $X\operatorname*{\bigtriangleup}Y$, also belongs to $\mathcal{S}$. By extension, a set family $\mathcal{S}$ is called weakly-partitive or partitive whenever $\left(\bigcup\mathcal{S},\mathcal{S}\right)$ is a set system which is weakly-partitive or partitive, respectively (note these also requires that $\emptyset\notin\mathcal{S}$, $\bigcup\mathcal{S}\in\mathcal{S}$, and $\{a\}\in\mathcal{S}$ for every $a\in\bigcup\mathcal{S}$).
A member of a set system $(U,\mathcal{S})$ is said to be strong if it does not overlap any other set from $\mathcal{S}$. The sub-family $\mathcal{S}_{!}$ of strong sets of $\mathcal{S}$ is thus laminar by definition. Hence, it induces a laminar tree $T$. By extension, we say that $T$ is induced by the set system $(U,\mathcal{S})$ (or simply by $\mathcal{S}$). The next result extends Theorem 3.1, by showing that $T$ can be $\operatorname{C}_{2}{\!\!}\operatorname{MSO}$-transduced from $\mathcal{S}$.
Lemma 4.1.
Let $\Sigma$ be an extended vocabulary, including a set unary predicate name $\mathsf{SET}$ and not including the binary relational symbol $\mathsf{ancestor}$. There exists a non-deterministic $\operatorname{C}_{2}{\!\!}\operatorname{MSO}$-transduction $\tau$ such that, for each set system $(U,\mathcal{S})$ represented as the $\{\mathsf{SET}\}$-structure $\mathbb{S}$ and inducing the laminar tree $T$ with ${L(T)=U}$, and for each $\Sigma$-structure $\mathbb{A}$ with $\mathbb{S}\sqsubseteq\mathbb{A}$, $\tau(\mathbb{A})$ is non-empty and every output in $\tau(\mathbb{A})$ is equal to some $\{\mathsf{ancestor}\}$-structure $\mathbb{T}$ representing $T$.
Proof 4.2.
On the $\{\mathsf{SET}\}$-structure $\mathbb{S}$, it is routine to define an $\operatorname{MSO}$-formula $\phi_{\mathsf{SET!}}(Z)$ that identifies those subsets $Z\subseteq U$ that are strong members of $\mathcal{S}$:
$\phi_{\mathsf{SET!}}(Z)$ $:=\mathsf{SET}(Z)\land\forall X\>\Big{(}\big{(}\mathsf{SET}(X)\land X\cap Z\neq\emptyset\big{)}\rightarrow\big{(}X\subseteq Z\lor Z\subseteq X\big{)}\Big{)}.$
Hence, we can design an $\operatorname{MSO}$-interpretation that outputs the $\Sigma\cup\{\mathsf{SET!}\}$-structure corresponding to $\mathbb{A}$ equipped with the set unary predicate $\mathsf{SET!}$ that selects strong members of $\mathcal{S}$. Thus, up to renaming the predicates $\mathsf{SET!}$ and $\mathsf{SET}$, by Theorem 3.1, we can produce, through a $\operatorname{C}_{2}{\!\!}\operatorname{MSO}$-transduction, the $\Sigma\cup\{\mathsf{ancestor}\}$-structure $\mathbb{A}\sqcup\mathbb{T}$ where $\mathbb{T}$ is the tree-structure modeling the laminar tree $T$ induced by $\mathcal{S}_{!}$ with $L(T)=U$ (once the output obtained, the interpretation drops the set predicate $\mathsf{SET!}$ which is no longer needed).
Clearly, the laminar tree $T$ of a weakly-partitive set system $(U,\mathcal{S})$ does not characterize $(U,\mathcal{S})$. However, as shown by the below theorem, a labeling of its inner nodes and a controlled partial ordering of its nodes are sufficient to characterize all the sets of $\mathcal{S}$. For $Z$ a set equipped with a partial order $<$ and $X$ a subset of $Z$, we say that $X$ is a $<$-interval whenever $<$ defines a total order on $X$ and for every $a,b\in X$ and every $c\in Z$, $a<c<b$ implies $c\in X$.
Theorem 4.3 ([7, 25]).
Let $\mathcal{S}$ be a weakly-partitive family, $\mathcal{S}_{!}$ be its subfamily of strong sets, and $T$ be the laminar tree it induces. There exists a total labeling function $\lambda$ from the set $V(T)\setminus L(T)$ of inner nodes of $T$ to the set $\{\mathsf{degenerate},\mathsf{prime},\mathsf{linear}\}$, and, for each inner node $t\in\lambda^{-1}(\mathsf{linear})$, a linear ordering $<_{t}$ of its children, such that every inner node having exactly two children is labeled by $\mathsf{degenerate}$ and the following two conditions are satisfied:
  • for each $X\in\mathcal{S}\setminus\mathcal{S}_{!}$, there exists $t\in V(T)$ and a subset $\mathcal{C}$ of children of $t$ such that $X=\bigcup_{c\in\mathcal{C}}L(T_{c})$ and either $\lambda(t)=\mathsf{linear}$ and $\mathcal{C}$ is a $<_{t}$-interval, or $\lambda(t)=\mathsf{degenerate}$;
  • conversely, for each inner node $t$ and each non-empty subset $\mathcal{C}$ of children of $t$, if either $\lambda(t)=\mathsf{linear}$ and $\mathcal{C}$ is a $<_{t}$-interval, or $\lambda(t)=\mathsf{degenerate}$, then $\bigcup_{c\in\mathcal{C}}L(T_{c})\in\mathcal{S}$.
Furthermore, $T$ and $\lambda$ are uniquely determined from $\mathcal{S}$, and, for each inner node $t$ of $T$ labeled by $\mathsf{linear}$, only two orders $<_{t}$ are possible, one being the inverse of the other (indeed, inverting an order $<$ does preserve the property of being a $<$-interval). Hence, every weakly-partitive family $\mathcal{S}$ is characterized by a labeled and partially-ordered tree $(T,\lambda,<)$ where $T$ is the laminar tree induced by the subfamily $\mathcal{S}_{!}$ of strong sets of $\mathcal{S}$, $\lambda:V(T)\setminus L(T)\to\{\mathsf{degenerate},\mathsf{prime},\mathsf{linear}\}$ is the labeling function, and $<$ is the partial order $\bigcup_{t\in\lambda^{-1}(\mathsf{linear})}<_{t}$ over $V(T)$. As, up-to inverting some of the $<_{t}$ orders, $(T,\lambda,<)$ is unique, we abusively call it the weakly-partitive tree induced by $\mathcal{S}$. Conversely, a weakly-partitive tree characterizes the unique weakly-partitive set system which induced it.
We naturally model a weakly-partitive tree $(T,\lambda,<)$ of a partitive set system $(U,\mathcal{S})$ by the $\{\mathsf{ancestor},\mathsf{degenerate},\mathsf{betweeness}\}$-structure $\mathbb{T}$ of universe $U_{\mathbb{T}}=V(T)$ such that $\langle V(T),\mathsf{ancestor}_{\mathbb{T}}\rangle\sqsubseteq\mathbb{T}$ models $T$ with ${L(T)=U}$, $\mathsf{degenerate}_{\mathbb{T}}$ is a unary relation which selects the inner nodes of $T$ of label $\mathsf{degenerate}$, i.e., ${\mathsf{degenerate}_{\mathbb{T}}=\lambda^{-1}(\mathsf{degenerate})}$, and $\mathsf{betweeness}_{\mathbb{T}}$ is a ternary relation selecting triples $(x,y,z)$ satisfying $x<y<z$ or $z<y<x$. (Although it is possible, through a non-deterministic $\operatorname{MSO}$-transduction, to define $<$ from $\mathsf{betweeness}_{\mathbb{T}}$, the use of $\mathsf{betweeness}_{\mathbb{T}}$ rather than $<_{\mathbb{T}}$ ensures unicity of the output weakly-partitive tree.) The inner nodes of $T$ that are labeled by $\mathsf{linear}$ can be recovered through an $\operatorname{MSO}$-formula as those inner nodes whose children are related by $\mathsf{betweeness}$. The inner nodes labeled by $\mathsf{prime}$ can be recovered through an $\operatorname{MSO}$-formula as those inner nodes that are labeled neither by $\mathsf{degenerate}$ nor by $\mathsf{linear}$. Using Theorem 4.3, it is routine to design an $\operatorname{MSO}$-transduction which takes as input a weakly-partitive tree and outputs the weakly-partitive set system which induced it. The inverse $\operatorname{C}_{2}{\!\!}\operatorname{MSO}$-transduction is the purpose of the next result.
Theorem 4.4.
There exists a non-deterministic $\operatorname{C}_{2}{\!\!}\operatorname{MSO}$-transduction $\tau$ such that, for every weakly-partitive set system $(U,\mathcal{S})$ represented as the $\{\mathsf{SET}\}$-structure $\mathbb{S}$ and inducing the weakly-partitive tree $(T,\lambda,<)$ represented as the $\{\mathsf{ancestor},\mathsf{degenerate},\mathsf{betweeness}\}$-structure $\mathbb{T}$, we have $\mathbb{T}\in\tau(\mathbb{S})$ and every output in $\tau(\mathbb{S})$ is a weakly-partitive tree of $(U,\mathcal{S})$.
Proof 4.5.
By Lemma 4.1, we have a $\operatorname{C}_{2}{\!\!}\operatorname{MSO}$-transduction which on inputs $\mathbb{S}$ outputs the $\{\mathsf{SET},\mathsf{ancestor}\}$-structure $\mathbb{S}\sqcup\mathbb{T^{\prime}}$, where $\mathbb{T^{\prime}}$ is the $\{\mathsf{ancestor}\}$-structure modeling $T$, with $L(T)=U=U_{\mathbb{S}}$. It thus remains to define a second $\operatorname{C}_{2}{\!\!}\operatorname{MSO}$-transduction which outputs $\mathbb{T}$ from $\mathbb{S}\sqcup\mathbb{T^{\prime}}$. Actually, an $\operatorname{MSO}$-interpretation is enough. Indeed, we can define an $\operatorname{MSO}$-formula $\phi_{\mathsf{degenerate}}$ which selects the inner nodes of $T$ that are labeled $\mathsf{degenerate}$ by $\lambda$, and an $\operatorname{MSO}$-formula $\phi_{\mathsf{betweeness}}$ which selects triples $(x,y,z)$ of elements such that $x<y<z$ or $z<y<x$. In order to define both formulas, we use a functional symbol $\mathsf{leafset}(t)$ which is interpreted as the set $L(T_{t})$, and is clearly $\operatorname{MSO}$-definable on $\{\mathsf{ancestor}\}$-structures such as $\mathbb{T}$. Indeed, an element $z$ belongs to $\mathsf{leafset}(t)$ if and only if it is a leaf and has $t$ as ancestor.
According to Theorem 4.3, an inner node $t$ is labeled $\mathsf{degenerate}$ if and only if, for every two children $s_{1},s_{2}$ of $t$ the set $L(T_{s_{1}})\cup L(T_{s_{2}})$ is in $\mathcal{S}$. We can define this in $\operatorname{MSO}$ using the following formula:
$\phi_{\mathsf{degenerate}}(x):=\forall y\forall z\Big{(}\big{(}\mathsf{parent}(x,y)\land\mathsf{parent}(x,z)\big{)}\rightarrow\mathsf{SET}\big{(}\mathsf{leafset}(y)\cup\mathsf{leafset}(z)\big{)}\Big{)}.$
Now, according to Theorem 4.3, we have $x<y<z$ or $z<y<x$ exactly when $x$, $y$, and $z$ are three distinct children of some node $t$ and the three following properties are satisfied:
  1. 1.
    $L(T_{x})\cup L(T_{z})$ is not a member of $\mathcal{S}$;
  2. 2.
    there exists a member of $\mathcal{S}$ which includes both $L(T_{x})$ and $L(T_{y})$ but excludes $L(T_{z})$;
  3. 3.
    there exists a member of $\mathcal{S}$ which includes both $L(T_{y})$ and $L(T_{z})$ but excludes $L(T_{x})$.
Indeed, when these conditions are satisfied, $t$ is necessarily labeled $\mathsf{linear}$ (Item 1 asserts that it is not labeled by $\mathsf{degenerate}$ and, e.g., Item 2 that it is not labeled by $\mathsf{prime}$), and, among the possible orderings of $\{x,y,z\}$, only the ones ensuring $x<y<z$ or $z<y<x$ are possible ($x<z<y$ and $y<z<x$ are not possible because of Item 2, and $y<x<z$ and $z<x<y$ are not possible because of Item 3). Each of these properties is $\operatorname{MSO}$-definable:
$\phi_{\mathsf{betweeness}}(x,y,z):=$ $\exists t\Big{[}\mathsf{parent}(t,x)\land\mathsf{parent}(t,y)\land\mathsf{parent}(t,z)\land$
$\neg\mathsf{SET}\big{(}\mathsf{leafset}(x)\cup\mathsf{leafset}(z)\big{)}\land$
$\exists S\Big{(}\mathsf{SET}(S)\land\big{(}\mathsf{leafset}(x)\cup\mathsf{leafset}(y)\big{)}\subseteq S\land\mathsf{leafset}(z)\cap S=\emptyset\Big{)}\land$
$\exists S\Big{(}\mathsf{SET}(S)\land\big{(}\mathsf{leafset}(z)\cup\mathsf{leafset}(y)\big{)}\subseteq S\land\mathsf{leafset}(x)\cap S=\emptyset\Big{)}\Big{]}.$
Once $\phi_{\mathsf{degenerate}}$ and $\phi_{\mathsf{betweeness}}$ are defined, our transduction drops the original predicate $\mathsf{SET}$ which is no longer needed in the output $\mathbb{T}$.
If $\mathcal{S}$ is partitive then the weakly-partitive tree it induces enjoys a simple form, and is truly unique. Indeed, the label $\mathsf{linear}$ and, thus, the partial order $<$, are not needed.
Theorem 4.6 ([7]).
Let $\mathcal{S}$ be a weakly-partitive family and $(T,\lambda,<)$ be the weakly-partitive tree it induces. If $\mathcal{S}$ is partitive, then $\lambda^{-1}(\mathsf{linear})=\emptyset$ and $<$ is empty.
Hence, in case of a partitive set systems $(U,\mathcal{S})$, we can consider the simpler object $(T,\lambda)$, called the partitive tree induced by $\mathcal{S}$ (or the partitive tree of $(U,\mathcal{S})$) in which $\lambda$ maps $V(T)\setminus L(T)$ to $\{\mathsf{degenerate},\mathsf{prime}\}$. As a direct consequence of Theorems 4.6 and 4.4, we can produce, through a $\operatorname{C}_{2}{\!\!}\operatorname{MSO}$-transduction, the partitive tree induced by a partitive set system and naturally modeled by an $\{\mathsf{ancestor},\mathsf{degenerate}\}$-structure.
Corollary 4.7.
There exists a non-deterministic $\operatorname{C}_{2}{\!\!}\operatorname{MSO}$-transduction $\tau$ such that, for each partitive set system $(U,\mathcal{S})$ represented as the $\{\mathsf{SET}\}$-structure $\mathbb{S}$ and inducing the partitive tree $(T,\lambda)$ represented as the $\{\mathsf{ancestor},\mathsf{degenerate}\}$-structure $\mathbb{T}$, we have $\mathbb{T}\in\tau(\mathbb{S})$ and every output in $\tau(\mathbb{S})$ is a partitive tree of $(U,\mathcal{S})$.
4.2 Application to modular decomposition
Let $G$ be a directed graph and let $M\subseteq V(G)$. We say that $M$ is a module (of $G$) if for every $u\notin M$ and every $v,w\in M$, ${uv\in E(G)\!\iff\!uw\in E(G)}$ and ${vu\in E(G)\!\iff\!wu\in E(G)}$. Clearly, the empty set, $V(G)$, and all the singletons $\{x\}$ for $x\in V(G)$ are modules; they are called the trivial modules of $G$. We say a non-empty module $M$ is maximal if it is not properly contained in any non-trivial module. Let $M$ and $M^{\prime}$ be two disjoint non-empty modules of $G$. Considering the edges that go from $M$ to $M^{\prime}$, namely edges from the set $(M\times M^{\prime})\cap E(G)$, we have two possibilities: either it is empty, or it is equal to $M\times M^{\prime}$. We write $M\text{$\not\to$}M^{\prime}$ in the former case and $M\text{$\to$}M^{\prime}$ in the latter. (It is of course possible to have both $M\text{$\to$}M^{\prime}$ and $M^{\prime}\text{$\to$}M$.) A modular partition of $G$ is a partition $\mathcal{P}=\{M_{1},\ldots,M_{\ell}\}$ of $V(G)$ such that every $M_{i}$ is a non-empty module. A modular partition $\mathcal{P}=\{M_{1},\dots,M_{\ell}\}$ is called maximal if it is non-trivial and every $M_{i}$ is strong and maximal. Note that every graph has exactly one maximal modular partition.
A modular decomposition of $G$ is a rooted tree $T$ in which the leaves are the vertices of $G$, and for each inner node $t\in T$, $t$ has at least two children and the set $L(T_{t})$ is a module of $G$. In a modular decomposition $T$ of $G$, for each inner node $t\in V(T)$ with children $c_{1},\ldots,c_{r}$, the family $\mathcal{P}_{t}=\{L(T_{c_{1}}),\ldots,L(T_{c_{r}})\}$ is a modular partition of $G[V(T_{t})]$. When each such partition is maximal, the decomposition is unique and it is called the maximal modular decomposition of $G$. The maximal modular decomposition $T$ of $G$ alone is not sufficient to characterize $G$. However, enriching $T$ with, for each inner node $t$ with children $c_{1},\ldots,c_{j}$, the information of which pair of modules $\big{(}L(T_{c_{i}}),L(T_{c_{j}})\big{)}$ is such that $L(T_{c_{i}})\text{$\to$}L(T_{c_{j}})$, yields a unique canonical representation of $G$. Formally, the enriched modular decomposition of $G$ is the pair $(T,F)$ where $T$ is the maximal modular decomposition of $G$ (with $L(T)=V(G)$) and $F\subset V(T)\times V(T)$ is a binary relation, that relates a pair $(s,t)$ of nodes of $T$, denoted $st\in F$, exactly when $s$ and $t$ are siblings and $L(T_{s})\text{$\to$}L(T_{t})$. The elements of $F$ are called $\mathsf{m}$-edges.
It should be mentioned that the family of all non-empty modules of $G$ is known to be weakly-partitive (or even partitive when $G$ is undirected). In particular, the maximal modular decomposition $T$ of $G$ is the laminar tree induced by the family of strong modules. Hence Theorem 4.4 could be used to produce a partially-ordered and labeled tree which displays all the modules of $G$. However, this weakly-partitive tree is not sufficient for being able to recover the graph $G$ from it. We now prove how to obtain the enriched modular decomposition of $G$ through a $\operatorname{C}_{2}{\!\!}\operatorname{MSO}$-transduction.
To model enriched modular decompositions as relational structures we use the relational vocabulary $\{\mathsf{ancestor},\textsf{m}\text{-}\mathsf{edge}\}$ where $\mathsf{ancestor}$ and $\textsf{m}\text{-}\mathsf{edge}$ are two binary relation names. An enriched modular decomposition $(T,F)$ of a graph $G$ is modeled by the $\{\mathsf{ancestor},\textsf{m}\text{-}\mathsf{edge}\}$-structure $\mathbb{M}$ with universe $U_{\mathbb{M}}=V(T)$, $\mathsf{ancestor}_{\mathbb{M}}$ being the set of pairs $(s,t)$ for which $s$ is an ancestor of $t$ in $T$, and $\textsf{m}\text{-}\mathsf{edge}_{\mathbb{M}}$ being the set of all pairs $(s,t)$ such that $st\in F$ (in particular, $s$ and $t$ are siblings in $T$). We use Lemma 4.1 in order to transduce the maximal modular decomposition of a graph.
Theorem 4.8.
There exists a non-deterministic $\operatorname{C}_{2}{\!\!}\operatorname{MSO}$-transduction $\tau$ such that for every directed graph $G$ represented as the $\{\mathsf{edge}\}$-structure $\mathbb{G}$, $\tau(\mathbb{G})$ is non-empty and every output in $\tau(\mathbb{G})$ is equal to some $\{\mathsf{ancestor},\textsf{m}\text{-}\mathsf{edge}\}$-structure $\mathbb{M}$ representing the enriched modular decomposition $(T,F)$ of $G$.
Proof 4.9.
Let $G$ be a graph represented by the $\{\mathsf{edge}\}$-structure $\mathbb{G}$. Several objects are associated to $G$, and each of them can be described by a structure:
  • let $\mathcal{M}$ be the family of non-empty modules of $G$ and let $\mathbb{S}$ be the $\{\mathsf{SET}\}$-structure modeling the weakly-partitive set system $(V(G),\mathcal{M})$ with $U_{\mathbb{S}}=V(G)$;
  • let $T$ be the laminar tree induced by the weakly-partitive family $\mathcal{M}$ and let $\mathbb{T}$ be the $\{\mathsf{ancestor}\}$-structure modeling it with $U_{\mathbb{T}}=V(T)$ and $L(T)=V(G)\subset U_{\mathbb{T}}$;
  • let $F$ be the $\mathsf{m}$-edge relation, namely the subset of $V(T)\times V(T)$ such that $(T,F)$ is the maximal modular decomposition of $G$, and let $\mathbb{M}$ be the $\{\mathsf{ancestor},\textsf{m}\text{-}\mathsf{edge}\}$-structure modeling it with $\mathbb{T}\sqsubset\mathbb{M}$ and $U_{\mathbb{T}}=U_{\mathbb{M}}$.
Our $\operatorname{C}_{2}{\!\!}\operatorname{MSO}$-transduction is obtained by composing the three following transductions:
  • $\tau_{1}$: an $\operatorname{MSO}$-interpretation which outputs the $\{\mathsf{edge},\mathsf{SET}\}$-structure $\mathbb{G}\sqcup\mathbb{S}$ from $\mathbb{G}$;
  • $\tau_{2}$: the non-deterministic $\operatorname{C}_{2}{\!\!}\operatorname{MSO}$-transduction given by Lemma 4.1 which produces the $\{\mathsf{edge},\mathsf{SET},\mathsf{ancestor}\}$-structure $\mathbb{G}\sqcup\mathbb{S}\sqcup\mathbb{T}$ from $\mathbb{G}\sqcup\mathbb{S}$;
  • $\tau_{3}$: an $\operatorname{MSO}$-interpretation which outputs the $\{\mathsf{ancestor},\textsf{m}\text{-}\mathsf{edge}\}$-structure $\mathbb{M}$ from $\mathbb{G}\sqcup\mathbb{S}\sqcup\mathbb{T}$.
In order to define $\tau_{1}$ it is sufficient to observe that there exists an $\operatorname{MSO}$-formula $\phi_{\mathsf{SET}}(Z)$ with one monadic free-variable, which is satisfied exactly when $Z$ is a non-empty module of $G$. Then, since $\tau_{2}$ is given by Lemma 4.1, it only remains to define $\tau_{3}$. Given an inner node $t$, we can select, within $\operatorname{MSO}$, the set $L(T_{t})$ of leaves of the subtree rooted in $t$. We thus assume a function $\mathsf{leafset}$, with one first-order free-variable which returns the set of leaves of the subtree rooted at the given node. Equipped with this function, we can define the $\operatorname{MSO}$-formula $\phi_{\textsf{m}\text{-}\mathsf{edge}}$ which selects pairs $(s,r)$ of siblings such that $sr\in F$. Remember that this happen exactly when there exist $u\in L(T_{s})$ and $v\in L(T_{r})$ such that $uv\in E(G)$. Hence, $\phi_{\textsf{m}\text{-}\mathsf{edge}}$ could be defined as:
$\phi_{\textsf{m}\text{-}\mathsf{edge}}(s,r):=s\neq r\>\land\>$ $\exists t\>\big{(}\mathsf{parent}(t,s)\land\mathsf{parent}(t,r)\big{)}\land$
$\exists x\exists y\big{(}x\in\mathsf{leafset}(s)\land y\in\mathsf{leafset}(r)\land\mathsf{edge}(x,y)\big{)}.$
Once defined, the $\operatorname{MSO}$-interpretation $\tau_{3}$ simply drops all non-necessary relations and predicates (namely $\mathsf{edge}$ and $\mathsf{SET}$) and keeps only the $\mathsf{ancestor}$ and $\textsf{m}\text{-}\mathsf{edge}$ relations.
Notice that it is routine to design a deterministic $\operatorname{MSO}$-transduction which, given an $\{\mathsf{ancestor},\textsf{m}\text{-}\mathsf{edge}\}$-structure $\mathbb{M}$ representing an enriched modular decomposition of some directed graph $G$, produces the $\{\mathsf{edge}\}$-structure $\mathbb{G}$ representing $G$.
Cographs
Let $G$ be a graph, $(T,F)$ be its modular decomposition, and $(T,\lambda,<)$ be the weakly-partitive tree induced by the (weakly-partitive) family of its modules. Let $t$ be an inner node of $T$, let $\mathcal{C}$ be its set of children, and let $C$ be the graph $\big{(}V(T)\setminus L(T),F\big{)}\left[\mathcal{C}\right]$ induced by $F$ on the set of children of $t$. It can be checked that, if $\lambda(t)=\mathsf{degenerate}$ then $C$ is either a clique or an independant, and if $\lambda(t)=\mathsf{linear}$ then $C$ is a tournament consistent with $<_{t}$ (i.e., for every $x,y\in\mathcal{C}$, $xy$ is an edge of $C$ if and only if $x<_{t}y$) or with the inverse of $<_{t}$. In the former case, we can refine the $\mathsf{degenerate}$ label into $\mathsf{series}$ and $\mathsf{parallel}$ labels, thus expressing that $C$ is a clique or an indenpendant, respectively. In the latter case, up-to reversing $<_{t}$, we can ensure that $C$ is a tournament consistent with $<_{t}$. This yields a refined weakly-partitive tree $(T,\gamma,<)$, where $\gamma$ maps inner nodes to $\{\mathsf{series},\mathsf{parallel},\mathsf{prime},\mathsf{linear}\}$ and $<$ is the order $\bigcup_{t\in\gamma^{-1}(\mathsf{linear})}<_{t}$ which ensures that tournaments are consistent with the corresponding $<_{t}$. Notice that this labeled and partially-ordered tree is now uniquely determined from $G$. Moreover, edges from $F$ that connect children of a node not labeled by $\mathsf{prime}$ can be recovered from the so-refined weakly-partitive tree. In particular, if no nodes of $T$ is labeled by $\mathsf{prime}$, $(T,F)$ and thus $G$ is fully characterized by $(T,\gamma,<)$. Graphs for which this property holds are known as directed cographs, and can be described by the refined weakly-partitive tree, called cotree, explained above and formalized in the following statement.
Theorem 4.10.
Let $G$ be a directed cograph and let $T$ be the laminar tree induced by the family of its strong modules. There exists a unique total labeling $\lambda$ from the set $V(T)\setminus L(T)$ of inner nodes of $T$ to the set $\{\mathsf{series},\mathsf{parallel},\mathsf{linear}\}$ of labels, and, for each inner node $t\in\lambda^{-1}(\mathsf{linear})$, a unique linear ordering $<_{t}$ of its children, such that every inner node having exactly two children is labeled $\mathsf{series}$ or $\mathsf{parallel}$, and the following condition is satisfied:
  • for every two leaves $x$ and $y$ of $T$, denoting by $t$ their least common ancestor, $xy$ is an edge of $G$ if and only if either $t$ is labeled by $\mathsf{linear}$ and $x<_{t}y$, or $t$ is labeled by $\mathsf{series}$.
We naturally model cotrees as $\{\mathsf{ancestor},\mathsf{series},\varepsilon\text{-}\mathsf{ord}\}$-structures as follow. A cotree $(T,\gamma,<)$ is modeled by $\mathbb{C}$ where $U_{\mathbb{C}}=V(T)$, $\mathsf{series}_{\mathbb{C}}=\gamma^{-1}(\mathsf{series})$, and $\varepsilon\text{-}\mathsf{ord}_{\mathbb{C}}=\{(x,y)\mid x<y\}$. The nodes that are labeled by $\mathsf{linear}$ could be recovered as those inner nodes whose children are related by $\varepsilon\text{-}\mathsf{ord}$, while the nodes that are labeled by $\mathsf{parallel}$ could be recovered as those inner nodes which are labeled neither by $\mathsf{series}$ nor by $\mathsf{linear}$. Based on Theorem 4.10 and as a consequence of Theorem 4.8, we can design a $\operatorname{C}_{2}{\!\!}\operatorname{MSO}$-transduction which produces the cotree of a cograph $G$ from $G$.
Corollary 4.11.
There exists a non-deterministic $\operatorname{C}_{2}{\!\!}\operatorname{MSO}$-transduction $\tau$ such that, for each cograph $G$ modeled by the $\{\mathsf{edge}\}$-structure $\mathbb{G}$, $\tau(\mathbb{G})$ is non-empty and every output in $\tau(\mathbb{G})$ is equal to some $\{\mathsf{ancestor},\mathsf{series},\varepsilon\text{-}\mathsf{ord}\}$-structure $\mathbb{C}$ representing the cotree $(T,\gamma,<)$ of $G$.
5 Transducing split and bi-join decompositions
In this section we consider systems of bipartitions. In graph theory, two known systems of bipartitions are splits and bi-joins. Both are instances of so called weakly-bipartitive systems of bipartitions (or partitive systems of bipartitions in case of undirected graphs). In this section, we first give a $\operatorname{C}_{2}{\!\!}\operatorname{MSO}$-transduction of the canonical tree-like decomposition of weakly-bipartitive systems of bipartitions and then derive a $\operatorname{C}_{2}{\!\!}\operatorname{MSO}$-transductions for split and bi-join decompositions of a graph.
5.1 Transducing weakly-bipartitive trees
A bipartition system is a pair $(U,\mathcal{B})$ consisting of a finite set $U$, the universe, and a family $\mathcal{B}$ of bipartitions of $U$ such that $\{\emptyset,U\}\notin\mathcal{B}$ and $\{\{a\},U\setminus\{a\}\}\in\mathcal{B}$ for all $a\in U$.
To model bipartition systems, we use the extended vocabulary $\{\mathsf{BIPART}\}$ where $\mathsf{BIPART}$ is a set predicate name of arity $2$. A bipartition system $(U,\mathcal{B})$ is thus naturally modeled as the $\{\mathsf{BIPART}\}$-structure $\mathbb{B}$ with universe $U_{\mathbb{B}}=U$ and interpretation $\mathsf{BIPART}_{\mathbb{B}}=\mathcal{B}$.
Two bipartitions $\{X,Y\}$ and $\{X^{\prime},Y^{\prime}\}$ of a set $U$ overlap if the four sets $X\cap X^{\prime}$, $X\cap Y^{\prime}$, $Y\cap X^{\prime}$ and $Y\cap Y^{\prime}$ are non-empty. A bipartition family is said to be laminar (aka overlap-free) if no two bipartitions in $\mathcal{B}$ overlap. By extension, we call a family $\mathcal{B}$ of bipartition of a set $U$ laminar whenever $(U,\mathcal{B})$ is a bipartition system which is laminar. Similarly to laminar set systems, we can associate a tree with each laminar bipartition system. The tree $T$, called the laminar tree induced by $(U,\mathcal{B})$ (or laminar tree of $(U,\mathcal{B})$), associated with a laminar bipartition system $(U,\mathcal{B})$ is unrooted, each element of $U$ corresponds to a leaf of $T$ and for each bipartition $\{X,Y\}$ in $\mathcal{B}$ there is exactly one edge $e$ such that $X=L(T_{1})$ and $Y=L(T_{2})$ for the two connected components $T_{1},T_{2}$ of $T-e$. We remark that there is a unique such tree and each inner node has degree at least three. Furthermore, the size of the laminar tree is linearly bounded in the size of the universe $U$.
A bipartition system $(U,\mathcal{B})$ is said to be weakly-bipartitive if for every two overlapping bipartitions $\{X,Y\},\{X^{\prime},Y^{\prime}\}\in\mathcal{B}$, the biparitions $\{X\cup X^{\prime},Y\cap Y^{\prime}\}$, $\{X\cup Y^{\prime},Y\cap X^{\prime}\}$, $\{Y\cup X^{\prime},X\cap Y^{\prime}\}$ and $\{Y\cup Y^{\prime},X\cap X^{\prime}\}$ are in $\mathcal{B}$. It is bipartitive, if additionally for every two overlapping bipartitions $\{X,Y\},\{X^{\prime},Y^{\prime}\}\in\mathcal{B}$, the bipartition $\{X\operatorname*{\bigtriangleup}X^{\prime},Y\operatorname*{\bigtriangleup}Y^{\prime}\}$ is also in $\mathcal{B}$. By extension, we call a family $\mathcal{B}$ of bipartitions of a set $U$ weakly-bipartitive or bipartitive whenever $(U,\mathcal{B})$ is a bipartition system which is weakly-bipartitive or partitive, respectively.
A bipartition $\{X,Y\}$ of a bipartition system $(U,\mathcal{B})$ is said to be strong if $\{X,Y\}$ does not overlap with any other bipartition in $\mathcal{B}$. We denote the subfamily of strong bipartitions in $\mathcal{B}$ by $\mathcal{B}_{!}$ and remark that $\mathcal{B}_{!}$ is a laminar family. Hence, it induces a laminar tree $T$. By extension, we say that $T$ is induced by the bipartition system $(U,\mathcal{B})$ (or simply by $\mathcal{B}$). As the laminar tree $T$ of a bipartition system $(U,\mathcal{B})$ is undirected, we model $T$ by the $\{\mathsf{t}\text{-}\mathsf{edge}\}$-structure $\mathbb{T}$ with universe $U_{\mathbb{T}}=V(T)$ and binary relation $\mathsf{t}\text{-}\mathsf{edge}_{\mathbb{T}}$ modeling the edge relation of $T$.
The following lemma gives a connection between laminar trees of set systems and laminar trees of bipartition systems. Note that a very similar statement follows from [25, Lemma 1.14] and hence we only provide a proof for sake of completeness in the appendix.
Lemma 5.1.
Let $(U,\mathcal{B})$ be a bipartition system and $a\in U$. Then $(U,\mathcal{S}_{a})$ is a set system where $\mathcal{S}_{a}:=\{X:a\notin X\text{ and }\{X,U\setminus X\}\in\mathcal{B}\}\cup\{\{a\},U\}$. Moreover, the laminar tree of $(U,\mathcal{B})$ can be obtained from the laminar tree of $(U,\mathcal{S}_{a})$ by deleting the root, making node $\{a\}$ adjacent to the node $U\setminus\{a\}$ and making the tree undirected.
Lemma 5.2.
Let $\Sigma$ be an extended vocabulary, including a set unary predicate name $\mathsf{BIPART}$ and not including the binary relational symbol $\mathsf{t}\text{-}\mathsf{edge}$. There exists a non-deterministic $\operatorname{C}_{2}{\!\!}\operatorname{MSO}$-transduction $\tau$ such that, for each bipartition system $(U,\mathcal{B})$ represented as the $\{\mathsf{BIPART}\}$-structure $\mathbb{B}$ and inducing the laminar tree $T$ with ${L(T)=U}$, and for each $\Sigma$-structure $\mathbb{A}$ with $\mathbb{B}\sqsubseteq\mathbb{A}$, $\tau(\mathbb{A})$ is non-empty and every output in $\tau(\mathbb{A})$ is equal to some $\{\mathsf{t}\text{-}\mathsf{edge}\}$-structure $\mathbb{T}$ representing $T$.
Proof 5.3.
Let $\mathcal{B}$ be a family of bipartitions of $U$ represented as the extended relational structure $\mathbb{B}$ and $\mathbb{A}$ a $\Sigma$-structure with $\mathbb{B}\sqsubseteq\mathbb{A}$. We first use the following two simple atomic $\operatorname{C}_{2}{\!\!}\operatorname{MSO}$-transduction to obtain a set system from $(U,\mathcal{B})$. Guess any colouring $A$ for which there is a single $a\in U$ with $A(a)$ being satisfied. Interpret using the following formula to obtain a $\mathsf{SET}$-predicate:
$\phi_{\mathsf{SET}}(X):=\forall a(A(a)\rightarrow(X=\{a\}\lor X=U\lor a\notin X))\land\mathsf{BIPART}(X,U\setminus X).$
The resulting set system $(U,\mathcal{S}_{a})$ represented by the $\{\mathsf{SET}\}$-structure $\mathbb{S}$ is a set system by Lemma 5.1. Hence, we can use the transduction from Lemma 4.1 to obtain the $\Sigma\cup\{\mathsf{ancestor}\}$-structure $\mathcal{A}\sqcup\mathcal{T}$ where $\mathcal{T}$ is the $\{\mathsf{ancestor}\}$-structure representing the laminar tree $T$ of $(U,\mathcal{S}_{a})$. We now use the following two atomic transductions to obtain the weakly-bipartitive tree of $(U,\mathcal{B})$ according to Lemma 5.1. We fist restrict the universe to all nodes which are not the root of $\mathbb{T}$. We then interpret using the following formula to obtain the $\mathsf{t}\text{-}\mathsf{edge}$-predicate which makes the graph undirected and adds an edge from $a$ to the node representing the set $U\setminus\{a\}$ which is the root of $T$ after removing the node corresponding to $U$:
$\phi_{\mathsf{t}}\text{-}\mathsf{edge}(x,y):=\mathsf{parent}(x,y)\lor\mathsf{parent}(y,x)\lor(x=a\land\lnot\exists z\mathsf{parent}(z,y))\lor(y=a\land\lnot\exists z\mathsf{parent}(z,x)).$
Finally, the transduction keeps only $\mathsf{t}\text{-}\mathsf{edge}$ and any $Q\in\Sigma$ as relations for its output.
Similar to the set system case, for each weakly-bipartitive family $\mathcal{B}$ we can equip the laminar tree with a labelling of its inner nodes and a partial order which characterizes the family $\mathcal{B}$.
Theorem 5.4 (Theorem 3 in [17]).
Let $\mathcal{B}$ be a weakly-bipartitive family, $\mathcal{B}_{!}$ be its subfamily of strong bipartitions, and $T$ the laminar tree it induces. There exists a total labeling function $\lambda$ from the set $V(T)\setminus L(T)$ of the inner nodes of $T$ to the set $\{\mathsf{degenerate},\mathsf{prime},\mathsf{linear}\}$, and, for each inner node $t\in\lambda^{-1}(\mathsf{linear})$, a linear ordering $<_{t}$ of its neighbors, such that every inner node of degree exactly three is labeled by $\mathsf{degenerate}$ and the following conditions are satisfied:
  • for each bipartition $\{X,Y\}\in\mathcal{B}\setminus\mathcal{B}_{!}$, there exists $t\in V(T)\setminus L(T)$ and a subset $\mathcal{C}$ of neighbors of $t$ such that either $X$ or $Y$ is equal to the set $\bigcup_{c\in\mathcal{C}}L(T^{t}_{c})$ and either $\lambda(t)=\mathsf{linear}$ and $\mathcal{C}$ is a $<_{t}$-interval, or $\lambda(t)=\mathsf{degenerate}$;
  • conversely, for each inner node $t$ and each non-empty subset $\mathcal{C}$, $\mathcal{C}\not=U$ of neighbors of $t$, the bipartition $\big{\{}\bigcup_{c\in\mathcal{C}}L(T^{t}_{c}),\bigcup_{c\notin\mathcal{C}}L(T^{t}_{c})\big{\}}$ is a member of $\mathcal{B}$ if either $\lambda(t)=\mathsf{linear}$ and $\mathcal{C}$ is $<_{t}$-interval or $\lambda(t)=\mathsf{degenerate}$.
The tree $T$ and its labelling function $\lambda$ are uniquely determined by $\mathcal{B}$. We note that the total orders $<_{t}$ are uniquely determined by $\mathcal{B}$ up to inverting and cyclic shifting, i.e., for a linear order $<$ of $X=\{x_{1},\dots,x_{\ell}\}$ with $x_{1}<\dots<x_{\ell}$ we call the order $<^{\prime}$ of $X$ with $x_{\ell}<^{\prime}x_{1}<^{\prime}\dots<^{\prime}x_{\ell-1}$ a cyclic shift of $<$. To see this, note that in the theorem above the bipartitions obtained from nodes with label $\mathsf{linear}$ can be equivalently obtained from a $<_{t}$-interval or the complement of a $<_{t}$-interval. Indeed, a cyclic shifting of a linear order $<$ maintains the property of being either a $<$-interval or the complement of a $<$-interval. Neglecting that $<_{t}$ are only unique up to inverting and cyclic shifting, we represent every weakly-bipartitive family $\mathcal{B}$ by a triple $(T,\lambda,(<_{t})_{t\in\lambda^{-1}(\mathsf{linear})})$ where $T$ is the laminar tree induced by $\mathcal{B}_{!}$, $\lambda:V(T)\setminus L(T)\to\{\mathsf{degenerate},\mathsf{prime},\mathsf{linear}\}$ is the labeling function, and $<_{t}$ is the total order of the neighbours of $t$ described above. Note that in this case $\bigcup_{t\in\lambda^{-1}(\mathsf{linear})}<_{t}$ is not a partial order as it is not necessarily transitive. We call the triple $(T,\lambda,(<_{t})_{t\in\lambda^{-1}(\mathsf{linear})})$ from Theorem 5.4 the bipartitive tree of $\mathcal{B}$.
We model weakly-bipartitive trees $(T,\lambda,(<_{t})_{t\in\lambda^{-1}(\mathsf{linear})})$ of a weakly-bipartitive family $\mathcal{B}$ of bipartition of $U$ by the $\{\mathsf{t}\text{-}\mathsf{edge},\mathsf{degenerate},\mathsf{betweeness}\}$-structure $\mathbb{T}$ of universe $U_{\mathbb{T}}=V(T)$ such that $\langle V(T),\mathsf{t}\text{-}\mathsf{edge}_{\mathbb{T}}\rangle\sqsubseteq\mathbb{T}$ models $T$ with ${L(T)=U}$, $\mathsf{degenerate}_{\mathbb{T}}$ is a unary relation which selects all inner nodes of $T$ of label $\mathsf{degenerate}$, i.e., ${\mathsf{degenerate}_{\mathbb{T}}=\lambda^{-1}(\mathsf{degenerate})}$, and $\mathsf{betweeness}_{\mathbb{T}}$ is a relation of arity four selecting all quadruples $(t,x,y,z)$ for which $y$ is between $x$ and $z$ in $<_{t}$. Note that in order to define $\mathsf{betweeness}$ we essentially choose one representative order amongst all orders obtained by inverting and cyclic shifting. Similarly to the weakly-partitive case, we can recover the inner nodes of $T$ that are labeled by $\mathsf{linear}$ through an $\operatorname{MSO}$-formula as those inner nodes whose neighbors are related by $\mathsf{betweeness}$. Then $\mathsf{prime}$ nodes are the ones which are labeled neither by $\mathsf{degenerate}$ nor by $\mathsf{linear}$ which can be obtained by an $\operatorname{MSO}$-transduction. Furthermore, Theorem 5.4, gives rise to an $\operatorname{MSO}$-transduction which takes as input a weakly-bipartitive tree and outputs the weakly-bipartitive set system which induced it.
We can now extend the transduction from Lemma 5.2 to obtain a $\operatorname{C}_{2}{\!\!}\operatorname{MSO}$-transduction producing the weakly-bipartitive tree of a weakly-bipartitive family.
Theorem 5.5.
There exists a non-deterministic $\operatorname{C}_{2}{\!\!}\operatorname{MSO}$-transduction $\tau$ such that, for every weakly-bipartitive bipartition system $(U,\mathcal{B})$ represented as the $\{\mathsf{BIPART}\}$-structure $\mathbb{B}$, $\tau(\mathbb{B})$ is non-empty and every output in $\tau(\mathbb{B})$ is equal to some $\{\mathsf{t}\text{-}\mathsf{edge},\mathsf{degenerate},\mathsf{betweeness}\}$-structure $\mathbb{T}$ representing the weakly-bipartitive tree $\left(T,\lambda,(<_{t})_{t\in\lambda^{-1}(\mathsf{linear})}\right)$ of $(U,\mathcal{B})$.
Proof 5.6.
Lemma 5.2 already provides a $\operatorname{C}_{2}{\!\!}\operatorname{MSO}$-transduction which on input $\mathbb{B}$ outputs the $\{\mathsf{BIPART},\mathsf{t}\text{-}\mathsf{edge}\}$-structure $\mathbb{B}\sqcup\mathbb{T}^{\prime}$, where $\mathbb{T}^{\prime}$ represents the $\{\mathsf{t}\text{-}\mathsf{edge}\}$-structure modeling the laminar tree $T$ of $(U,\mathcal{B})$. To complete the proof of the theorem, we provide an $\operatorname{MSO}$-transduction which outputs $\mathbb{T}$ given input $\mathbb{B}\sqcup\mathbb{T}^{\prime}$. For the transduction we define an $\operatorname{MSO}$-formula $\phi_{\mathsf{degenerate}}$ which selects the inner nodes of $T$ which are labeled $\mathsf{degenerate}$, and a formula $\phi_{\mathsf{betweeness}}$ which selects quadruples of nodes $(t,x,y,z)$ such that either $x<_{t}y<_{t}z$ or $z<_{t}y<_{t}x$. Recall, that the ordering $<_{t}$ for each inner node $t$ of the neighbours of $t$ are only unique up to inverting the order and cyclic shift. In order to select a representative of $<_{t}$ which is unique only up to inverting the order, we need non-determinism (details follow). We use a functional symbol $\mathsf{leafset}(t,s)$ which is interpreted as the set $L(T_{s}^{t})$. The functional predicate $\mathsf{leafset}$ clearly can be defined in $\operatorname{MSO}$ as a element $z$ belongs to $\mathsf{leafset}(t,s)$ if and only if $z$ is a leaf and the path from $z$ to $s$ does not contain $t$.
Observe that by Theorem 5.4 an inner node $t$ is labeled $\mathsf{degenerate}$ if and only if for every pair of neighbours $s_{1},s_{2}$ of $t$ the bipartition $\big{\{}L(T^{t}_{s_{1}})\cup L(T^{t}_{s_{2}}),U\setminus\big{(}L(T^{t}_{s_{1}})\cup L(T^{t}_{s_{2}})\big{)}\big{\}}$ is in $\mathcal{B}$. We can define this in $\operatorname{MSO}$ using the following formula:
$\phi_{\mathsf{degenerate}}$ $(x):=\forall y\forall z\Big{(}\big{(}\mathsf{t}\text{-}\mathsf{edge}(x,y)\land\mathsf{t}\text{-}\mathsf{edge}(x,z)\big{)}\rightarrow$
$\mathsf{BIPART}\big{(}\mathsf{leafset}(x,y)\cup\mathsf{leafset}(x,z),U\setminus(\mathsf{leafset}(x,y)\cup\mathsf{leafset}(x,z))\big{)}\Big{)}.$
By Theorem 5.4, two neighbors $x$ and $y$ of $t$ are consecutive in $<_{t}$ if and only if, the bipartition $\big{\{}L(T^{t}_{x})\cup L(T^{t}_{y}),U\setminus\big{(}L(T^{t}_{x})\cup L(T^{t}_{y})\big{)}\big{\}}$ is a member of $\mathcal{B}$ and at least one of $x$ or $y$ is not extremal in $<_{t}$. Hence, fixing, for each $\mathsf{linear}$-labeled node $t$, the two extremals of $<_{t}$ is enough to define $<_{t}$ on $\mathbb{B}\sqcup\mathbb{T^{\prime}}$ in $\operatorname{MSO}$. To this end, we consider a rooted version of $T$, in which the non-deterministically-guessed root $r$ is not labeled by $\mathsf{linear}$, and we fix the parent $p$ of each $\mathsf{linear}$-labeled node $t$ as the minimum of $<_{t}$. We also guess a set $M$ of nodes, such that in $T$ rooted in $r$, exactly one child $m$ of each $\mathsf{linear}$-labeled node $t$ with parent $p$ belongs to $M$, and furthermore, the pair $\big{\{}L(T^{t}_{p})\cup L(T^{t}_{m}),U\setminus\big{(}L(T^{t}_{p})\cup L(T^{t}_{m})\big{)}\big{\}}$ is a member of $\mathcal{B}$. Thus, $m$ can be considered as the maximum of $<_{t}$. Equipped with the relation $\mathsf{ancestor}$ (induced by the rooting of $T$ in $r$) and the colouring $M$, we are now able to define $\phi_{\mathsf{betweeness}}$ in $\operatorname{MSO}$. Indeed, for two neighbours $x,y$ of a node $t$ labeled $\mathsf{linear}$, with parent $p$, and child $m\in M$, it holds that $x<_{t}y$ if and only if there is a set $X$ containing both $L(T_{p}^{t})$ and $L(T_{x}^{t})$, excluding $L(T_{y}^{t})$ and $L(T_{m}^{t})$, and satisfying $(X,U\setminus X)\in\mathcal{B}$. Hence, we can define in $\operatorname{MSO}$, $x<_{t}y$ as follows:
$x<_{t}y:=\exists p\exists m\Big{[}\mathsf{t}\text{-}\mathsf{edge}(t,x)\land\mathsf{t}\text{-}\mathsf{edge}(t,y)\land\mathsf{parent}(p,t)\land\mathsf{t}\text{-}\mathsf{edge}(t,m)\land M(m)\land\exists X\exists Y$
$\Big{(}\mathsf{BIPART}(X,Y)\land\big{(}\mathsf{leafset}(t,p)\cup\mathsf{leafset}(t,x)\big{)}\subseteq X\land\big{(}\mathsf{leafset}(t,y)\cup\mathsf{leafset}(t,m)\big{)}\subseteq Y\Big{)}\Big{]}.$
Note that any node with a child in $M$ must necessarily be labeled $\mathsf{linear}$ and hence:
$\phi_{\mathsf{betweeness}}(t,x,y,z)$ $:=(x<_{t}y<_{t}z)\lor(z<_{t}y<_{t}x).$
Finally, the transduction forgets the predicate $\mathsf{BIPART}$ to produce the output $\mathbb{T}$.
Indeed, considering a bipartitive family $\mathcal{B}$ makes the use of nodes labeled $\mathsf{linear}$ in the associated tree obsolete, as the following theorem states.
Theorem 5.7 (Theorem 4 in[17]).
Let $\mathcal{B}$ be a weakly-biaprtitive family and $(T,\lambda,(<_{t})_{t\in\lambda^{-1}(\mathsf{linear})})$ be the weakly-bipartitive tree it induces. If $\mathcal{B}$ is bipartitive, then $\lambda^{-1}(\mathsf{linear})=\emptyset$ and all $<_{t}$ are empty.
As a consequence, we can capture any bipartitive family $\mathcal{B}$ by the simpler structure $(T,\lambda)$, called the bipartitive tree induced by $\mathcal{B}$ (or or the bipartitive tree of $\mathcal{B}$), where $\lambda$ is now a labeling function labeling each inner nodes of $T$ by either $\mathsf{degenerate}$ or $\mathsf{prime}$. Hence, we obtain the following as a corollary from Theorem 5.5, where naturally the bipartitive tree $(T,\lambda)$ of a bipartitive family $\mathcal{B}$ is modeled by a $\{\mathsf{t}\text{-}\mathsf{edge},\mathsf{degenerate}\}$-structure $\mathbb{B}$.
Corollary 5.8.
There exists a non-deterministic $\operatorname{C}_{2}{\!\!}\operatorname{MSO}$-transduction $\tau$ such that, for each bipartitive set system $(U,\mathcal{B})$ represented as the $\{\mathsf{BIPART}\}$-structure $\mathbb{B}$, $\tau(\mathbb{B})$ is non-empty and every output in $\tau(\mathbb{B})$ is equal to some $\{\mathsf{t}\text{-}\mathsf{edge},\mathsf{degenerate}\}$-structure $\mathbb{T}$ representing the bipartitive tree $(T,\mathsf{degenerate})$ of $(U,\mathcal{B})$.
5.2 Application to split decomposition
Let $G$ be a directed graph. A split in $G$ is a bipartition $\{X,Y\}$ of $V(G)$ such that there are subsets $X^{\operatorname{in}},X^{\operatorname{out}}\subseteq X$, $Y^{\operatorname{in}},Y^{\operatorname{out}}\subseteq Y$ such that for all pairs of vertices $x\in X$ and $y\in Y$ the edge $xy\in E(G)$ if and only if either $x\in X^{\operatorname{out}}$ and $y\in Y^{\operatorname{in}}$ or $x\in Y^{\operatorname{out}}$ and $y\in X^{\operatorname{in}}$. Note that if $G$ is connected then $X^{\operatorname{in}}\cup X^{\operatorname{out}}\not=\emptyset$ and $Y^{\operatorname{in}}\cup Y^{\operatorname{out}}\not=\emptyset$. We call splits $\{X,Y\}$ for which either $|X|=1$ or $|Y|=1$ trivial splits. In a family $\mathcal{B}$ of splits we use the notion of strong splits to coincide with the notion of strong bipartitions.
With a split $\{X,Y\}$ in a connected graph $G$ we associate the following decomposition of $G$ into graphs $G_{X}$ and $G_{Y}$. Let $v_{X}^{Y},v_{Y}^{X}\notin V(G)$ be two new vertices, called marker vertices. Define $G_{X}$ to be the graph obtained by adding $v_{X}^{Y}$ to $G[X]$ and edge $xv_{X}^{Y}$ for every $x\in X^{\operatorname{out}}$ and edge $v_{X}^{Y}x$ for every $x\in X^{\operatorname{in}}$. The graph $G_{Y}$ is defined equivalently. Note that this decomposition captures the graph $G$ as $G$ can be obtained as the union of $G_{X}[X]$ and $G_{Y}[Y]$ with additional edges $xy$ whenever either $xv_{X}^{Y}\in V(G_{X})$ and $v_{Y}^{X}y\in V(G_{Y})$ or $xv_{Y}^{X}\in V(G_{Y})$ and $v_{X}^{Y}y\in V(G_{X})$. We can recursively decompose $G$ along splits to obtain a decomposition into graphs $G_{1},\dots,G_{k}$. We call such a decomposition of $G$ into graphs $G_{1},\dots,G_{k}$ a split decomposition if every graph $G_{i}$ is either a singleton consisting of an original (not a marker) vertex or does not have any non-trivial split. We call the graphs $G_{i}$ components of the split decomposition. To be able to reconstruct the graph, we represent a split decomposition of $G$ into graphs $G_{1},\dots,G_{k}$ by an enriched split decomposition which is a pair $(H,F)$ consisting of a directed graph $H$ with an additional (symmetric) edge relation $F$. The graph $H$ consisting of the disjoint union $G_{1},\dots,G_{k}$. The edge relation $F$ contains an (undirected) edge between every pair of markers corresponding to one step of the decomposition. For a clear distinction, we call the edges in $H$ $\mathsf{c}$-edges and the edges in $F$ $\mathsf{t}$-edges. Note that the splits used in the decomposition are in one-to-one correspondence with the $\mathsf{t}$-edges.
We require the following theorem for our transduction.
Theorem 5.9 ([16]).
If $G$ is connected then the family of splits of $G$ is weakly-bipartitive.
Hence, there is a unique split decomposition of $G$, which is called the canonical split decomposition of $G$, in which we only decompose along strong splits. Note that the tree we obtain by contracting all edges of the enriched canonical split decomposition $(H,F)$ of $G$ (only leaving the t-edges) is isomorphic to the bipartitive tree of the family $\mathcal{B}$.
The following (standard) technical lemma is needed both to show that a graph can be recovered from an enriched split decomposition (by an $\operatorname{MSO}$ transduction) and to define the set of edges in the enriched canonical split decomposition $(H,F)$ when transducing $(H,F)$ from $G$.
Lemma 5.10 (See e.g. Lemma 2.10 in [1]).
Let $G$ be a connected, directed graph and $(H,F)$ an enriched split decomposition of $G$. Then the following hold:
  • $uv\in E(G)$ if and only if there exists a directed path from $u$ to $v$ in $(H,F)$ on which c-edges and t-edges alternate;
  • $v_{X}^{Y}v_{X^{\prime}}^{Y^{\prime}}$ is a c-edge in $(H,F)$ for vertices $v_{X}^{Y},v_{X^{\prime}}^{Y^{\prime}}$ in the same component of $H$ if and only if there is $uu^{\prime}\in E(G)$ such that $u\in Y$ and $u^{\prime}\in Y^{\prime}$.
To model enriched split decompositions as relational structures we use the relational vocabulary $\{\mathsf{t}\text{-}\mathsf{edge},\mathsf{c}\text{-}\mathsf{edge}\}$ where $\mathsf{t}\text{-}\mathsf{edge}$ and $\mathsf{c}\text{-}\mathsf{edge}$ are binary relational symbols. Am enriched split decomposition $(H,F)$ of a graph $G$ is the $\{\mathsf{t}\text{-}\mathsf{edge},\mathsf{c}\text{-}\mathsf{edge}\}$-relation $\mathbb{H}$ with universe $U_{\mathbb{H}}=V(H)$, $\mathsf{t}\text{-}\mathsf{edge}_{\mathbb{H}}$ the set of t-edges in $(H,F)$ and $\mathsf{c}\text{-}\mathsf{edge}_{\mathbb{H}}$ the set of c-edges of $(H,F)$.
We use Lemma 5.2 to transduce the canonical split decomposition of a graph.
Theorem 5.11.
There exists a non-deterministic $\operatorname{C}_{2}{\!\!}\operatorname{MSO}$-transduction $\tau$ such that for any connected, directed graph $G$ represented as the $\{\mathsf{edge}\}$-structure $\mathbb{G}$; $\tau(\mathbb{G})$ is non-empty and every output in $\tau(\mathbb{G})$ is equal to some $\{\mathsf{t}\text{-}\mathsf{edge},\mathsf{c}\text{-}\mathsf{edge}\}$-structure $\mathbb{H}$ representing the canonical split decomposition of $G$.
Proof 5.12.
Let $G$ be a connected, directed graph represented by the $\{\mathsf{edge}\}$-structure $\mathbb{G}$. We associate the following objects with $G$:
  • let $\mathcal{B}$ be the family of splits of $G$ and $\mathbb{B}$ the $\{\mathsf{BIPART}\}$-structure modeling the weakly-bipartitive bipartition system $(V(G),\mathcal{B})$ with $U_{\mathbb{B}}=V(G)$;
  • let $T$ be the laminar tree induced by the weakly-bipartitive family $\mathcal{B}$ and let $\mathbb{T}$ be the $\{\mathsf{t}\text{-}\mathsf{edge}\}$-structure modeling it with $U_{\mathbb{T}}=V(T)$ and $L(T)=V(G)\subset U_{\mathbb{T}}$;
  • let $(H,F)$ be the enriched canonical split decomposition of $G$ represented by the $\{\mathsf{t}\text{-}\mathsf{edge},\mathsf{c}\text{-}\mathsf{edge}\}$-structure $\mathbb{H}$.
Our $\operatorname{C}_{2}{\!\!}\operatorname{MSO}$-transduction is obtained by composing the following transductions:
  • $\tau_{1}$: an $\operatorname{MSO}$-interprestation which outputs the $\{\mathsf{edge},\mathsf{BIPART}\}$-structure $\mathbb{G}\sqcup\mathbb{B}$ from $\mathbb{G}$;
  • $\tau_{2}$: the non-deterministic $\operatorname{C}_{2}{\!\!}\operatorname{MSO}$-transduction from Lemma 5.2 which outputs the $\{\mathsf{edge},\mathsf{BIPART},\mathsf{t}\text{-}\mathsf{edge}\}$-structure $\mathbb{G}\sqcup\mathbb{B}\sqcup\mathbb{T}$ on input $\mathbb{G}\sqcup\mathbb{S}$;
  • $\tau_{3}$: an $\operatorname{MSO}$-transduction which produces the $\{\mathsf{t}\text{-}\mathsf{edge},\mathsf{c}\text{-}\mathsf{edge}\}$-structure $\mathbb{H}$ from $\mathbb{G}\sqcup\mathbb{B}\sqcup\mathbb{T}$.
To define $\tau_{1}$ it is sufficient to observe that there is an $\operatorname{MSO}$-formula $\phi_{\mathsf{BIPART}}(X,Y)$ with two monadic free-variables $X$ and $Y$ such that $\phi_{\mathsf{BIPART}}$ is satisfied exactly when $\{X,Y\}$ is a split in $G$. As $\tau_{2}$ is constructed in Lemma 5.2, we are left with describing how to obtain $\tau_{3}$. We first need to find unique representative for the vertices in $V(H)$. Recall that $V(H)$ consists of two marker vertices $v_{X}^{Y}$ and $v_{Y}^{X}$ for every strong bipartition $\{X,Y\}\in\mathcal{B}$. For ease of notation, we use the name $v^{V(G)\setminus\{u\}}_{\{u\}}$ for the leaf $u\in V(G)$ of $T$. Additionally, we set $v_{s}^{t}:=v_{X}^{Y}$ and $v_{t}^{s}:=v_{Y}^{X}$ where $\{s,t\}\in E(T)$ is the edge for which $L(T_{s})=X$ and $L(T_{t})=Y$ where $T_{s}$ is the component of $T-e$ containing $s$ and $T_{t}$ the component containing $t$. We use the following atomic transductions to ensure unique representatives for all vertices in $V(H)$. We first guess a colouring $R$ which identifies one non-leaf element of $\mathbb{G}\sqcup\mathbb{B}\sqcup\mathbb{T}$ as root. We use this root element to interpret an auxiliary binary relation $\mathsf{parent}$ in the usual way. Now we copy the structure $\mathbb{G}\sqcup\mathbb{B}\sqcup\mathbb{T}$ once introducing the binary relation $\mathsf{copy}_{1}$ in which $\mathsf{copy}_{1}(x,y)$ indicates that $x$ is the copy of $y$. Assume that $\{s,t\}$ is an edge in $T$ corresponding to a strong bipartition in $\mathcal{B}$ and $\mathsf{parent}(s,t)$ is satisfied. We use the original node $t$ to represent the marker $v_{t}^{s}$ and we use the copy of $t$ to represent the marker $v_{s}^{t}$. As every node has a unique parent in a rooted tree, for each marker the chosen representatives are unique. To express this we use the predicate $\mathsf{marker}(x,y,z)$ which expresses that $x$ is the representative of the marker $v_{y}^{z}$ defined as follows:
$\mathsf{marker}(x,y,z):=\Big{(}\mathsf{parent}(z,y)\land x=y\Big{)}\lor\Big{(}\mathsf{parent}(y,z)\land\mathsf{copy}_{1}(x,z)\Big{)}.$
We now filter the universe only keeping all representatives of markers (recall that markers $v_{\{v\}}^{V(G)\setminus\{v\}}$ are the original vertices of $G$). We are left with defining the two edge relations $\mathsf{t}\text{-}\mathsf{edge}$ and $\mathsf{c}\text{-}\mathsf{edge}$. To do so, we use the previously defined functional predicate $\mathsf{leafset}(t,s)$.
First note that markers $v_{t}^{s}$ and $v_{t^{\prime}}^{s^{\prime}}$ are in the same component of $H$ if and only if $t=t^{\prime}$. By Lemma 5.10 $v_{t}^{s}v_{t}^{s^{\prime}}$ is an c-edge in $(H,F)$ if and only if there is an edge $uu^{\prime}\in E(G)$ with $u\in L(T^{t}_{s})$ and $u^{\prime}\in L(T^{t}_{s^{\prime}})$. Consequently, we can express $\mathsf{c}\text{-}\mathsf{edge}$ as follows:
$\phi_{\mathsf{c}\text{-}\mathsf{edge}}(x,x^{\prime}):=$ $\exists s\exists s^{\prime}\exists t\exists y\exists y^{\prime}\Big{(}\mathsf{t}\text{-}\mathsf{edge}(t,s)\land\mathsf{t}\text{-}\mathsf{edge}(t,s^{\prime})\land\mathsf{marker}(x,t,s)\land$
$\mathsf{marker}(x^{\prime},t,s^{\prime})\land y\in\mathsf{leafset}(t,s)\land y^{\prime}\in\mathsf{leafset}(t,s^{\prime})\land E(y,y^{\prime})\Big{)}.$
On the other hand, the $\mathsf{t}\text{-}\mathsf{edge}$-relation in $\mathbb{H}$ is the relation between pairs of markers and hence can be easily (re)-defined using the $\mathsf{marker}$ predicate by:
$\phi_{\mathsf{t}\text{-}\mathsf{edge}}(x,x^{\prime}):=\exists y\exists z\big{(}\mathsf{marker}(x,y,z)\land\mathsf{marker}(x^{\prime},z,y)\big{)}.$
Finally, the transduction only keeps the relation $\mathsf{t}\text{-}\mathsf{edge}$ and $\mathsf{c}\text{-}\mathsf{edge}$.
5.3 Application to bi-join decomposition
Bi-joins were introduced in [18, 25] as a generalisation of modules and splits in undirected graphs and we follow their notation. A bi-join in an undirected graph $G$ is a bipartition $\{X,Y\}$ of $V(G)$ such that $\left|X\right|$, $\left|Y\right|\geq 1$ and there are subsets $X^{\prime}\subseteq X$ and $Y^{\prime}\subseteq Y$ such that $X^{\prime}\times Y^{\prime}\cup(X\setminus X^{\prime})\times(Y\setminus Y^{\prime})\subseteq E(G)$, and there are no additional edges between $X$ and $Y$ in $G$.
As for splits, a bi-join $\{X,Y\}$ is considered trivial if $\left|X\right|=1$ or $\left|Y\right|=1$, and a graph is considered prime wrt bi-join if it has only trivial bi-joins. Let us denote by $\mathcal{B}_{\mathsf{join}}^{G}$ the set of bi-joins of a graph $G$. A proof of the bipartitiveness of bi-joins can be found in [25].
Theorem 5.13 ([25]).
For every undirected graph $G$, the system $\mathcal{B}_{\mathsf{join}}^{G}$ is bipartitive.
A corollary of Theorem 5.13 and Lemma 5.2 we get the existence of a $\operatorname{C}_{2}{\!\!}\operatorname{MSO}$-transduction taking as input an undirected graph $G$ represented as the $\{\mathsf{edge}\}$-structure $\mathbb{G}$ and outputting the laminar tree representing the set of strong bi-joins of $G$. We would like however, as in the case of splits, to output a relational structure that not only represents the laminar tree induced by $\mathcal{B}_{\mathsf{join}}^{G}$, but also allows to $\operatorname{C}_{2}{\!\!}\operatorname{MSO}$-transduce $\mathbb{G}$. Let’s now explain how to construct such a structure from the laminar tree induced by $\mathcal{B}_{\mathsf{join}}^{G}$. There are two ways to compute such a structure and both are similar to the one for the enhanced canonical split decomposition. Let’s first explain how to compute the one from [17].
Transducing the Skeleton graph [17]
If $X$ is a subset of $V(G)$, let $\equiv_{X}$ be the relation on $X\times X$ where $x\equiv_{X}y$ if $N(x)\cap V(G)\setminus X=N(y)\cap V(G)\setminus X$. One easily checks that $\equiv_{X}$ is an equivalence relation and, if $\{X,Y\}$ is a bi-join, then $\equiv_{X}$ and $\equiv_{Y}$ both have at most two equivalence classes. Let $T$ be the bipartitive tree of $\mathcal{B}_{\mathsf{join}}^{G}$, and for a node $u$, let $u_{1}$ and $u_{2}$ be the, at most two, equivalence classes of $\equiv_{L(T_{u})}$. For every node $u$ of $T$, let’s denote by $G_{u}$ the graph whose vertex set is $\{v_{1},v_{2}\mid v$ is a neighbor of $u\}$, and there is an edge $z_{i}z_{j}$ in $G_{u}$ if there is $x\in z_{i}$ and $y\in z_{j}$ such that $xy\in E(G)$. The skeleton graph associated with $T$ is the graph obtained by taking first the disjoint union of the graphs $G_{u}$, for all nodes $u$ of $T$, and then adding t-edges between copies of same equivalence classes. It is worth mentioning that the skeleton graph is uniquely defined from $T$, and since $T$ is unique, up to isomorphism, one can conclude that the skeleton graph is unique, up to isomorphism. It is routine to check the following.
Lemma 5.14 ([17]).
Let $G$ be a connected graph and let $H$ be its skeleton graph. Then the following hold:
  1. 1.
    two vertices $x$ and $y$ of $G$ are adjacent if and only if there is an alternating path between $x$ and $y$ in $H$;
  2. 2.
    the bipartitive tree of $\mathcal{B}_{\mathsf{join}}^{G}$ can be obtained from $H$ by contracting the graphs $G_{u}$, except vertices of $G$, and considering the $\mathsf{t}$-edges as its set of edges.
We are now ready to transduce the skeleton graph of a graph, the proof of which is similar to the one for the canonical split decomposition.
Theorem 5.15.
There exists a non-deterministic $\operatorname{C}_{2}{\!\!}\operatorname{MSO}$-transduction $\tau$ such that for any connected graph $G$ represented as the $\{\mathsf{edge}\}$-structure $\mathbb{G}$, $\tau(\mathbb{G})$ is non-empty and every output in $\tau(\mathbb{G})$ is equal to some $\{\mathsf{t}\text{-}\mathsf{edge},\mathsf{c}\text{-}\mathsf{edge}\}$-structure $\mathbb{H}$ representing the skeleton graph of $G$.
Proof 5.16.
Let $G$ be a connected graph represented by the $\{\mathsf{edge}\}$-structure $\mathbb{G}$, $(V(G),\mathcal{B}_{\mathsf{join}}^{G})$ the bipartitive set system consisting of all bi-joins of $G$ represented by the $\{\mathsf{BIPART}\}$-structure $\mathbb{B}$, $H$ its skeleton graph represented by the $\{\mathsf{t}\text{-}\mathsf{edge},\mathsf{c}\text{-}\mathsf{edge}\}$-structure $\mathbb{H}$ and $T$ represented by the $\{\mathsf{t}\text{-}\mathsf{edge},\mathsf{degenerate}\}$-structure, the bipartitive tree of $(V(G),\mathcal{B}_{\mathsf{join}}^{G})$.
Our first step is to transduce $\mathbb{B}$ from $\mathbb{G}$. For this it is sufficient to observe that we can easily define an $\operatorname{MSO}$-formula $\phi_{\mathsf{BIPART}}(X,Y)$ with two monadic free-variables $X$ and $Y$ which defines bi-joins $\{X,Y\}$ of $G$.
Since $\mathbb{B}$ is bipartitive we can apply Lemma 5.2 to transduce $\mathbb{G}\sqcup\mathbb{B}\sqcup\mathbb{T}$ from $\mathbb{G}\sqcup\mathbb{B}$. In the following we describe how to transduce $\mathbb{H}$ from $\mathbb{G}\sqcup\mathbb{B}\sqcup\mathbb{T}$.
Contrary to the case of the canonical split decomposition, we cannot deterministically deduce, from $T$, the four markers associated with the bi-join displayed by $T$. However, since the set $\{L(T_{u})\mid u$ is a node of $T\}$ is a laminar set system, one can use Corollary 3.17 to non-deterministically associate with each node $u$ of $T$ a leaf as done in the proof of Theorem 3.1. For each node $u$, let’s denote by $u_{1}$ this representative. Knowing $u_{1}$ and $u$, one can make a copy of $u$ and relate it to the equivalence of $\equiv_{L(T_{u})}$ not containing $u_{1}$. Now that we can create copies, for each node $u$, representing the two equivalence classes of $\equiv_{u}$, we leave, to the reader, the description of the transduction outputting the desired skeleton graph, as one can use the same copying technique as the one in Theorem 5.11.
Transducing a decomposition of its Seidel switch equivalence class [25]
Another way to represent the bipartitive tree and the graph itself is described in [18, 25] which uses graphs characteristic graphs as the ones used to describe a graph with its modular decomposition (see Section 4.2). Again, let $T$ be the bipartitive tree of $\mathcal{B}_{\mathsf{join}}^{G}$. For a node $u$ and $x_{1},\ldots,x_{p}$ leaves in the connected components of $T\setminus u$, not two of them in the same connected component, let $G[x_{1},\ldots,x_{p}]$ be the characteristic graph associated with $u$ and $\{x_{1},\ldots,x_{p}\}$. Unfortunately, while in the case of modular decomposition this graph is unique, and in the case of split decomposition one can choose it in a canonical way in order to describe the edges of $G$, there may be many characteristic graphs that can be associated with $u$. However, they are all degenerate if and only if $u$ is.
Lemma 5.17 ([18, 25]).
Let $G$ be a graph and let $T$ be the partitive tree of $\mathcal{B}_{\mathsf{join}}^{G}$. For every node $u$ of $T$ and every characteristic graph $G_{u}$ of $u$, $G_{u}$ is degenerate if and only if $u$ is degenerate.
One can moreover show that they are equivalent with respect to Seidel switch [25]. For a graph $H$ and $W$ a subset of $V(H)$, the Seidel switch of $H$ at $W$, denoted by $\overline{H}^{W}$, is the graph with $V(H)$ as vertex set and edge set $E(H[W])\cup E(H\setminus W)\cup\{xy\mid x\in W,y\notin W$ and $xy\notin E(H)\}$ [26]. One checks that if $H^{\prime}$ is obtained by a Seidel switch on $W^{\prime}$ in $\overline{H}^{W}$, then $H^{\prime}=\overline{H}^{W\operatorname*{\bigtriangleup}W^{\prime}}$ (see for instance [19, Lemma 11.1]). Now, two graphs $H$ and $H^{\prime}$ are Seidel equivalent if one is isomorphic to a Seidel switch of the other. The following tells that all the graphs in a Seidel equivalece class have the same set of bi-joins.
Lemma 5.18 ([25]).
Let $H$ be a graph and let $W$ be a subset of $V(H)$. Then, $\{X,Y\}$ is a bi-join in $H$ if and only if $\{X,Y\}$ is a bi-join of $\overline{H}^{W}$.
It has been proved that Seidel equivalence classes are in bijection with the set of two-graphs [26]. A two-graph is a pair $(V,\Omega)$ where $\Omega$ is a family of subsets of size $3$ on the finite set $V$ such that each subset of size $4$ of $V$ contains an even number of subsets in $\Omega$. For every undirected graph $H$, one can associate a two-graph $\Omega(H)$ where $\{x,y,z\}$ belongs to $\Omega(H)$ if $x,y$ and $z$ are all pairwise distinct and the number of edges in $H[\{x,y,z\}]$ is odd. Conversely, if $(V,\Omega)$ is a two-graph and $x$ is in $V$, then one can construct a graph $H(\Omega,x)$ with vertex set $V$ and edge set $\{yz\mid\{x,y,z\}\in\Omega\}$. The following is proved in [19, 26].
Theorem 5.19 ([19, Theorem 11.7]).
Let $V$ be a finite set of size at least $3$. Then the following hold:
  1. 1.
    for every two-graph $(V,\Delta)$, the graph $H(\Delta,x)$ is such that $\Omega(H(\Delta,x))=(V,\Delta)$;
  2. 2.
    two graphs $G$ and $H$ with vertex set $V$ are Seidel equivalent if and only if $\Omega(G)=\Omega(H)$.
For two two-graphs $(V_{1},\Omega_{1})$ and $(V_{2},\Omega_{2})$ and $v_{1}\in V_{1}$ and $v_{2}\in V_{2}$, their join wrt $v_{1}$ and $v_{2}$ is the two-graph $(V_{1}\cup V_{2}\setminus\{v_{1},v_{2}\},\Omega)$ where $\Omega=\{\{x,y,z\}\subseteq V_{1}\cup V_{2}\setminus\{v_{1},v_{2}\}\mid\{x,y,z\}\in\Omega_{1}\cup\Omega_{2}$ or $\{x,y,v_{1}\}\in\Omega_{1},\{z,t,v_{2}\}\in\Omega_{2}$ for some $t$, or $\{x,y,v_{2}\}\in\Omega_{2},\{z,t,v_{1}\}\in\Omega_{1}$ for some $t\}$.
A corollary of Theorem 5.19 is that if $\{X,Y\}$ is a bi-join of a graph $H$ with $X\times Y\cap E(H)=X^{\prime}\times Y^{\prime}\cup(X\setminus X^{\prime})\times(Y\setminus Y^{\prime})$, if we let $H[X]$ (resp. $H[Y]$) as defined for split decomposition with the marker $v_{X}^{Y}$ (resp. $v_{Y}^{X}$) adjacent to $X^{\prime}$ (resp. $Y^{\prime}$), $\Omega(H)$ is equal to the join of $\Omega(H[X])$ and of $\Omega(H[Y])$ wrt $v_{X}^{Y}$ and $v_{Y}^{X}$. As observed in [25], making $v_{X}^{Y}$ (resp. $v_{Y}^{X}$) adjacent to $X\setminus X^{\prime}$ (resp. $Y\setminus Y^{\prime}$) gives the same two-graph. As a corollary, if we associate with each node $u$ of the bipartitive tree $T$ of $G$ any characteristic graph and join all of them using t-edges as in the split decomposition, one obtains a structure, called Seidel skeleton, representing the Seidel equivalence class of $G$, which can by definition transduces $G$. We can summarise then the construction of the desired structure with the following lemma.
Lemma 5.20.
Let $T$ be the bipartive tree of $\mathcal{B}_{\mathsf{join}}^{G}$ and let $S(G,T)$ be the Seidel skeleton of $T$. If $S(G,T)$ is considered as a split decomposition and $H$ is the graph it defines, then $G$ is Seidel equivalent to $H$.
A direct corollary is the desired transduction.
Theorem 5.21.
There exists a non-deterministic $\operatorname{C}_{2}{\!\!}\operatorname{MSO}$-transduction $\tau$ such that for any connected graph $G$ represented as the $\{\mathsf{edge}\}$-structure $\mathbb{G}$, $\tau(\mathbb{G})$ is non-empty and every output in $\tau(\mathbb{G})$ is equal to some $\{\mathsf{t}\text{-}\mathsf{edge},\mathsf{c}\text{-}\mathsf{edge}\}$-structure $\mathbb{H}$ representing the Seidel skeleton of $G$.
Notice that the skeleton graph allows to transduce deterministically $G$ while the structure described here transduces $G$ non-deterministically by guessing the set $W$ such that Seidel switching it gives $G$.
6 Conclusion
We provide transductions for obtaining tree-like graph decompositions such as modular decompositions, cotrees, split decompositions and bi-join decompositions from a graph using $\operatorname{C}{\!\!}\operatorname{MSO}$. This improves upon results of Courcelle [10, 12] who gave such transductions for ordered graphs. It is worth mentionning that Theorem 5.5 can be also used to $\operatorname{C}{\!\!}\operatorname{MSO}$-transduce canonical decompositions of other structures such as Tutte’s decomposition of matroids or generally split-decompositions of submodular functions [16] or modular decompositions of $2$-structures [19] or of hypergraphs [22]. As shown by the application given in [12] for transducing Whitney’s isomorphism class of a graph, a line of research is to more investigate which structures can be $\operatorname{C}{\!\!}\operatorname{MSO}$-transduced from a graph or a set system by using the transductions from Theorem 1.1. Also, naturally, the question arises whether counting is necessary or whether $\operatorname{MSO}$ is sufficient to transduce such decompositions. Furthermore, our results include that transducing rank decompositions for graphs of rank-width 1 is possible using $\operatorname{C}{\!\!}\operatorname{MSO}$. But it is not known whether rank-decompositions can be transduced in general using some suitable extension of $\operatorname{MSO}$. Nevertheless, a corollary of our results is that $\operatorname{C}{\!\!}\operatorname{MSO}$-transducing rank-decompositions can be now reduced to consider $\operatorname{C}{\!\!}\operatorname{MSO}$-transducing rank-decompositions of prime graphs wrt either modular decomposition or split-decomposition as if a graph has rank-width at least $2$, then its rank-width is equal to the rank-width of its prime induced graphs wrt either modular or split decomposition. Thus, our results imply that, for any fixed $k$, there is a $\operatorname{C}{\!\!}\operatorname{MSO}$-transduction that computes a clique-decomposition of small width for any graph belonging to a graph class whose prime graphs have sizes bounded by $k$ or prime graphs have linear clique-width bounded by $k$, e.g., many $H$-free graphs have small prime graphs or have small linear clique-width (see for instance [6] or [24] for prominent such examples).
Appendix A Deferred proofs of Section 5
Proof A.1 (Proof of Lemma 5.1).
Let $\mathcal{B}$ be a family of bipartitions of a set $U$ and $a\in U$ any element. First observe that $(U,\mathcal{S}_{a})$ is a set system as $\emptyset\notin\mathcal{S}_{a}$ (since $\{\emptyset,U\}\notin\mathcal{B}$), $U\in\mathcal{S}_{a}$, and $\{b\}\in\mathcal{S}_{a}$ for every $b\in U$ (since $\{a\}\in\mathcal{S}_{a}$ and for $b\neq a$, $\{\{b\},U\setminus\{b\}\}\in\mathcal{B}$).
To argue the second assertion, assume that $T$ is the laminar tree of the set system $(U,\mathcal{S}_{a})$. First observe that in the set system $(U,\mathcal{S}_{a})$ the element $a$ is not contained in any set apart from $U$ and $\{a\}$ and hence the leaf $\{a\}$ is a child of the root $U$ of $T$. Additionally, the set $U\setminus\{a\}$ is in $S_{a}$ as the bipartition ($(\{a\},U\setminus\{a\})$ is in $\mathcal{B}$ by definition and hence $U\setminus\{a\}$ is also a child of the root $U$ of $T$. Let $T^{\prime}$ be the tree obtained from $T$ by removing node $U$, making all edges undirected and adding an edge from node $\{a\}$ to node $U\setminus\{a\}$. Indeed, by our previous observations, $T^{\prime}$ is an undirected tree. Furthermore, there is a correspondence between elements of $U$ and the leaves of $T^{\prime}$ (element $b\in U$ corresponds to leaf $\{b\}$ of $T^{\prime}$). To check that $T^{\prime}$ is indeed the laminar tree of $T$, we need to verify that for each bipartition $\{X,Y\}$ in $\mathcal{B}$ there is exactly one edge $e$ of $T^{\prime}$ such that $X=L(T^{\prime}_{1})$ and $Y=L(T^{\prime}_{2})$ for the two connected components $T^{\prime}_{1},T^{\prime}_{2}$ of $T^{\prime}-e$. We say that the edge $e$ implements the bipartition $\{X,Y\}$. Assume that $\{X,Y\}$ is any bipartition from $\mathcal{B}$. First consider the case that $\{X,Y\}$ is the bipartition $\{\{a\},U\setminus\{a\}\}$. Then $\{X,Y\}$ is clearly implemented by the edge $\{\{a\},U\setminus\{a\}\}$ of $T^{\prime}$. Now assume that $\{X,Y\}$ is any other bipartition in $\mathcal{B}$ and assume, without loss of generality, that $a\in Y$. Hence, $X\in\mathcal{S}_{a}$. Let $X^{\prime}$ be the parent of $X$ in $T$. Note that $X$ must be a proper subset of $U\setminus\{a\}$ as by our assumption $\{X,Y\}$ is not the partition $\{\{a\},U\setminus\{a\}$ and hence $X^{\prime}$ cannot be $U$. As we only deleted edge incident to $U$ in our construction of $T^{\prime}$, $\{X,X^{\prime}\}$ is an edge in $T^{\prime}$ which clearly implements the bipartition $\{X,Y\}$. Finally, every bipartition $\{X,Y\}\in\mathcal{B}$ is clearly implemented by only one edge as every inner node of the tree $T^{\prime}$ has degree at least three. Therefore, $T^{\prime}$ is the laminar tree of $(U,\mathcal{B})$.