### v4

Abstract.
We give $\operatorname{C}{\!\!}\operatorname{MSO}$-transductions that, given a graph $G$, output its modular decomposition, its split decomposition and its bi-join decomposition. This improves results by Courcelle [Logical Methods in Computer Science, 2006] who gave such transductions using order-invariant $\operatorname{MSO}$, a strictly more expressive logic than $\operatorname{C}{\!\!}\operatorname{MSO}$. Our methods more generally yield $\operatorname{C}_{2}{\!\!}\operatorname{MSO}$-transductions that output the canonical decompositions of weakly-partitive set systems and weakly-bipartitive systems of bipartitions.
Key words and phrases:
MSO-transduction, MSO-definability, graph decomposisions
Rutger Campbell: Supported by the Institute for Basic Science (IBS-R029-C1).
Mamadou Moustapha Kanté: Supported by the French National Research Agency (ANR-18-CE40-0025-01 and ANR-20-CE48-0002).
1. Introduction
A decomposition of a graph, especially a tree-like decomposition, is a result of recursive separations of a graph and is extremely useful for investigating combinatorial properties such as colourability, and for algorithm design. Such a decomposition also plays a fundamental role when one wants to understand the relationship between logic and a graph class. Different notions of the complexity of a separation motivate different ways to decompose, such as tree-decomposition, branch-decomposition, rank-decomposition and carving-decomposition. Furthermore, some important graph classes can be defined through the tree-like decomposition they admit; cographs with cotrees and distance-hereditary graphs with split decompositions being prominent examples.
For a logic $\mathsf{L}$, an $\mathsf{L}$-transduction is a non-deterministic map from a class of relational structures to a new class of relational structures using $\mathsf{L}$-formulas. Transducing a tree-like decomposition is of particular interest. Notably, transducing a decomposition of a graph implies that any property that is definable using a decomposition, is also definable on graphs that admit such a decomposition.To provide an example, a transduction constructing a modular decomposition of a graph $G$ can be used to define a sentence determining that the number of modules in $G$ is even (see Section 4.2 for definitions and Section 4.2 for details). Moreover, tree-like decompositions can be often represented by labelled trees, for which the equivalence of recognisability by a tree automaton and definability in $\operatorname{MSO}$ with modulo counting predicates, denoted $\operatorname{C}{\!\!}\operatorname{MSO}$, is well known [Cou90]. Hence, it is an interesting question to consider what kind of graph decompositions can be transduced using $\mathsf{L}$-transductions for some extension $\mathsf{L}$ of $\operatorname{MSO}$.
In a series of papers [Cou90, Cou91, Cou96, Cou06], Courcelle investigated the relationship between the graph properties that can be defined in an extension of $\operatorname{MSO}$ and the graph properties that can be recognized by a tree automaton where the tree-automaton receives a term representing the input graph. In particular, Courcelle’s theorem states that any graph property that is definable in the logic $\operatorname{C}{\!\!}\operatorname{MSO}$ can be recognized by a tree automaton receiving a tree-decompositions of bounded width [Cou90]. Combining this result with the linear time algorithm for computing tree-decompositions [Bod93], yields that $\operatorname{C}{\!\!}\operatorname{MSO}$ model-checking can be done in linear time on graphs of bounded treewidth. The converse statement – whether recognisability by a tree automaton implies definability in $\operatorname{C}{\!\!}\operatorname{MSO}$ on graphs of bounded treewidth – was conjectured by Courcelle in [Cou90] and finally settled by Bojańczyk and Pilipczuk [BP16]. The key step to obtain this result is obtaining a tree-decomposition of a graph via an $\operatorname{MSO}$-transduction, a strategy which was initially proposed in [Cou91] and is now standard.
The obvious next question is whether an analogous result can be proved for graphs of bounded clique-width and for more general combinatorial objects, most notably, matroids representable over a fixed finite field and of bounded branch-width. Due to the above-mentioned strategy, the key challenge is to produce corresponding tree-like decompositions by $\operatorname{MSO}$-transduction. It is known that clique-width decompositions can be $\operatorname{MSO}$-transduced for graphs of bounded linear clique-width [BGP21]. However, it is unknown if clique-width decompositions can be $\operatorname{MSO}$-transduced in general. In fact, this question remained open even for distance-hereditary graphs, which are precisely graphs of rank-width 1 (thus, of constant clique-width).
Besides tree-decompositions, the problem of transducing cotrees, and in general hierarchical decompositions such as modular decompositions and split decompositions were considered in the literature [Cou96, Cou99, Cou06, Cou13]. In [Cou96], Courcelle provides transductions using order-invariant $\operatorname{MSO}$ for cographs and modular decompositions of graphs of bounded modular width. Order-invariant $\operatorname{MSO}$ allows the use of a linear order of the set of vertices and is more expressive than $\operatorname{C}{\!\!}\operatorname{MSO}$ [GR08]. The applicability of these transductions was later generalized using the framework of weakly-partitive set systems [^1] Weakly-partitive set systems are set systems enjoying some nice closure properties, which were then used to show that some set systems allow canonical tree representations, see for instance the thesis by Montgolfier and Rao [dM03, Rao06]. to obtain order-invariant $\operatorname{MSO}$-transductions of modular and split decompositions [Cou06]. It was left as an open question whether one can get rid of the order (see for instance [Cou12] where an overview of the result on hierarchical decompositions was given).
1.1. Our results
In this paper, we consider decompositions given by nested partitions. We view partitions of a given kind as a ‘set system’. A set system consists of a set $U$, the universe, and a set $\mathcal{S}$ of subsets of $U$. Two sets overlap if they have non-empty intersection but neither of them is contained in the other. If no two elements in a set system $(U,\mathcal{S})$ overlap, i.e. the set system is laminar, then the subset relation in $(U,\mathcal{S})$ can be described by a rooted tree $T$, called the laminar tree of $(U,\mathcal{S})$. For any set system $(U,\mathcal{S})$ we can look at the subset of strong sets, i.e., sets that do not overlap with any other set, and the laminar tree $T$ they induce.
Given a graph $G$ we can consider the set system $(V(G),\mathcal{M})$ where $\mathcal{M}$ is the set of all modules in $G$. We obtain the modular decomposition, or the cotree in case $G$ is a cograph, by equipping the laminar tree of the suitable set system with some additional structure. The additional structure allows us to recover the graph from the respective decomposition.
Abstractly, the set systems mentioned are instances of weakly-partitive set systems. Roughly speaking, if a set system $(U,\mathcal{S})$ is well behaved, i.e. $(U,\mathcal{S})$ is weakly-partitive (definition in Section 2), then there is a tree $T$, a labelling $\lambda$ of $T$ and a partial order << of its nodes such that $(U,\mathcal{S})$ is completely described by $(T,\lambda,<)$ [CHM81, Rao06]. See Figure 1 for an example.
Figure 1. Left: A weakly-partitive set system for which strong sets are indicated by red, thick lines and singletons are omitted. Right: The laminar tree obtain by considering the laminar set system consisting of the strong sets. Here the node labeled $\mathsf{linear}$ plus the linear ordering of its children indicate that the leaves of every << interval corresponds to a set in the set system (e.g. $\{a,b,c\}$ is in the set system while $\{h,f,e,c\}$ is not). Note that some labels of nodes are omitted.
We show the following, wherein each item $\lambda$ is a suitable labelling of the nodes of the laminar tree $T$, << is a partial ordering of its nodes, and $F$ is an additional edge relation defined only on pairs of siblings in $T$. A visualization how results depend on each other is given in Figure 3.
Theorem 1.
There are non-deterministic $\operatorname{C}_{2}{\!\!}\operatorname{MSO}$-transductions $\tau_{1},\ldots,\tau_{4}$ such that:
  1. (1)
    For any laminar set system $(U,\mathcal{S})$, $\tau_{1}(U,\mathcal{S})$ is non-empty and every output in $\tau_{1}(U,\mathcal{S})$ is a laminar tree $T$ of $(U,\mathcal{S})$ (3);
  2. (2)
    For any weakly-partitive set system $(U,\mathcal{S})$, $\tau_{2}(U,\mathcal{S})$ is non-empty and every output in $\tau_{2}(U,\mathcal{S})$ is a weakly-partitive tree $(T,\lambda,<)$ of $(U,\mathcal{S})$ (15);
  3. (3)
    For any graph $G$, $\tau_{3}(G)$ is non-empty and every output in $\tau_{3}(G)$ is a modular decomposition $(T,F)$ of $G$ (17);
  4. (4)
    For any cograph $G$, $\tau_{4}(G)$ is non-empty and every output in $\tau_{4}(G)$ is a cotree $(T,\lambda)$ of $G$ (19).
Other tree-like graph decompositions can be obtained by considering systems of bipartitions. Aiming to study such decompositions, we can apply our techniques to systems of bipartitions. A systems of bipartitions consists of a universe $U$ and a set $\mathcal{B}$ of bipartitions of $U$. Two bipartitions of $U$ overlap if neither side of one of the bipartition is contained in either side of the other bipartition. In case $(U,\mathcal{B})$ has no overlapping bipartitions, then $(U,\mathcal{B})$ can be described by an undirected tree, also called laminar tree, in which bipartitions correspond to edge cuts. We can define the concept of strong bipartitions equivalently and consider the laminar tree induced by the strong bipartitions in $(U,\mathcal{B})$.
Given a graph $G$, we consider the system of bipartition $(V(G),\mathcal{S})$ where $\mathcal{S}$ contains all splits in $G$, or the system of bipartitions $(V(G),\mathcal{B})$ where $\mathcal{B}$ contains all bi-joins in $G$. Equipping the laminar tree of the respective systems of bipartitions with additional structure yields the split decomposition or bi-join decomposition of the graph. These systems of bipartitions are examples of weakly-bipartitive systems of bipartitions. Similar to weakly-partitive set systems, we can completely describe any weakly-bipartitive system of biparitions by a tree $T$, a labelling $\lambda$ and linear order $_{t}$ of the children of some particular nodes $t$ [dM03]. See Figure 1 for an example.
Figure 2. Left: A weakly-bipartitive system of bipartitions for which strong bipartitions are indicated by red, thick lines and bipartitions of the form $\{\{a\},U\setminus\{a\}\}$ are omitted. Right: The laminar tree obtain by considering the laminar system of bipartitions consisting of the strong bipartitions. The node labeled $\mathsf{linear}$ plus the linear ordering of its children indicate that the leaves of every << interval corresponds to a bipartition in the system of biparitions (e.g. $\{\{a,b,c,i,g\},\{d,e,f,h\}\}$ is in the system of bipartitions while $\{\{a,b,d\},\{c,e,f,g,h,i\}\}$ is not). Note that some labels of nodes are omitted.
We prove the following theorem, where similarly to before each item $\lambda$ is a suitable labelling of the nodes of the laminar tree $T$, << is a partial ordering of its nodes, and $F$ is an additional edge relation defined only on pairs of siblings in $T$.
Theorem 2.
There are non-deterministic $\operatorname{C}_{2}{\!\!}\operatorname{MSO}$-transductions $\tau_{1},\ldots,\tau_{3}$ such that:
  1. (1)
    For any weakly-bipartitive set system $(U,\mathcal{B})$, $\tau_{1}(U,\mathcal{B})$ is non-empty and every output in $\tau_{1}(U,\mathcal{B})$ is a weakly-bipartitive tree $(T,\lambda,<)$ of $(U,\mathcal{B})$ (22);
  2. (2)
    For any graph $G$, $\tau_{2}(G)$ is non-empty and every output in $\tau_{2}(G)$ is a split decomposition $(T,F)$ of $G$ (25);
  3. (3)
    For any graph $G$, $\tau_{3}(G)$ is non-empty and every output in $\tau_{3}(G)$ is a bi-join decomposition $(T,F)$ of $G$ (26).
The key step in obtaining these transductions is to transduce the laminar tree $T$ of a set system $(U,\mathcal{S})$. The crux here is to find a suitable representative of each node of $T$ amongst the elements of $U$ and a non-deterministic coloring which allows the assignment of representatives to nodes by means of a $\operatorname{C}_{2}{\!\!}\operatorname{MSO}$-formula. It should be mentioned that a similar result is claimed in the preprint [Boj23], where a proof sketch designing a $\operatorname{C}_{3}{\!\!}\operatorname{MSO}$-transduction is described. Once the laminar tree is obtained, the additional relations for each decomposition can be obtained using a deterministic $\operatorname{MSO}$-transduction. Notice that for each of these transductions, there exists an inverse deterministic $\operatorname{MSO}$-transduction, namely a transduction that from the tree-like decomposition outputs the original structure. It is worth mentionning that a corollary of 1(2) is a $\operatorname{C}{\!\!}\operatorname{MSO}$-transduction outputting a rank-decomposition of width $1$ for every graph of rank-width $1$.
Section 3Section 4Section 5 3blablabla laminar set system $\rightarrow$ laminar tree 14blablabla weakly-partitive set system $\rightarrow$ laminar tree 21blablabla weakly-bipartitive set system $\rightarrow$ laminar tree 15blablabla weakly-partitive set system $\rightarrow$ weakly-partitive tree 17blablabla graph $\rightarrow$ modular decomposition 22blablabla weakly-bipartitive set system $\rightarrow$ weakly-bipartitive tree 25blablabla graph $\rightarrow$ split decomposition 19blablabla cograph $\rightarrow$ cotree 26 graph $\rightarrow$ bi-join decomposition
Figure 3. Overview of the various transductions in the paper. An arrow from $x$ to $y$ indicates that result $x$ is used in the proof of result $y$.
1.2. Organization
The paper is organized as follows. In Section 2 we introduce terminology and notation needed. In Section 3 we prove 3. In Section 4 we provide the proof of 15 and 17 and obtain 19. In Section 5 we provide the proofs of $2$, 25 and 26.
2. Preliminaries
2.1. Graphs, trees, set systems
Graphs
We use standard terminology of graph theory, and we fix some notations. We consider graphs to be finite. Given a directed graph $G$, its sets of vertices and edges are denoted by $V(G)$ and $E(G)$, respectively. We denote by uv an edge $(u,v)\in E(G)$. An undirected graph is no more than a directed graph for which $E(G)$ is symmetric (i.e., $uv\in E(G)\iff vu\in E(G)$). The notions of paths, connected components, etc… are defined as usual. Given a subset $Z$ of $V(G)$, we denote by $G[Z]$ the sub-graph of $G$ induced by $Z$.
Trees
A tree is a connected undirected graph without cycles. In the context of trees, we use a slightly different terminology than for graphs. In particular, vertices are called nodes, nodes of degree at most $1$ are called leaves, and nodes of degree greater than $1$ are called inner. The set of leaves is denoted $L(T)$; thus the set of inner nodes is $V(T)\setminus L(T)$. For a node $t$ of a tree $T$ and a neighbour $s$ of $t$, we denote by $T^{t}_{s}$ the connected component of TtT-t containing $s$. We sometimes consider rooted trees, namely trees with a distinguished node, called the root. Rooted trees enjoy a natural orientation of their edges toward the root, which induces the usual notions of parent, child, sibling, ancestor and descendant. Hence, we represent a rooted tree by a set of nodes with an ancestor/descendant relationship (instead of specifying the root). We use the convention that every node is one of its own ancestors and descendants. We refer to ancestors (resp. descendants) of a node that are not the node itself as proper ancestors (resp. proper descendants). For a node $t$ of a rooted tree $T$, we denote by $T_{t}$ the subtree of $T$ rooted in $t$ (i.e., the restriction of $T$ to the set of descendants of $t$).
Set systems and laminar trees
A set system is a pair $(U,\mathcal{S})$ where $U$ is a finite set, called the universe, and $\mathcal{S}$ is a family of subsets of $U$ where $\emptyset\notin\mathcal{S}$, $U\in\mathcal{S}$, and $\{a\}\in\mathcal{S}$ for every $a\in U$. [^2] Though these restrictions on $\mathcal{S}$ are not usual for set systems, it is convenient for our contribution and it does not significantly impact the generality of set systems: every family $\mathcal{F}$ of subsets of $U$ can be associated with a set system $(U,\mathcal{S})$ where $\mathcal{S}=\big{(}\mathcal{F}\setminus\{\emptyset\}\big{)}\cup\big{\{}U\big{\}}\cup\big{\{}\{a\}\mid a\in U\big{\}}$. Two sets $X$ and $Y$ overlap if they are neither disjoint nor related by containment. A set system $(U,\mathcal{S})$ is said to be laminar (aka overlap-free) when no two sets from $\mathcal{S}$ overlap. By extension, a set family $\mathcal{S}$ is laminar if $\left(\bigcup\mathcal{S},\mathcal{S}\right)$ is a laminar set system (note this also requires that $\emptyset\notin\mathcal{S}$, $\bigcup\mathcal{S}\in\mathcal{S}$, and $\{a\}\in\mathcal{S}$ for every $a\in\bigcup\mathcal{S}$).
A laminar family $\mathcal{S}$ of subsets of $U$ naturally defines a rooted tree where the nodes are the sets from $\mathcal{S}$, the root is $U$, and the ancestor relation corresponds to set inclusion, i.e., nodes corresponding to sets $S\subset S^{\prime}\in\mathcal{S}$ are adjacent in the tree if there is no proper superset of $S$ in $\mathcal{S}$ that is a proper subset of $S^{\prime}$. We call this rooted tree the laminar tree of $U$ induced by $\mathcal{S}$ (or laminar tree of $(U,\mathcal{S})$). In this rooted tree, the leaves are the singletons $\{x\}$ for $x\in U$, which we identify with the elements themselves. That is to say, $L(T)=U$. Laminar trees have the property that each inner node has at least two children due to the definition of set systems demanding each singleton being in the set system. Observe also that the size of a laminar tree is linearly bounded in the size of the universe.
2.2. Logic and transductions
We use relational structures to model both graphs and the various tree-like decompositions used in this paper. In order to concisely model set systems we use the more general notion of extended relational structures, namely relational structure extended with set predicate names. Such structures also naturally arise as outputs of $\operatorname{MSO}$-transductions defined below.
Define an (extended) vocabulary to be a set of symbols, each being either a relation name $R$ with associated arity $\operatorname{ar}(R)\in\mathbb{N}$, or a set predicate name $P$ with associated arity $\operatorname{ar}(P)\in\mathbb{N}$. Set predicate names are aimed to describe relations between sets, e.g., one may have a unary set predicate for selecting finite sets of even size, or a binary set predicate for selecting pairs of disjoint sets. We use capital $R$ or names starting with a lowercase letter (e.g., $\mathsf{edge}$, $\mathsf{ancestor}$, $\mathsf{t}\text{-}\mathsf{edge}$) for relation names, and capital $P$ or uppercase names (e.g., $\mathsf{SET}$, $\operatorname{C}_{2}$) for set predicate names. To refer to an arbitrary symbol of undetermined kind, we use capital $Q$. A relational vocabulary is an extended vocabulary in which every symbol is a relation name.
Let $\Sigma$ be a vocabulary. An extended relational structure over $\Sigma$ ($\Sigma$-structure) is a structure $\mathbb{A}=\langle U_{\mathbb{A}},(Q_{\mathbb{A}})_{Q\in\Sigma}\rangle$ consisting on the one hand of a set $U_{\mathbb{A}}$ called universe, and on the other hand, for each symbol $Q$ in $\Sigma$, an interpretation $Q_{\mathbb{A}}$ of $Q$, which is a relation of arity $\operatorname{ar}(Q)$ either over the universe if $Q$ is a relation name, or over the family of subsets of the universe if $Q$ is a set predicate name. When $\Sigma$ is not extended, $\mathbb{A}$ is simply a relational structure.
Given a $\Sigma$-structure $\mathbb{A}$ and, for some vocabulary $\Gamma$, a $\Gamma$-structure $\mathbb{B}$, we write $\mathbb{A}\sqsubseteq\mathbb{B}$ if $\Sigma\subseteq\Gamma$, $U_{\mathbb{A}}\subseteq U_{\mathbb{B}}$ and for each symbol $Q$ in $\Sigma$, $Q_{\mathbb{A}}=Q_{\mathbb{B}}$. [^3] We require equality here (in particular only elements or subsets of $U_{\mathbb{A}}$ are related in $Q_{\mathbb{B}}$). This differs from classical notions of inclusions of relational structures which typically require equality only on the restriction of the universe to $U_{\mathbb{A}}$, i.e., $Q_{\mathbb{A}}=Q_{\mathbb{B}_{/U_{\mathbb{A}}}}$, e.g., in order to correspond to induced graphs. We write $\mathbb{A}\sqcup\mathbb{B}$ to denote the $(\Sigma\cup\Gamma)$-structure consisting of the universe $U_{\mathbb{A}}\cup U_{\mathbb{B}}$ and, for each symbol $Q\in\Sigma\cup\Gamma$, the interpretation $Q_{\mathbb{A}\sqcup\mathbb{B}}$ which is $Q_{\mathbb{A}}$, $Q_{\mathbb{B}}$, or $Q_{\mathbb{A}}\cup Q_{\mathbb{B}}$ according to whether $Q$ belongs to $\Sigma\setminus\Gamma$, to $\Gamma\setminus\Sigma$, or to $\Sigma\cap\Gamma$. [^4] We do not require $\mathbb{A}$ and $\mathbb{B}$ to be disjoint structures whence we may have $\mathbb{A}\not\sqsubseteq\mathbb{A}\sqcup\mathbb{B}$.
[…]
We now fix some (extended) vocabularies that we will use.
Graphs:
To model both graphs, unrooted trees, and directed graphs, we use the relational vocabulary $\{\mathsf{edge}\}$ where $\mathsf{edge}$ is a relation name of arity $2$. A (directed) graph $G=(V,E)$ is modeled as the $\{\mathsf{edge}\}$-structure $\mathbb{G}$ with universe $U_{\mathbb{G}}=V$ and interpretation $\mathsf{edge}_{\mathbb{G}}=E$. In particular if $G$ is undirected, then $\mathsf{edge}_{\mathbb{G}}$ is symmetric.
Rooted trees:
We use the relational vocabulary $\{\mathsf{ancestor}\}$ to model rooted trees where $\mathsf{ancestor}$ is a relation name of arity $2$. A rooted tree $T$ is modeled as the $\{\mathsf{ancestor}\}$-structure $\mathbb{T}$ with universe $U_{\mathbb{T}}=V(T)$ and the interpretation $\mathsf{ancestor}_{\mathbb{T}}$ being the set of pairs $(u,v)$ such that $u$ is an ancestor of $v$ in $T$. It is routine to define $\operatorname{FO}$-formulas over this vocabulary to express the binary relations parent, child, proper ancestor, (proper) descendant, as well as the unary relations leaf and root.
Set systems:
To model set systems, we use the extended vocabulary $\{\mathsf{SET}\}$ where $\mathsf{SET}$ is a set predicate name of arity $1$. A set system $S=(U,\mathcal{S})$ is thus naturally modeled as the $\{\mathsf{SET}\}$-structure $\mathbb{S}$ with universe $U_{\mathbb{S}}=U$ and interpretation $\mathsf{SET}_{\mathbb{S}}=\mathcal{S}$.
Transductions
Let $\Sigma$ and $\Gamma$ be two extended vocabularies. A $\Sigma$-to-$\Gamma$ transduction is a set $\tau$ of pairs formed by a $\Sigma$-structure, call the input, and a $\Gamma$-structure, called the output. We write $\mathbb{B}\in\tau(\mathbb{A})$ when $(\mathbb{A},\mathbb{B})\in\tau$. We say that a class of $\Sigma$-structures $\mathcal{C}$ transduces a class of $\Gamma$-structures $\mathcal{D}$ if there is a $\Sigma$-to-$\Gamma$ transduction $\tau$ with $\mathcal{D}\subseteq\tau(\mathcal{C})$. When for every pair $(\mathbb{A},\mathbb{B})\in\tau$ we have $\mathbb{A}\sqsubseteq\mathbb{B}$, we call $\tau$ an overlay transduction. Overlay transductions are used to augment a relational structure by adding additional relations while not altering existing relations in any way. Some transductions can be defined by means of $\operatorname{MSO}$- or $\operatorname{C}_{2}{\!\!}\operatorname{MSO}$-formulas. This leads to the notion of $\operatorname{MSO}$- and $\operatorname{C}_{2}{\!\!}\operatorname{MSO}$-transductions. Following the presentation of [BGP21], for $\mathsf{L}$ denoting $\operatorname{MSO}$ or $\operatorname{C}_{2}{\!\!}\operatorname{MSO}$, define an $\mathsf{L}$-transduction to be a transduction obtained by composing a finite number of atomic $\mathsf{L}$-transductions of the following kinds.
Filtering:
An overlay transduction specified by an $\mathsf{L}$-sentence $\phi$ over the input vocabulary $\Sigma$, which discards the inputs that do not satisfy $\phi$ and keeps the other unchanged. Hence, it defines a partial function (actually, a partial identity) from $\Sigma$-structures to $\Sigma$-structures.
Universe restriction:
A transduction specified by an $\mathsf{L}$-formula $\phi$ over the input vocabulary $\Sigma$, with one free first-order variable, which restricts the universe to those elements that satisfy $\phi$. The output vocabulary is $\Sigma$ and the interpretation of every relation (resp. every predicate) in the output structure is defined as the restriction of its interpretation in the input structure to those tuples of elements satisfying $\phi$ (resp. tuples of sets of elements that satisfy $\phi$). This defines a total function from $\Sigma$-structures to $\Sigma$-structures.
Interpretation:
A transduction specified by a family $(\phi_{Q})_{Q\in\Gamma}$ over the input vocabulary $\Sigma$ where $\Gamma$ is the output vocabulary and each $\phi_{Q}$ has $\operatorname{ar}(Q)$ free variables which are first-order if $Q$ is a relation name and monadic if it is a set predicate name. The transduction outputs the $\Gamma$-structure that has the same universe as the input structure and in which each relation or predicate $Q$ is interpreted as those set of tuples that satisfy $\phi_{Q}$. This defines a total function from $\Sigma$-structures to $\Gamma$-structures.
Copying:
An overlay transduction parametrized by a positive integer $k$ that adds $k$ copies of each element to the universe. The output vocabulary consists in the input vocabulary $\Sigma$ extended with $k$ binary relational symbols $(\mathsf{copy}_{i})_{i\in[k]}$ interpreted as pairs of elements $(x,y)$ saying that “$y$ is the $i$-th copy of xx”. The interpretation of the relations (resp. predicates) of the input structure are preserved, on original elements. This defines a total function from $\Sigma$-structures to $\Gamma$-structures, where $\Gamma=\Sigma\cup\{\mathsf{copy}_{i}\mid i\in[k]\}$.
Colouring:
An overlay transduction that adds a new unary relation $\mathsf{color}\notin\Sigma$ to the structure. Any possible interpretation yields an output; indeed the interpretation is chosen non-deterministically. The interpretation of the relations (resp. predicates) of the input structure are preserved. Hence, it defines a total (non-functional) relation from $\Sigma$-structures to $\Gamma$-structures where $\Gamma=\Sigma\cup\{\mathsf{color}\}$. Note that by composing several transductions of this type any given element may receive multiple colors.
We say an $\mathsf{L}$-transduction is deterministic if it does not use colouring, it is non-deterministic otherwise. By definition, deterministic $\mathsf{L}$-transductions define partial functions. It is known that every $\operatorname{C}{\!\!}\operatorname{MSO}$-transduction can be put in a standard form in which each of the above atomic transductions occurs once in a fixed order, namely: first colouring, then filtering, copying, interpreting the output relations, and finally restricting the universe [CE12].
3. Transducing the laminar-tree
In this section, we present an overlay $\operatorname{C}_{2}{\!\!}\operatorname{MSO}$-transduction that takes as input a laminar set system and outputs the laminar tree it induces.
Theorem 3.
Let $\Sigma$ be an extended vocabulary, including a unary set predicate name $\mathsf{SET}$ and not including the binary relational symbol $\mathsf{ancestor}$. There exists a non-deterministic overlay $\operatorname{C}_{2}{\!\!}\operatorname{MSO}$-transduction $\tau$ such that, for each laminar set system $(U,\mathcal{S})$ represented as the $\{\mathsf{SET}\}$-structure $\mathbb{S}$ and inducing the laminar tree $T$ with L(T)=U{L(T)=U}, and for each $\Sigma$-structure $\mathbb{A}$ with $\mathbb{S}\sqsubseteq\mathbb{A}$, $\tau(\mathbb{A})$ is non-empty and every output in $\tau(\mathbb{A})$ is equal to $\mathbb{A}\sqcup\mathbb{T}$ for some $\{\mathsf{ancestor}\}$-structure $\mathbb{T}$ representing $T$.
Since the sets from $\mathcal{S}$ are precisely the sets of leaves of the subtrees of $T$, there is an $\operatorname{MSO}$-transduction which is the inverse of the above $\operatorname{C}_{2}{\!\!}\operatorname{MSO}$-transduction; that is to say, given an $\{\mathsf{ancestor}\}$-structure $\mathbb{T}$ representing the laminar tree, it outputs the original set system $\mathbb{S}$. Namely,
  1. (1)
    An interpretation $\Phi_{\mathsf{SET}}(S)$ that is true for a set $S$ when there exists an element $a$ such that an element $x$ is in $S$ if and only if $a$ is an ancestor of $x$.
  2. (2)
    The filtering $\phi(x)$ keeping leaves only.
We prove 3 in Section 3.2, using the key tools developed in Section 3.1, that allow us to represent each inner node of $T$ with a pair of leaves from its subtree, while keeping, for each leaf, the number of inner nodes it represents bounded.
3.1. Inner node representatives
In this section, the root of any rooted tree is always an inner node (unless the tree is a unique node). We fix a rooted tree $T$, in which every inner node has at least two children (a necessary assumption that is satisfied by laminar trees), and we let $V$ denote the set of its nodes ($V=V(T)$) and $L\subseteq V$ denote the subset of its leaves ($L=L(T)$).
Let $S\subseteq V\setminus L$ be a set of inner nodes, and let $(\pi,\sigma)$ be a pair of injective mappings from $S$ to $L$. We say that the pair $(\pi,\sigma)$ identifies $S$ if for each $\in S, ss$ is the least common ancestor of $\pi(s)$ and $\sigma(s)$. For $s\in S$ and $x\in V$, we say that $x$ is $s$-requested in $(\pi,\sigma)$ if $x$ lies on the path from $\pi(s)$ to $\sigma(s)$ (namely, on either of the paths from $\pi(s)$ to $s$ and from $\sigma(s)$ to $s$). We say that $x$ is requested in $(\pi,\sigma)$ if it is $s$-requested for some $s$. The pair $(\pi,\sigma)$ has unique request if every node $x$ of $T$ is requested at most once in $(\pi,\sigma)$, i.e., there exists at most one $s\in S$ such that $x$ is $s$-requested in $(\pi,\sigma)$. Note that if $(\pi,\sigma)$ has unique request then the paths from $\pi(s)$ to $\sigma(s)$ are pairwise disjoint for all the $s\in S$. We now state a few basic observations.
Remark 4.
Let $(\pi,\sigma)$ identifying some subset $S$ of $V\setminus L$.
  1. (1)
    The reversed pair $(\sigma,\pi)$ also identifies $S$ and has unique request whenever $(\pi,\sigma)$ does.
  2. (2)
    If $S^{\prime}\subset S$, then $(\pi_{|S^{\prime}},\sigma_{|S^{\prime}})$ identifies $S^{\prime}$, and has unique request if $(\pi,\sigma)$ does.
  3. (3)
    For each $\in S, ss$ is $s$-requested in $(\pi,\sigma)$.
  4. (4)
    If $(\sigma,\pi)$ has unique request, then $\pi(S)$ and $\sigma(S)$ are disjoint subsets of $L$.
Lemma 5.
Let $(\pi,\sigma)$ identifying some subset $S$ of $V\setminus L$ with unique request. Then for each $a\in\pi(S)$, the node $\pi^{-1}(a)$ is the least ancestor of $a$ which belongs to $S$.
Proof 3.1.
Let $a\in\pi(S)$ and let $s=\pi^{-1}(a)$. By definition, $s$ is an ancestor of $a$ which belongs to $S$. Let $y$ be the least ancestor of $a$ that is contained in $S$. As $s$ is an ancestor of $a$ belonging to $S$, $y$ must be a descendant of $s$. Hence, $y$ is $s$-requested in $(\pi,\sigma)$. Additionally, by item 3 of 4, $y$ is $y$-requested in $(\pi,\sigma)$. Since $(\pi,\sigma)$ has unique request, $y=s$. Thus, $s$ is the least ancestor of $a$ which belongs to $S$.
Notice that, by item 1 of 4, a similar result holds for each $b\in\sigma(S)$. It follows that the sets $\pi(S)$ and $\sigma(S)$ characterize $(\pi,\sigma)$.
Lemma 6.
Let $S\subseteq V\setminus L$, and $(\pi,\sigma)$ and $(\pi^{\prime},\sigma^{\prime})$ be two pairs of injections from $S$ to $L$ identifying $S$ with unique request. If $\pi(S)=\pi^{\prime}(S)$ and $\sigma(S)=\sigma^{\prime}(S)$, then $\pi=\pi^{\prime}$ and $\sigma=\sigma^{\prime}$.
Proof 3.2.
Let $s\in S$, $a=\pi(s)$, and $s^{\prime}=\pi^{\prime-1}(a)$. Both $s$ and $s^{\prime}$ are the least ancestor of $a$ which belongs to $S$, hence $s=s^{\prime}$ and $\pi^{\prime}(s)=a$. Thus $\pi=\pi^{\prime}$. By item 1 of 4, we also obtain that $\sigma=\sigma^{\prime}$.
Let $A$ and $B$ be two subsets of $L$ that are disjoint and of same cardinality. We call such a pair $(A,B)$ a bi-colouring. We say that $(A,B)$ identifies $S$ if there exists a pair $(\pi,\sigma)$ identifying $S$ with unique request such that $\pi(S)=A$ and $\sigma(S)=B$. By the previous lemma, for a fixed set $S\subseteq V\setminus L$ and a fixed bi-colouring $(A,B)$ of $L$, the pair $(\pi,\sigma)$ identifying $S$ with $\pi(S)=A$ and $\sigma(S)=B$ is unique when it exists. We will also prove that $S$ is actually uniquely determined from $(A,B)$. Before that, we state the following technical lemma.
Lemma 7.
Let $(A,B)$ identify some subset $S$ of $V\setminus L$ through a pair $(\pi,\sigma)$ of injections having unique request. Then, for each inner node $x$, exactly one of the three following cases holds:
  1. (1)
    $\notin S, xx$ is not requested in $(\pi,\sigma)$, and for each child $c$ of $x$, $\left|A\cap V(T_{c})\right|=\left|B\cap V(T_{c})\right|$;
  2. (2)
    $\notin S, xx$ is requested in $(\pi,\sigma)$, and there exists one leaf $z\in(A\cup B)\cap V(T_{x})$ such that, for each child $c$ of $x$, $\left|(A\setminus\{z\})\cap V(T_{c})\right|=\left|(B\setminus\{z\})\cap V(T_{c})\right|$;
  3. (3)
    $\in S, xx$ is requested in $(\pi,\sigma)$, and there exists two leaves $a\in A\cap V(T_{x})$ and $b\in B\cap V(T_{x})$ such that, for each child $c$ of $x$, $\left|(A\setminus\{a\})\cap V(T_{c})\right|=\left|(B\setminus\{b\}\cap V(T_{c}))\right|$ and $\{a,b\}\nsubseteq V(T_{c})$.
In particular, $x\in S$ if and only if there exists $a\in A\cap V(T_{x})$ and $b\in B\cap V(T_{x})$ such that, for each child $c$ of $x$, $\left|(A\setminus\{a\})\cap V(T_{c})\right|=\left|(B\setminus\{b\}\cap V(T_{c}))\right|$ and $\{a,b\}\nsubseteq V(T_{c})$.
For an illustration of the different cases see Figure 4 where $x_{1}$ is an example of a node satisfying the first case, $x_{2}$ an example for the second case and $x_{3}$ for the third.
Proof 3.3.
Let $x$ be an inner node. We consider the set $S^{\prime}=S\setminus(V(T_{x})\setminus\{x\})$ of all nodes which are not proper descendants of $x$ and the restrictions $\pi^{\prime}$ and $\sigma^{\prime}$ of, respectively, $\pi$ and $\sigma$ to $S^{\prime}$. By item 2 of 4 $(\pi^{\prime},\sigma^{\prime})$ identify $S^{\prime}$ with unique request. We denote $A^{\prime}=\pi^{\prime}(S^{\prime})$ and $B^{\prime}=\sigma^{\prime}(S^{\prime})$, thus $(A^{\prime},B^{\prime})$ identify $S^{\prime}$. Let $A^{\prime}_{x}=A^{\prime}\cap V(T_{x})$ and $B^{\prime}_{x}=B^{\prime}\cap V(T_{x})$ be the sets of images of nodes in $S^{\prime}$ under $\pi^{\prime}$, $\sigma^{\prime}$, respectively, contained in $T_{x}$. In case $a$ is an element of $A^{\prime}_{x}$, the element $s_{a}=\pi^{-1}(a)$ is an ancestor of $x$ because it is an ancestor of $a\in V(T_{x})$ and belongs to $S^{\prime}$ whence not to $V(T_{x})\setminus\{x\}$. Therefore, $x$ is $s_{a}$-requested. As $(\pi^{\prime},\sigma^{\prime})$ has unique request, there exists at most one element in $A^{\prime}_{x}$. Similarly, $B^{\prime}_{x}$ has size at most $1$. Moreover, $A^{\prime}_{x}\cap B^{\prime}_{x}=\emptyset$ by item 4 of 4. We thus we have three cases:
Case $A^{\prime}_{x}\cup B^{\prime}_{x}=\emptyset$:
Then there is no $s\in S^{\prime}$ such that $\pi(s)\in V(T_{x})$ or $\sigma(s)\in V(T_{x})$. Hence, $x$ is not requested in $(\pi,\sigma)$, in particular, $x\notin S$.
Let $c$ be a child of $x$. If $\left|A\cap V(T_{c})\right|\neq\left|B\cap V(T_{c})\right|$, then there exists $a\in(A\cup B)\cap V(T_{c})$ such that, assuming without loss of generality that $a\in A$ and denoting $s_{a}=\pi^{-1}(a)$, $\sigma(s_{a})\notin V(T_{c})$. Hence, $s_{a}$ is a proper ancestor of $c$, thus, equivalently, an ancestor of $x$, implying that $x$ is $s_{a}$-requested in $(\pi,\sigma)$. This contradicts the above argument. So, $\left|A\cap V(T_{c})\right|=\left|B\cap V(T_{c})\right|$ for each child $c$ of $x$.
Case $A^{\prime}_{x}\cup B^{\prime}_{x}=\{a\}$:
Assume, without loss of generality, that $a\in A$, and denote $s_{a}=\pi^{-1}(a)$ and $b=\sigma(s_{a})$. We have that $s_{a}$ is a proper ancestor of $x$, since it is the least common ancestor of $a\in V(T_{x})$ and $b\notin V(T_{x})$. Thus, $x$ is $s_{a}$-requested and $s_{a}\neq x$ so $x\notin S$.
Let $c$ be a child of $x$. If $\left|A\cap V(T_{c})\right|\neq\left|B\cap V(T_{c})\right|$, then it means that there exists $z\in(A\cup B)\cap V(T_{c})$, such that either $z\in A$ and $s_{z}=\pi^{-1}(z)\notin V(T_{c})$, or $z\in B$ and $s_{z}=\sigma^{-1}(z)\notin V(T_{c})$. In both cases, $s_{z}$ is a proper ancestor of $c$, whence an ancestor of $x$. Thus, $x$ is $s_{z}$-requested in $(\pi,\sigma)$. However, $x$ is $s_{a}$-requested in $(\pi,\sigma)$ which has unique request, hence $s_{z}=s_{a}$. If $z\in B$, it follows that $z=b$, which contradicts the fact $b\notin V(T_{x})$. Hence, $z\in A$ and thus, $z=\pi(s_{a})=a$. On the other hand, if $\left|A\cap V(T_{c})\right|=\left|B\cap V(T_{c})\right|$ then $a\notin V(T_{c})$. If $a\in V(T_{c})$ then $c$ would be $s_{a}$ requested and additionally (as $\left|(A\setminus\{a\}\cap V(T_{c})\right|<\left|B\cap V(T_{c})\right|$) there has to be $z\in T_{c}$, $z\neq a$ for which $s_{z}=\sigma(z)^{-1}$ is an ancestor of $c$. In particular, $s_{a}\neq s_{z}$ as $c$ is the least common ancestor of $a$ and $z$, but $s_{a}$ is a proper ancestor of $c$. Hence, $c$ is requested by both $s_{a}$ and $s_{z}$ contradicting the assumption that $(\pi,\sigma)$ has unique requests. Therefore, $\left|(A\setminus\{a\})\cap V(T_{c})\right|=\left|B\cap V(T_{c})\right|$ for each child $c$ of $x$.
Case $A^{\prime}_{x}=\{a\}$ and $B^{\prime}_{x}=\{b\}$:
Let $s_{a}=\pi^{-1}(a)$ and $s_{b}=\sigma^{-1}(b)$. The node $x$ is $s_{a}$-requested and $s_{b}$-requested in $(\pi,\sigma)$, so, by the unique request property, $s_{a}=s_{b}$. Since $s_{a}$ is the least common ancestor of $a$ and $b$, it belongs to $V(T_{x})$ whence $s_{a}=x$ implying $x\in S$.
Let $c$ be a child of $x$. If $\left|A\cap V(T_{c})\right|\neq\left|B\cap V(T_{c})\right|$, then there exists $z\in(A\cup B)\cap V(T_{c})$ such that either $z\in A$ and $s_{z}=\pi^{-1}(z)\notin V(T_{c})$, or $z\in B$ and $s_{z}=\sigma^{-1}(z)\notin V(T_{c})$. In both cases, $s_{z}$ is a proper ancestor of $c$, whence an ancestor of $x$, and thus $x$ is $s_{z}$-requested in $(\pi,\sigma)$. However, $x$ is $x$-requested in $(\pi,\sigma)$ which has unique request, hence $s_{z}=x$ and thus $z\in\{a,b\}$. On the other hand, similarly to the previous case, if $\left|A\cap V(T_{c})\right|=\left|B\cap V(T_{c})\right|$, then neither $a$ nor $b$ can be contained in $V(T_{c})$. Therefore, $\left|(A\setminus\{a\})\cap V(T_{c})\right|=\left|(B\setminus\{b\})\cap V(T_{c})\right|$ for each child $c$ of $x$. Because the least common ancestor of $a$ and $b$ is $x$, there is no child $c$ of $x$ containing both $a$ and $b$ as leaves, i.e., $\{a,b\}\nsubseteq V(T_{c})$.
This concludes the proof of the statement.
It follows that for each bi-colouring $(A,B)$, there exists at most one set $S$ of inner nodes identified by $(A,B)$.
Lemma 8.
Let $(A,B)$ be two disjoint subsets of $L$ and let $S$ and $S^{\prime}$ be two subsets of $V\setminus L$. If $(A,B)$ identify both $S$ and $S^{\prime}$, then $S=S^{\prime}$.
Proof 3.4.
We proceed by contradiction and thus assume $S\neq S^{\prime}$. Let $s\in S\setminus S^{\prime}$. Since $s\in S$ by 7(3), there are two leaves $a\in A\cap V(T_{c})$ and $b\in B\cap V(T_{c})$ such that, for each child $c$ of $x$, $\left|(A\setminus\{a\})\cap V(T_{c})\right|=\left|(B\setminus\{b\}\cap V(T_{c}))\right|$ and $\{a,b\}\nsubseteq V(T_{c})$. Let $c_{a}$ and $c_{b}$ be the two children of $s$ with $a\in V(T_{c_{a}})$ and $b\in V(T_{c_{b}})$. Since $\{a,b\}\nsubseteq V(T_{c})$ for every child $c$ of $s$, we get $c_{a}\neq c_{b}$. In particular, $\left|A\cap V(T_{c_{a}})\right|=\left|(B\cap V(T_{c_{a}}))\right|+1$, $\left|A\cap V(T_{c_{b}})\right|+1=\left|(B\cap V(T_{c_{b}}))\right|$ and $\left|A\cap V(T_{c})\right|=\left|(B\cap V(T_{c}))\right|$ for every other children $c$ of $s$. Since $A$ and $B$ are disjoint by item 4 of 4, the number of children $c$ of $s$ for which the set $(A\cup B)\cap V(T_{c})$ has odd size is $2$.
On the other hand, since $s\notin S^{\prime}$, 7(1) and (2) imply, using a similar argument to above, that there are no children $c$ of $s$ for which $(A\cup B)\cap V(T_{c})$ has odd size (in case (1)) or there is $1$ child $c$ of $s$ for which $(A\cup B)\cap V(T_{c})$ has odd size. This contradicts our conclusion from the previous paragraph that $s$ should have $2$ such children.
When $(A,B)$ identifies a set $S$, we call $A$-representative (resp. $B$-representative) of $s\in S$ the leaf $\pi(s)\in A$ (resp. $\sigma(s)\in B$), where $(\pi,\sigma)$ witnessing that $(A,B)$ identifies $S$. An example of bi-colouring identifying a subset of inner nodes is given in Figure 4.
Figure 4. Illustration of a bi-colouring $(A,B)$ identifying some set $S\subseteq V\setminus L$ in a binary tree $T$. Leaves from $A$ are blue diamonds, leaves from $B$ are orange filled circles, and inner nodes from $S$ are filled purple squares. Furthermore, the two paths connecting an inner node $s\in S$ to its AA- and $B$-representatives are coloured blue and orange, respectively. Nodes labeled $x_{1},x_{2},x_{3}$ constitute examples illustrating the different cases in 7.
Not every set of inner nodes has a bi-colouring identifying it. To ensure that such a pair exists, we consider thin sets of inner nodes. While thin sets always have bi-colourings identifying them, it is also guaranteed that the set of inner nodes can be partitioned into only 4 thin sets. A subset $X\subseteq V\setminus L$ is thin when, for each $x\in X$ not being the root, on the one hand, the parent $p_{x}$ of $x$ does not belong to $X$, and on the other hand, $x$ admits at least one sibling (including possible leaves) that does not belong to $X$. Having a thin set $X$ allows to find branches avoiding it.
Lemma 9.
Let $X\subseteq V\setminus L$ and $s\in V\setminus X$. If $X$ is thin, then there exists a leaf $t\in V(T_{s})$ such that the path from $t$ to $s$ avoids $X$ (i.e., none of the nodes along this path belong to $X$).
Proof 3.5.
If $T_{s}$ has height 0, then $s$ is a leaf and taking $t=s$ trivially gives the expected path. Otherwise, $s$ is an inner node and, because $X$ is thin, $s$ has at least one child $c_{s}$ not belonging to $X$. By induction, there is a path from some leaf $t\in V(T_{c_{s}})$ to $c_{s}$ avoiding $X$ and, since $s\notin X$, this path could be extended into a path from $t$ to $s$ avoiding $X$.
The following lemma allows to identify every thin set.
Lemma 10.
If $X$ is a thin set, then there exists a pair $(\pi,\sigma)$ of injections from $X$ to $L$ that has unique request and that identifies $X$.
Proof 3.6.
[…]
A family $F=(A_{1},B_{1}),\ldots,(A_{n},B_{n})$ of bi-colourings identifies a set $S\subseteq V\setminus L$, if there exists a partition $(S_{1},\ldots,S_{n})$ of $S$ such that, for each $i\in[n]$, $(A_{i},B_{i})$ identifies $S_{i}$. Whenever $S=V\setminus L$ we say that $F$ identifies $T$. Note that while $A_{i}$ and $B_{i}$ must be disjoint, for $i\not=j$ it is possible (and sometimes even necessary) that $A_{i}$ and $A_{j}$ or $A_{i}$ and $B_{j}$ are not disjoint (see e.g. Figure 5). A collection of subsets of $V\setminus L$ is thin if each of its subsets is thin. We now show that there exists a thin $4$-partition.
Lemma 11.
There exists a thin $4$-partition of $V\setminus L$.
Proof 3.7.
We build such a thin $4$-partition as follows. First, consider the partition $(D_{e},D_{o})$ of $V\setminus L$ in which $D_{e}$ (resp. $D_{o}=V\setminus(D_{e}\cup L)$) is the set of all inner nodes of even (resp. odd) depth. Second, arbitrarily fix one child $c_{x}$ of $x$ for each inner node $x$, and consider the set $C=\{c_{x}\mid x\in V\setminus L\}\setminus L$, inducing a partition $(C,\overline{C})$ of $V\setminus L$ where $\overline{C}=V\setminus(L\cup C)$. By refining these two bi-partitions, we obtain a $4$-partition which is thin by construction.5
Hence, four bi-colourings are enough to identify $T$. For an example see Figure 5.
Corollary 12.
There exists a family of four bi-colourings identifying $T$.
Proof 3.8.
The result immediately follows from 10 and $1$.
Figure 5. Illustration of a partition of the inner nodes into 4 thin sets indicated by different shapes/fillings. The four colourings identifying each thin set are indicated in the table below the leaves where an orange slash indicates the leaf being in the corresponding set $A$ and a blue cross indicates the leaf being in $B$.
3.2. The transduction
The goal of this section is to prove 3, that is, to design a $\operatorname{C}_{2}{\!\!}\operatorname{MSO}$-transduction that produces the laminar tree induced by an input laminar set system. We fix a laminar set system $(U,\mathcal{S})$, represented by a $\{\mathsf{SET}\}$-structure $\mathbb{S}$. Before proving the theorem, we make the following basic observation. On $\mathbb{S}$, we can define two $\operatorname{MSO}$-formulas $\mathsf{desc}(X,Y)$ and $\mathsf{child}(X,Y)$ expressing that in the laminar tree induced by $\mathcal{S}$, $X$ and $Y$ are nodes and $X$ is a descendant or a child of $Y$, respectively:
$\mathsf{desc}(X,Y)$ $:=\mathsf{SET}(X)\land\mathsf{SET}(Y)\land X\subseteq Y;$
$\mathsf{child}(X,Y)$ $:=\mathsf{desc}(X,Y)\land X\neq Y\land\forall Z\>\Big{(}\big{(}\mathsf{desc}(Z,Y)\land Z\neq Y\big{)}\rightarrow\mathsf{desc}(Z,X)\Big{)}.$
The key point of our construction consists in defining a $\operatorname{C}_{2}{\!\!}\operatorname{MSO}$-formula ${\mathsf{repr}}_{A,B}(a,X)$ which, assuming a bi-colouring $(A,B)$ of the universe modeled as disjoint unary relations and identifying a subset $S$ of inner nodes of $T$, is satisfied exactly when $X\in S$ and $a$ is its $A$-representative.
Lemma 13.
Let $(A,B)$ be a bi-colouring of $L(T)$ identifying a subset $S$ of inner nodes of $T$. There exists a $\operatorname{C}_{2}{\!\!}\operatorname{MSO}$-formula ${\mathsf{repr}}_{A,B}(a,X)$ that is satisfied exactly when $X$ is an inner node of $T$ that belongs to $S$ and is $A$-represented by $a$.
Proof 3.9.
[…]
$\phi_{A,B}(X):=\>$ $\mathsf{SET}(X)\land\exists a\exists b\Big{[}a\in(X\cap A)\land b\in(X\cap B)\land$
$\qquad\forall Z\>\Big{(}\mathsf{child}(Z,X)\rightarrow\Big{(}\{a,b\}\nsubseteq Z\land\operatorname{C}_{2}\big{(}(Z\setminus\{a,b\})\cap(A\cup B)\big{)}\Big{)}\Big{)}\Big{]}.$
Now, we can easily define ${\mathsf{repr}}_{A,B}(a,X)$ based on 5:
${\mathsf{repr}}_{A,B}(a,X):=\>$ $\phi_{A,B}(X)\land a\in(X\cap A)\land\forall Z\subsetneq X\>\Big{(}a\in Z\rightarrow\neg\phi_{A,B}(Z)\Big{)}.$
This concludes the proof.
We are now ready to prove the theorem.
Proof 3.10 (Proof of 3).
The $\operatorname{C}_{2}{\!\!}\operatorname{MSO}$-transduction is obtained by composing the following atomic $\operatorname{C}_{2}{\!\!}\operatorname{MSO}$-transductions. The transduction makes use of the formulas ${\mathsf{repr}}_{A,B}(a,X)$ given by 13.
  1. (1)
    Guess a family of four bi-colourings $(A_{i},B_{i})_{i\in[4]}$ identifying $T$ (which exists by 12).
  2. (2)
    Copy the input graph four times, thus introducing four binary relations $(\mathsf{copy}_{i})_{i\in[4]}$ where $\mathsf{copy}_{i}(x,y)$ indicates that $x$ is the $i$-th copy of the original element $y$.
  3. (3)
    Filter the universe keeping only the original elements as well as the $i$-th copy of each vertex $a$ for which there exists $X$ such that ${\mathsf{repr}}_{A_{i},B_{i}}(a,X)$.
  4. (4)
    Use an interpretation adding relation $\mathsf{ancestor}$ and keeping every other relation in $\Sigma$ unchanged outputting $\Sigma\cup\{\mathsf{ancestor}\}$-structures as follows. Define the relation $\mathsf{ancestor}(x,y)$ so that it is satisfied exactly when there exist $x^{\prime}$, $X$, $i$, and $Y$ such that, on the one hand $\mathsf{desc}(Y,X)$ and $\mathsf{copy}_{i}(x,x^{\prime})\land{\mathsf{repr}}_{A_{i},B_{i}}(x^{\prime},X)$, and, on the other hand, either $y$ is an original element and $Y=\{y\}$, or there exists $y^{\prime}$ and $j$ such that $\mathsf{copy}_{j}(y,y^{\prime})\land{\mathsf{repr}}_{A_{j},B_{j}}(y^{\prime},Y)$.
  5. (5)
    Use a filtering transduction keeping only those outputs $\mathbb{B}$ from the previous steps that satisfy that for every set $X$ in $\mathsf{SET}_{\mathbb{B}}$ there is an element $t$ in $U_{\mathbb{B}}$ such that the set of descendants of $t$ is equal to $X$ and for every element $t$ of $U_{\mathbb{B}}$ the set of descendants of $t$ is in $\mathsf{SET}_{\mathbb{B}}$ and hence verifying that $\mathsf{ancestor}_{\mathbb{B}}$ defines a laminar tree of $(U,\mathcal{S})$. Clearly, this can be expressed using a $\operatorname{MSO}$-sentence and hence outputs are of the form $\mathbb{A}\sqcup\mathbb{T}$ for some $\{\mathsf{ancestor}\}$-structure $\mathbb{T}$ representing the laminar tree $T$ of $(U,\mathcal{S})$.
4. Transducing modular decompositions
The set of modules of a directed graph is a specific example of a particular type of set system, a “weakly-partitive set system” (and a “partitive set system” in the case of an undirected graph). In this section, we first give a general $\operatorname{C}_{2}{\!\!}\operatorname{MSO}$-transduction to obtain the canonical tree-like decomposition of a weakly-partitive set system from the set system itself. We then show how to obtain the modular decomposition of a graph via a $\operatorname{C}_{2}{\!\!}\operatorname{MSO}$-transduction as an application.
4.1. Transducing weakly-partitive trees
A set system $(U,\mathcal{S})$ is weakly-partitive if for every two overlapping sets $X,Y\in\mathcal{S}$, the sets $X\cup Y$, $X\cap Y$, $X\setminus Y$, and $Y\setminus X$ belong to $\mathcal{S}$. It is partitive if, moreover, for every two overlapping sets $X,Y\in\mathcal{S}$, their symmetric difference, denoted $X\operatorname*{\bigtriangleup}Y$, also belongs to $\mathcal{S}$. By extension, a set family $\mathcal{S}$ is called weakly-partitive or partitive whenever $\left(\bigcup\mathcal{S},\mathcal{S}\right)$ is a set system which is weakly-partitive or partitive, respectively (note these also requires that $\emptyset\notin\mathcal{S}$, $\bigcup\mathcal{S}\in\mathcal{S}$, and $\{a\}\in\mathcal{S}$ for every $a\in\bigcup\mathcal{S}$).
Figure 6. Sets to be included in a weakly-partitive set system (excluds the rightmost) or in a partitive set system (includes the rightmost) for two overlapping sets $X$ and $Y$.
A member of a set system $(U,\mathcal{S})$ is said to be strong if it does not overlap any other set from $\mathcal{S}$. The sub-family $\mathcal{S}_{!}$ of strong sets of $\mathcal{S}$ is thus laminar by definition. Hence, it induces a laminar tree $T$. By extension, we say that $T$ is induced by the set system $(U,\mathcal{S})$ (or simply by $\mathcal{S}$). The next result extends 3, by showing that $T$ can be $\operatorname{C}_{2}{\!\!}\operatorname{MSO}$-transduced from $\mathcal{S}$.
Lemma 14.
Let $\Sigma$ be an extended vocabulary, including a unary set predicate name $\mathsf{SET}$ and not including the binary relational symbol $\mathsf{ancestor}$. There exists a non-deterministic overlay $\operatorname{C}_{2}{\!\!}\operatorname{MSO}$-transduction $\tau$ such that, for each set system $(U,\mathcal{S})$ represented as the $\{\mathsf{SET}\}$-structure $\mathbb{S}$ and inducing the laminar tree $T$ with L(T)=U{L(T)=U}, and for each $\Sigma$-structure $\mathbb{A}$ with $\mathbb{S}\sqsubseteq\mathbb{A}$, $\tau(\mathbb{A})$ is non-empty and every output in $\tau(\mathbb{A})$ is equal to $\mathbb{A}\sqcup\mathbb{T}$ for some $\{\mathsf{ancestor}\}$-structure $\mathbb{T}$ representing $T$.
Proof 4.1.
On the $\{\mathsf{SET}\}$-structure $\mathbb{S}$, it is routine to define an $\operatorname{MSO}$-formula $\phi_{\mathsf{SET!}}(Z)$ that identifies those subsets $Z\subseteq U$ that are strong members of $\mathcal{S}$:
$\phi_{\mathsf{SET!}}(Z)$ $:=\mathsf{SET}(Z)\land\forall X\>\Big{(}\big{(}\mathsf{SET}(X)\land X\cap Z\neq\emptyset\big{)}\rightarrow\big{(}X\subseteq Z\lor Z\subseteq X\big{)}\Big{)}.$
Hence, we can design an $\operatorname{MSO}$-interpretation that outputs the $\Sigma\cup\{\mathsf{SET!}\}$-structure corresponding to $\mathbb{A}$ equipped with the set unary predicate $\mathsf{SET!}$ that selects strong members of $\mathcal{S}$. Thus, up to renaming the predicates $\mathsf{SET!}$ and $\mathsf{SET}$, by 3, we can produce, through a $\operatorname{C}_{2}{\!\!}\operatorname{MSO}$-transduction, the $\Sigma\cup\{\mathsf{ancestor}\}$-structure $\mathbb{A}\sqcup\mathbb{T}$ where $\mathbb{T}$ is the tree-structure modeling the laminar tree $T$ induced by $\mathcal{S}_{!}$ with $L(T)=U$ (once the output obtained, the interpretation drops the set predicate $\mathsf{SET!}$ which is no longer needed).
Clearly, the laminar tree $T$ of a weakly-partitive set system $(U,\mathcal{S})$ does not characterize $(U,\mathcal{S})$. However, as shown by the below theorem, a labeling of its inner nodes and a controlled partial ordering of its nodes are sufficient to characterize all the sets of $\mathcal{S}$. For $Z$ a set equipped with a partial order << and $X$ a subset of $Z$, we say that $X$ is a <<-interval whenever << defines a total order on $X$ and for every $a,b\in X$ and every $c\in Z$, $a<c<b$ implies $c\in X$.{thmC}[[CHM81, Rao06]] Let $\mathcal{S}$ be a weakly-partitive family, $\mathcal{S}_{!}$ be its subfamily of strong sets, and $T$ be the laminar tree it induces. There exists a total labeling function $\lambda$ from the set $V(T)\setminus L(T)$ of inner nodes of $T$ to the set $\{\mathsf{degenerate},\mathsf{prime},\mathsf{linear}\}$, and, for each inner node $t\in\lambda^{-1}(\mathsf{linear})$, a linear ordering $_{t}$ of its children, such that every inner node having exactly two children is labeled by $\mathsf{degenerate}$ and the following two conditions are satisfied:
  • for each $X\in\mathcal{S}\setminus\mathcal{S}_{!}$, there exists $t\in V(T)$ and a subset $\mathcal{C}$ of children of $t$ such that $X=\bigcup_{c\in\mathcal{C}}L(T_{c})$ and either $\lambda(t)=\mathsf{linear}$ and $\mathcal{C}$ is a $_{t}-$interval, or $\lambda(t)=\mathsf{degenerate}$;
  • conversely, for each inner node $t$ and each non-empty subset $\mathcal{C}$ of children of $t$, if either $\lambda(t)=\mathsf{linear}$ and $\mathcal{C}$ is a $_{t}-$interval, or $\lambda(t)=\mathsf{degenerate}$, then $\bigcup_{c\in\mathcal{C}}L(T_{c})\in\mathcal{S}$.
Furthermore, $T$ and $\lambda$ are uniquely determined from $\mathcal{S}$, and, for each inner node $t$ of $T$ labeled by $\mathsf{linear}$, only two orders $_{t}$ are possible, one being the inverse of the other (indeed, inverting an order << does preserve the property of being a <<-interval). Hence, every weakly-partitive family $\mathcal{S}$ is characterized by a labeled and partially-ordered tree $(T,\lambda,<)$ where $T$ is the laminar tree induced by the subfamily $\mathcal{S}_{!}$ of strong sets of $\mathcal{S}$, $\lambda:V(T)\setminus L(T)\to\{\mathsf{degenerate},\mathsf{prime},\mathsf{linear}\}$ is the labeling function, and << is the partial order $\bigcup_{t\in\lambda^{-1}(\mathsf{linear})}<_{t}$ over $V(T)$. As, up-to inverting some of the $_{t}$ orders, $(T,\lambda,<)$ is unique, we abusively call it the weakly-partitive tree induced by $\mathcal{S}$. Conversely, a weakly-partitive tree characterizes the unique weakly-partitive set system which induced it.
We naturally model a weakly-partitive tree $(T,\lambda,<)$ of a weakly-partitive set system $(U,\mathcal{S})$ by the $\{\mathsf{ancestor},\mathsf{degenerate},\mathsf{betweeness}\}$-structure $\mathbb{T}$ of universe $U_{\mathbb{T}}=V(T)$ such that $\langle V(T),\mathsf{ancestor}_{\mathbb{T}}\rangle\sqsubseteq\mathbb{T}$ models $T$ with L(T)=U{L(T)=U}, $\mathsf{degenerate}_{\mathbb{T}}$ is a unary relation which selects the inner nodes of $T$ of label $\mathsf{degenerate}$, i.e., $\mathsf{degenerate}_{\mathbb{T}}=\lambda^{-1}(\mathsf{degenerate})$}, and $\mathsf{betweeness}_{\mathbb{T}}$ is a ternary relation selecting triples $(x,y,z)$ satisfying $x<y<z$ or $z<y<x$. (Although it is possible, through a non-deterministic $\operatorname{MSO}$-transduction, to define << from $\mathsf{betweeness}_{\mathbb{T}}$, the use of $\mathsf{betweeness}_{\mathbb{T}}$ rather than $_{\mathbb{T}}$ ensures uniqueness of the output weakly-partitive tree.) The inner nodes of $T$ that are labeled by $\mathsf{linear}$ can be recovered through an $\operatorname{MSO}$-formula as those inner nodes whose children are related by $\mathsf{betweeness}$. The inner nodes labeled by $\mathsf{prime}$ can be recovered through an $\operatorname{MSO}$-formula as those inner nodes that are labeled neither by $\mathsf{degenerate}$ nor by $\mathsf{linear}$. Using 4.1, it is routine to design an $\operatorname{MSO}$-transduction which takes as input a weakly-partitive tree and outputs the weakly-partitive set system which induced it. The inverse $\operatorname{C}_{2}{\!\!}\operatorname{MSO}$-transduction is the purpose of the next result.
Theorem 15.
There exists a non-deterministic $\operatorname{C}_{2}{\!\!}\operatorname{MSO}$-transduction $\tau$ such that, for every weakly-partitive set system $(U,\mathcal{S})$ represented as the $\{\mathsf{SET}\}$-structure $\mathbb{S}$ and inducing the weakly-partitive tree $(T,\lambda,<)$ represented as the $\{\mathsf{ancestor},\mathsf{degenerate},\mathsf{betweeness}\}$-structure $\mathbb{T}$, we have $\mathbb{T}\in\tau(\mathbb{S})$ and every output in $\tau(\mathbb{S})$ is a weakly-partitive tree of $(U,\mathcal{S})$.
Proof 4.2.
By 14, we have a $\operatorname{C}_{2}{\!\!}\operatorname{MSO}$-transduction which on inputs $\mathbb{S}$ outputs the $\{\mathsf{SET},\mathsf{ancestor}\}$-structure $\mathbb{S}\sqcup\mathbb{T^{\prime}}$, where $\mathbb{T^{\prime}}$ is the $\{\mathsf{ancestor}\}$-structure modeling $T$, with $L(T)=U=U_{\mathbb{S}}$. It thus remains to define a second $\operatorname{C}_{2}{\!\!}\operatorname{MSO}$-transduction which outputs $\mathbb{T}$ from $\mathbb{S}\sqcup\mathbb{T^{\prime}}$. Actually, an $\operatorname{MSO}$-interpretation is enough. Indeed, we can define an $\operatorname{MSO}$-formula $\phi_{\mathsf{degenerate}}$ which selects the inner nodes of $T$ that are labeled $\mathsf{degenerate}$ by $\lambda$, and an $\operatorname{MSO}$-formula $\phi_{\mathsf{betweeness}}$ which selects triples $(x,y,z)$ of elements such that $x<y<z$ or $z<y<x$. In order to define both formulas, we use a functional symbol $\mathsf{leafset}(t)$ which is interpreted as the set $L(T_{t})$, and is clearly $\operatorname{MSO}$-definable on $\{\mathsf{ancestor}\}$-structures such as $\mathbb{T}$. Indeed, an element $z$ belongs to $\mathsf{leafset}(t)$ if and only if it is a leaf and has $t$ as ancestor.
According to 4.1, an inner node $t$ is labeled $\mathsf{degenerate}$ if and only if, for every two children $s_{1},s_{2}$ of $t$ the set $L(T_{s_{1}})\cup L(T_{s_{2}})$ is in $\mathcal{S}$. We can define this in $\operatorname{MSO}$ using the following formula:
$\phi_{\mathsf{degenerate}}(x):=\forall y\forall z\Big{(}\big{(}\mathsf{parent}(x,y)\land\mathsf{parent}(x,z)\big{)}\rightarrow\mathsf{SET}\big{(}\mathsf{leafset}(y)\cup\mathsf{leafset}(z)\big{)}\Big{)}.$
Now, according to 4.1, we have $x<y<z$ or $z<y<x$ exactly when $x$, $y$, and $z$ are three distinct children of some node $t$ and the three following properties are satisfied:
  1. (1)
    $L(T_{x})\cup L(T_{z})$ is not a member of $\mathcal{S}$;
  2. (2)
    there exists a member of $\mathcal{S}$ which includes both $L(T_{x})$ and $L(T_{y})$ but excludes $L(T_{z})$;
  3. (3)
    there exists a member of $\mathcal{S}$ which includes both $L(T_{y})$ and $L(T_{z})$ but excludes $L(T_{x})$.
Indeed, when these conditions are satisfied, $t$ is necessarily labeled $\mathsf{linear}$ (item 1 asserts that it is not labeled by $\mathsf{degenerate}$ and, e.g., item 2 that it is not labeled by $\mathsf{prime}$), and, among the possible orderings of $\{x,y,z\}$, only the ones ensuring $x<y<z$ or $z<y<x$ are possible ($x<z<y$ and $y<z<x$ are not possible because of item 2, and $y<x<z$ and $z<x<y$ are not possible because of item 3). Each of these properties is $\operatorname{MSO}$-definable:
$\phi_{\mathsf{betweeness}}(x,y,z):=$ $\exists t\Big{[}\mathsf{parent}(t,x)\land\mathsf{parent}(t,y)\land\mathsf{parent}(t,z)\land$
$\neg\mathsf{SET}\big{(}\mathsf{leafset}(x)\cup\mathsf{leafset}(z)\big{)}\land$
$\exists S\Big{(}\mathsf{SET}(S)\land\big{(}\mathsf{leafset}(x)\cup\mathsf{leafset}(y)\big{)}\subseteq S\land\mathsf{leafset}(z)\cap S=\emptyset\Big{)}\land$
$\exists S\Big{(}\mathsf{SET}(S)\land\big{(}\mathsf{leafset}(z)\cup\mathsf{leafset}(y)\big{)}\subseteq S\land\mathsf{leafset}(x)\cap S=\emptyset\Big{)}\Big{]}.$
Once $\phi_{\mathsf{degenerate}}$ and $\phi_{\mathsf{betweeness}}$ are defined, our transduction drops the original predicate $\mathsf{SET}$ which is no longer needed in the output $\mathbb{T}$.
If $\mathcal{S}$ is partitive then the weakly-partitive tree it induces enjoys a simple form, and is unique. Indeed, the label $\mathsf{linear}$ and, thus, the partial order <<, are not needed.{thmC}[[CHM81]] Let $\mathcal{S}$ be a weakly-partitive family and $(T,\lambda,<)$ be the weakly-partitive tree it induces. If $\mathcal{S}$ is partitive, then $\lambda^{-1}(\mathsf{linear})=\emptyset$ and << is empty. Hence, in case of a partitive set systems $(U,\mathcal{S})$, we can consider the simpler object $(T,\lambda)$, called the partitive tree induced by $\mathcal{S}$ (or the partitive tree of $(U,\mathcal{S})$) in which $\lambda$ maps $V(T)\setminus L(T)$ to $\{\mathsf{degenerate},\mathsf{prime}\}$. As a direct consequence of 4.2 and of 15, we can produce, through a $\operatorname{C}_{2}{\!\!}\operatorname{MSO}$-transduction, the partitive tree induced by a partitive set system and naturally modeled by an $\{\mathsf{ancestor},\mathsf{degenerate}\}$-structure.
Corollary 16.
There exists a non-deterministic $\operatorname{C}_{2}{\!\!}\operatorname{MSO}$-transduction $\tau$ such that, for each partitive set system $(U,\mathcal{S})$ represented as the $\{\mathsf{SET}\}$-structure $\mathbb{S}$ and inducing the partitive tree $(T,\lambda)$ represented as the $\{\mathsf{ancestor},\mathsf{degenerate}\}$-structure $\mathbb{T}$, we have $\mathbb{T}\in\tau(\mathbb{S})$ and every output in $\tau(\mathbb{S})$ is a partitive tree of $(U,\mathcal{S})$.
4.2. Application to modular decomposition
Let $G$ be a directed graph and let $M\subseteq V(G)$. We say that $M$ is a module (of $G$) if for every $u\notin M$ and every $v,w\in M$, $uv\in E(G)\!\iff\!uw\in E(G)$} and $vu\in E(G)\!\iff\!wu\in E(G)$}. Clearly, the empty set, $V(G)$, and all the singletons $\{x\}$ for $x\in V(G)$ are modules; they are called the trivial modules of $G$. We say a non-empty module $M$ is maximal if it is not properly contained in any non-trivial module. Furthermore, we use the notion of strong modules to coincide with the strong sets in the set system consisting of all modules of $G$. Let $M$ and $M^{\prime}$ be two disjoint non-empty modules of $G$. Considering the edges that go from $M$ to $M^{\prime}$, namely edges from the set $(M\times M^{\prime})\cap E(G)$, we have two possibilities: either it is empty, or it is equal to $M\times M^{\prime}$. We write $M\text{$\not\to$}M^{\prime}$ in the former case and $M\text{$\to$}M^{\prime}$ in the latter. (It is of course possible to have both $M\text{$\to$}M^{\prime}$ and $M^{\prime}\text{$\to$}M$.) A modular partition of $G$ is a partition $\mathcal{P}=\{M_{1},\ldots,M_{\ell}\}$ of $V(G)$ such that every $M_{i}$ is a non-empty module. A modular partition $\mathcal{P}=\{M_{1},\dots,M_{\ell}\}$ is called maximal if it is non-trivial and every $M_{i}$ is strong and maximal. Note that every graph has exactly one maximal modular partition.
A modular decomposition of $G$ is a rooted tree $T$ in which the leaves are the vertices of $G$, and for each inner node $\in T, tt$ has at least two children and the set $L(T_{t})$ is a module of $G$. In a modular decomposition $T$ of $G$, for each inner node $t\in V(T)$ with children $c_{1},\ldots,c_{r}$, the family $\mathcal{P}_{t}=\{L(T_{c_{1}}),\ldots,L(T_{c_{r}})\}$ is a modular partition of $G[L(T_{t})]$. When each such partition is maximal, the decomposition is unique and it is called the maximal modular decomposition of $G$. The maximal modular decomposition $T$ of $G$ alone is not sufficient to characterize $G$. However, enriching $T$ with, for each inner node $t$ with children $c_{1},\ldots,c_{j}$, the information of which pair of modules $\big{(}L(T_{c_{i}}),L(T_{c_{j}})\big{)}$ is such that $L(T_{c_{i}})\text{$\to$}L(T_{c_{j}})$, yields a unique canonical representation of $G$. Formally, the enriched modular decomposition of $G$ is the pair $(T,F)$ where $T$ is the maximal modular decomposition of $G$ (with $L(T)=V(G)$) and $F\subset V(T)\times V(T)$ is a binary relation, that relates a pair $(s,t)$ of nodes of $T$, denoted $st\in F$, exactly when $s$ and $t$ are siblings and $L(T_{s})\text{$\to$}L(T_{t})$. The elements of $F$ are called $\mathsf{m}$-edges.
It should be mentioned that the family of all non-empty modules of $G$ is known to be weakly-partitive (or even partitive when $G$ is undirected). In particular, the maximal modular decomposition $T$ of $G$ is the laminar tree induced by the family of strong modules. Hence, 15 could be used to produce a partially-ordered and labeled tree which displays all the modules of $G$. However, this weakly-partitive tree is not sufficient for being able to recover the graph $G$ from it. We now prove how to obtain the enriched modular decomposition of $G$ through a $\operatorname{C}_{2}{\!\!}\operatorname{MSO}$-transduction.
To model enriched modular decompositions as relational structures we use the relational vocabulary $\{\mathsf{ancestor},\textsf{m}\text{-}\mathsf{edge}\}$ where $\mathsf{ancestor}$ and $\textsf{m}\text{-}\mathsf{edge}$ are two binary relation names. An enriched modular decomposition $(T,F)$ of a graph $G$ is modeled by the $\{\mathsf{ancestor},\textsf{m}\text{-}\mathsf{edge}\}$-structure $\mathbb{M}$ with universe $U_{\mathbb{M}}=V(T)$, $\mathsf{ancestor}_{\mathbb{M}}$ being the set of pairs $(s,t)$ for which $s$ is an ancestor of $t$ in $T$, and $\textsf{m}\text{-}\mathsf{edge}_{\mathbb{M}}$ being the set of all pairs $(s,t)$ such that $st\in F$ (in particular, $s$ and $t$ are siblings in $T$). We use 14 in order to transduce the maximal modular decomposition of a graph.
Theorem 17.
There exists a non-deterministic $\operatorname{C}_{2}{\!\!}\operatorname{MSO}$-transduction $\tau$ such that for every directed graph $G$ represented as the $\{\mathsf{edge}\}$-structure $\mathbb{G}$, $\tau(\mathbb{G})$ is non-empty and every output in $\tau(\mathbb{G})$ is equal to some $\{\mathsf{ancestor},\textsf{m}\text{-}\mathsf{edge}\}$-structure $\mathbb{M}$ representing the enriched modular decomposition $(T,F)$ of $G$.
Proof 4.3.
Let $G$ be a graph represented by the $\{\mathsf{edge}\}$-structure $\mathbb{G}$. Several objects are associated to $G$, and each of them can be described by a structure:
  • let $\mathcal{M}$ be the family of non-empty modules of $G$ and let $\mathbb{S}$ be the $\{\mathsf{SET}\}$-structure modeling the weakly-partitive set system $(V(G),\mathcal{M})$ with $U_{\mathbb{S}}=V(G)$;
  • let $T$ be the laminar tree induced by the weakly-partitive family $\mathcal{M}$ and let $\mathbb{T}$ be the $\{\mathsf{ancestor}\}$-structure modeling it with $U_{\mathbb{T}}=V(T)$ and $L(T)=V(G)\subset U_{\mathbb{T}}$;
  • let $F$ be the $\mathsf{m}$-edge relation, namely the subset of $V(T)\times V(T)$ such that $(T,F)$ is the maximal modular decomposition of $G$, and let $\mathbb{M}$ be the $\{\mathsf{ancestor},\textsf{m}\text{-}\mathsf{edge}\}$-structure modeling it with $\mathbb{T}\sqsubset\mathbb{M}$ and $U_{\mathbb{T}}=U_{\mathbb{M}}$.
Our $\operatorname{C}_{2}{\!\!}\operatorname{MSO}$-transduction is obtained by composing the three following transductions:
  • $\tau_{1}$: an $\operatorname{MSO}$-interpretation which outputs the $\{\mathsf{edge},\mathsf{SET}\}$-structure $\mathbb{G}\sqcup\mathbb{S}$ from $\mathbb{G}$;
  • $\tau_{2}$: the non-deterministic $\operatorname{C}_{2}{\!\!}\operatorname{MSO}$-transduction given by 14 which produces the $\{\mathsf{edge},\mathsf{SET},\mathsf{ancestor}\}$-structure $\mathbb{G}\sqcup\mathbb{S}\sqcup\mathbb{T}$ from $\mathbb{G}\sqcup\mathbb{S}$;
  • $\tau_{3}$: an $\operatorname{MSO}$-interpretation which outputs the $\{\mathsf{ancestor},\textsf{m}\text{-}\mathsf{edge}\}$-structure $\mathbb{M}$ from $\mathbb{G}\sqcup\mathbb{S}\sqcup\mathbb{T}$.
In order to define $\tau_{1}$ it is sufficient to observe that there exists an $\operatorname{MSO}$-formula $\phi_{\mathsf{SET}}(Z)$ with one monadic free-variable, which is satisfied exactly when $Z$ is a non-empty module of $G$. Then, since $\tau_{2}$ is given by 14, it only remains to define $\tau_{3}$. Given an inner node $t$, we can select, within $\operatorname{MSO}$, the set $L(T_{t})$ of leaves of the subtree rooted in $t$. We thus assume a function $\mathsf{leafset}$, with one first-order free-variable which returns the set of leaves of the subtree rooted at the given node. Equipped with this function, we can define the $\operatorname{MSO}$-formula $\phi_{\textsf{m}\text{-}\mathsf{edge}}$ which selects pairs $(s,r)$ of siblings such that $sr\in F$. Remember that this happen exactly when there exist $u\in L(T_{s})$ and $v\in L(T_{r})$ such that $uv\in E(G)$. Hence, $\phi_{\textsf{m}\text{-}\mathsf{edge}}$ could be defined as:
$\phi_{\textsf{m}\text{-}\mathsf{edge}}(s,r):=s\neq r\>\land\>$ $\exists t\>\big{(}\mathsf{parent}(t,s)\land\mathsf{parent}(t,r)\big{)}\land$
$\exists x\exists y\big{(}x\in\mathsf{leafset}(s)\land y\in\mathsf{leafset}(r)\land\mathsf{edge}(x,y)\big{)}.$
Once defined, the $\operatorname{MSO}$-interpretation $\tau_{3}$ simply drops all non-necessary relations and predicates (namely $\mathsf{edge}$ and $\mathsf{SET}$) and keeps only the $\mathsf{ancestor}$ and $\textsf{m}\text{-}\mathsf{edge}$ relations.
Notice that it is routine to design a deterministic $\operatorname{MSO}$-transduction which, given an $\{\mathsf{ancestor},\textsf{m}\text{-}\mathsf{edge}\}$-structure $\mathbb{M}$ representing an enriched modular decomposition of some directed graph $G$, produces the $\{\mathsf{edge}\}$-structure $\mathbb{G}$ representing $G$.
In the following we provide an example of a natural property that can be $\operatorname{C}_{2}{\!\!}\operatorname{MSO}$-defined easily using the modular decomposition and hence (though backwards translation) is $\operatorname{C}_{2}{\!\!}\operatorname{MSO}$-definable on graphs. {exa} We want to define the property of an undirected graph $G$ having an even number of modules in $\operatorname{C}_{2}{\!\!}\operatorname{MSO}$. Let $T$ be the modular decomposition of $G$ and $\mathcal{M}$ the set of modules of $G$. As $\mathcal{M}$ is partitive, we can understand $T$ as a partitive tree and let $\lambda:V(T)\rightarrow\{\mathsf{prime},\mathsf{degenerate}\}$ be the labeling of $T$ that exists due to 4.1 and 4.2. By 5.2 it holds that the number of modules of $G[L(T_{t})]$ for a node $t$ with children $s_{1},\dots,s_{\ell}$ is exactly the sum of modules of the $G[L(T_{s_{i}})]$ if $\lambda(t)=\mathsf{prime}$ and the product of the $G[L(T_{s_{i}})]$ if $\lambda(t)=\mathsf{degenerate}$. Hence, given a set $X$ of nodes of $T$ we can check whether $X$ is precisely the set of nodes $t$ for which $G[L(T_{t})]$ has an even number of modules in $\operatorname{C}_{2}{\!\!}\operatorname{MSO}$. For this we check for each node $t$ their labeling and either the parity of children of $t$ that are in $X$ (if $\lambda(t)=\mathsf{prime}$) or the existence of a child in $X$ (if $\lambda(t)=\mathsf{degenerate}$) and can conclude whether $t$ should be included in $X$ or not. Note that additionally leaves of $T$ are not included in $X$ as a graph with one vertex only contains one module. The number of modules of $G$ is even if and only if there is a set $X$ which consist of all nodes $t$ for which $G[L(T_{t})]$ is even and the root of $T$ is in $X$.
To define the sentence, we let $\mathbb{T}$ be the $\{\mathsf{ancestor},\mathsf{prime},\mathsf{degenerate}\}$-structure modeling $T$ equipped with $\lambda$. While we did not provide a transduction producing this particular structure, it is easy to modify the above transductions to obtain the desired structure. We further use auxiliary predicates $\mathsf{children}(Y,y)$ expressing that $Y$ is the set of children of $y$, $\mathsf{root}(y)$ expressing that $y$ is the root of $T$ and $\mathsf{leaf}(y)$ expressing that $y$ is a leaf of $T$, which are routine to implement. Hence, we can define a sentence $\phi$ that is satisfied by any graph that has an even number of modules as follows.
$\phi:=$ $\exists X\forall y\exists Y\Bigg{(}\mathsf{children}(Y,y)\land\Big{[}\mathsf{root}(y)\rightarrow y\in X\Big{]}\land\Big{[}\mathsf{leaf}(y)\rightarrow y\notin X\Big{]}\land$
$\Big{[}\Big{(}\big{(}\mathsf{prime}(y)\land\operatorname{C}_{2}(X\cap Y)\big{)}\lor\big{(}\mathsf{degenerate}(y)\land X\cap Y\neq\emptyset\big{)}\Big{)}\leftrightarrow y\in X\Big{]}.$
Cographs
Let $G$ be a graph, $(T,F)$ be its modular decomposition, and $(T,\lambda,<)$ be the weakly-partitive tree induced by the (weakly-partitive) family of its modules. Let $t$ be an inner node of $T$, let $\mathcal{C}$ be its set of children, and let $C$ be the graph $\big{(}V(T)\setminus L(T),F\big{)}\left[\mathcal{C}\right]$ induced by $F$ on the set of children of $t$. It can be checked that, if $\lambda(t)=\mathsf{degenerate}$ then $C$ is either a clique or an independant, and if $\lambda(t)=\mathsf{linear}$ then $C$ is a tournament consistent with $_{t}$ (i.e., for every $\in\mathcal{C}, xyxy$ is an edge of $C$ if and only if $x<_{t}y$) or with the inverse of $_{t}$. In the former case, we can refine the $\mathsf{degenerate}$ label into $\mathsf{series}$ and $\mathsf{parallel}$ labels, thus expressing that $C$ is a clique or an independent set, respectively. In the latter case, up-to reversing $_{t}$, we can ensure that $C$ is a tournament consistent with $_{t}$. This yields a refined weakly-partitive tree $(T,\gamma,<)$, where $\gamma$ maps inner nodes to $\{\mathsf{series},\mathsf{parallel},\mathsf{prime},\mathsf{linear}\}$ and << is the order $\bigcup_{t\in\gamma^{-1}(\mathsf{linear})}<_{t}$ which ensures that tournaments are consistent with the corresponding $_{t}$. Notice that this labeled and partially-ordered tree is now uniquely determined from $G$. Moreover, edges from $F$ that connect children of a node not labeled by $\mathsf{prime}$ can be recovered from the so-refined weakly-partitive tree. In particular, if no nodes of $T$ is labeled by $\mathsf{prime}$, $(T,F)$ and thus $G$ is fully characterized by $(T,\gamma,<)$. Graphs for which this property holds are known as directed cographs. Directed cographs equivalently can be defined by a finite set of forbidden subgraphs (see [CP06]) and in the undirected case, cographs are exactly the $P_{4}$-free graphs. Directed cographs can be described by the refined weakly-partitive tree, called cotree, explained above and formalized in the following statement.
Theorem 18.
Let $G$ be a directed cograph and let $T$ be the laminar tree induced by the family of its strong modules. There exists a unique total labeling $\lambda$ from the set $V(T)\setminus L(T)$ of inner nodes of $T$ to the set $\{\mathsf{series},\mathsf{parallel},\mathsf{linear}\}$ of labels, and, for each inner node $t\in\lambda^{-1}(\mathsf{linear})$, a unique linear ordering $_{t}$ of its children, such that every inner node having exactly two children is labeled $\mathsf{series}$ or $\mathsf{parallel}$, and the following condition is satisfied:
  • for every two leaves $x$ and $y$ of $T$, denoting by $t$ their least common ancestor and $s_{x},s_{y}$ the children of $t$ that are ancestors of $x,y$ respectively, $xy$ is an edge of $G$ if and only if either $t$ is labeled by $\mathsf{linear}$ and $s_{x}<_{t}s_{y}$, or $t$ is labeled by $\mathsf{series}$.
We naturally model cotrees as $\{\mathsf{ancestor},\mathsf{series},\mathsf{ord}\}$-structures as follow. A cotree $(T,\gamma,<)$ is modeled by $\mathbb{C}$ where $U_{\mathbb{C}}=V(T)$, $\mathsf{series}_{\mathbb{C}}=\gamma^{-1}(\mathsf{series})$, and $\mathsf{ord}_{\mathbb{C}}=\{(x,y)\mid x<y\}$. The nodes that are labeled by $\mathsf{linear}$ could be recovered as those inner nodes whose children are related by $\mathsf{ord}$, while the nodes that are labeled by $\mathsf{parallel}$ could be recovered as those inner nodes which are labeled neither by $\mathsf{series}$ nor by $\mathsf{linear}$. Based on 18 and as a consequence of 17, we can design a $\operatorname{C}_{2}{\!\!}\operatorname{MSO}$-transduction which produces the cotree of a cograph $G$ from $G$.
Corollary 19.
There exists a non-deterministic $\operatorname{C}_{2}{\!\!}\operatorname{MSO}$-transduction $\tau$ such that, for each directed cograph $G$ modeled by the $\{\mathsf{edge}\}$-structure $\mathbb{G}$, $\tau(\mathbb{G})$ is non-empty and every output in $\tau(\mathbb{G})$ is equal to some $\{\mathsf{ancestor},\mathsf{series},\mathsf{ord}\}$-structure $\mathbb{C}$ representing the cotree $(T,\gamma,<)$ of $G$.
As for undirected graphs the set system of modules is partitive, a simpler result in which the label $\mathsf{linear}$ is not needed holds for undirected cographs.
5. Transducing split and bi-join decompositions
In this section we consider systems of bipartitions. In graph theory, two known systems of bipartitions are splits and bi-joins. Both are instances of so called weakly-bipartitive systems of bipartitions (or partitive systems of bipartitions in case of undirected graphs). In this section, we first give a $\operatorname{C}_{2}{\!\!}\operatorname{MSO}$-transduction of the canonical tree-like decomposition of weakly-bipartitive systems of bipartitions and then derive a $\operatorname{C}_{2}{\!\!}\operatorname{MSO}$-transductions for split and bi-join decompositions of a graph.
5.1. Transducing weakly-bipartitive trees
A bipartition system is a pair $(U,\mathcal{B})$ consisting of a finite set $U$, the universe, and a family $\mathcal{B}$ of bipartitions of $U$ such that $\{\emptyset,U\}\notin\mathcal{B}$ and $\{\{a\},U\setminus\{a\}\}\in\mathcal{B}$ for all $a\in U$.
To model bipartition systems, we use the extended vocabulary $\{\mathsf{BIPART}\}$ where $\mathsf{BIPART}$ is a unary set predicate name. A bipartition system $(U,\mathcal{B})$ is thus naturally modeled as the $\{\mathsf{BIPART}\}$-structure $\mathbb{B}$ with universe $U_{\mathbb{B}}=U$ and interpretation $\{\{X,U\setminus X\}\mid X\in\mathsf{BIPART}_{\mathbb{B}}\}=\mathcal{B}$.
Two bipartitions $\{X,Y\}$ and $\{X^{\prime},Y^{\prime}\}$ of a set $U$ overlap if the four sets $X\cap X^{\prime}$, $X\cap Y^{\prime}$, $Y\cap X^{\prime}$ and $Y\cap Y^{\prime}$ are non-empty. A bipartition family is said to be laminar (aka overlap-free) if no two bipartitions in $\mathcal{B}$ overlap. By extension, we call a family $\mathcal{B}$ of bipartition of a set $U$ laminar whenever $(U,\mathcal{B})$ is a bipartition system which is laminar. Similarly to laminar set systems, we can associate a tree with each laminar bipartition system. The tree $T$, called the laminar tree induced by $(U,\mathcal{B})$ (or laminar tree of $(U,\mathcal{B})$), associated with a laminar bipartition system $(U,\mathcal{B})$ is unrooted, each element of $U$ corresponds to a leaf of $T$ and for each bipartition $\{X,Y\}$ in $\mathcal{B}$ there is exactly one edge $e$ such that $X=L(T_{1})$ and $Y=L(T_{2})$ for the two connected components $T_{1},T_{2}$ of TeT-e. We remark that there is a unique such tree and each inner node has degree at least three. Furthermore, the size of the laminar tree is linearly bounded in the size of the universe $U$.
A bipartition system $(U,\mathcal{B})$ is said to be weakly-bipartitive if for every two overlapping bipartitions $\{X,Y\},\{X^{\prime},Y^{\prime}\}\in\mathcal{B}$, the biparitions $\{X\cup X^{\prime},Y\cap Y^{\prime}\}$, $\{X\cup Y^{\prime},Y\cap X^{\prime}\}$, $\{Y\cup X^{\prime},X\cap Y^{\prime}\}$ and $\{Y\cup Y^{\prime},X\cap X^{\prime}\}$ are in $\mathcal{B}$. It is bipartitive, if additionally for every two overlapping bipartitions $\{X,Y\},\{X^{\prime},Y^{\prime}\}\in\mathcal{B}$, the bipartition $\{X\operatorname*{\bigtriangleup}X^{\prime},Y\operatorname*{\bigtriangleup}X^{\prime}\}$ is also in $\mathcal{B}$. By extension, we call a family $\mathcal{B}$ of bipartitions of a set $U$ weakly-bipartitive or bipartitive whenever $(U,\mathcal{B})$ is a bipartition system which is weakly-bipartitive or partitive, respectively.
Figure 7. Bipartitions to be included in a weakly-bipartitive system of bipartitions (excluds the rightmost) or in a bipartitive system bipartitions (includes the rightmost) for two overlapping bipartitions $\{X,Y\}$ and $\{X^{\prime},Y^{\prime}\}$.
A bipartition $\{X,Y\}$ of a bipartition system $(U,\mathcal{B})$ is said to be strong if $\{X,Y\}$ does not overlap with any other bipartition in $\mathcal{B}$. We denote the subfamily of strong bipartitions in $\mathcal{B}$ by $\mathcal{B}_{!}$ and remark that $\mathcal{B}_{!}$ is a laminar family. Hence, it induces a laminar tree $T$. By extension, we say that $T$ is induced by the bipartition system $(U,\mathcal{B})$ (or simply by $\mathcal{B}$). As the laminar tree $T$ of a bipartition system $(U,\mathcal{B})$ is undirected, we model $T$ by the $\{\mathsf{t}\text{-}\mathsf{edge}\}$-structure $\mathbb{T}$ with universe $U_{\mathbb{T}}=V(T)$ and binary relation $\mathsf{t}\text{-}\mathsf{edge}_{\mathbb{T}}$ modeling the edge relation of $T$.
The following lemma gives a connection between laminar trees of set systems and laminar trees of bipartition systems. Note that a very similar statement follows from [Rao06, Lemma 1.14] and hence we only provide a proof for sake of completeness.
Lemma 20.
Let $(U,\mathcal{B})$ be a bipartition system and $a\in U$. Then $(U,\mathcal{S}_{a})$ is a set system where $\mathcal{S}_{a}:=\{X:a\notin X\text{ and }\{X,U\setminus X\}\in\mathcal{B}\}\cup\{\{a\},U\}$. Moreover, the laminar tree of $(U,\mathcal{B})$ can be obtained from the laminar tree of $(U,\mathcal{S}_{a})$ by deleting the root, making node $\{a\}$ adjacent to the node $U\setminus\{a\}$ and making the tree undirected.
Proof 5.1.
Let $\mathcal{B}$ be a family of bipartitions of a set $U$ and $a\in U$ any element. First observe that $(U,\mathcal{S}_{a})$ is a set system as $\emptyset\notin\mathcal{S}_{a}$ (since $\{\emptyset,U\}\notin\mathcal{B}$), $U\in\mathcal{S}_{a}$, and $\{b\}\in\mathcal{S}_{a}$ for every $b\in U$ (since $\{a\}\in\mathcal{S}_{a}$ and for $b\neq a$, $\{\{b\},U\setminus\{b\}\}\in\mathcal{B}$).
To argue the second assertion, assume that $T$ is the laminar tree of the set system $(U,\mathcal{S}_{a})$. First observe that in the set system $(U,\mathcal{S}_{a})$ the element $a$ is not contained in any set apart from $U$ and $\{a\}$ and hence the leaf $\{a\}$ is a child of the root $U$ of $T$. Additionally, the set $U\setminus\{a\}$ is in $S_{a}$ as the bipartition $(\{a\},U\setminus\{a\})$ is in $\mathcal{B}$ by definition and hence $U\setminus\{a\}$ is also a child of the root $U$ of $T$. Let $T^{\prime}$ be the tree obtained from $T$ by removing node $U$, making all edges undirected and adding an edge from node $\{a\}$ to node $U\setminus\{a\}$. Indeed, by our previous observations, $T^{\prime}$ is an undirected tree. Furthermore, there is a correspondence between elements of $U$ and the leaves of $T^{\prime}$ (element $b\in U$ corresponds to leaf $\{b\}$ of $T^{\prime}$). To check that $T^{\prime}$ is indeed the laminar tree of $T$, we need to verify that for each bipartition $\{X,Y\}$ in $\mathcal{B}$ there is exactly one edge $e$ of $T^{\prime}$ such that $X=L(T^{\prime}_{1})$ and $Y=L(T^{\prime}_{2})$ for the two connected components $T^{\prime}_{1},T^{\prime}_{2}$ of $T^{\prime}-e$. We say that the edge $e$ implements the bipartition $\{X,Y\}$. Assume that $\{X,Y\}$ is any bipartition from $\mathcal{B}$. First consider the case that $\{X,Y\}$ is the bipartition $\{\{a\},U\setminus\{a\}\}$. Then $\{X,Y\}$ is clearly implemented by the edge $\{\{a\},U\setminus\{a\}\}$ of $T^{\prime}$. Now assume that $\{X,Y\}$ is any other bipartition in $\mathcal{B}$ and assume, without loss of generality, that $a\in Y$. Hence, $X\in\mathcal{S}_{a}$. Let $X^{\prime}$ be the parent of $X$ in $T$. Note that $X$ must be a proper subset of $U\setminus\{a\}$ as by our assumption $\{X,Y\}$ is not the partition $\{\{a\},U\setminus\{a\}$ and hence $X^{\prime}$ cannot be $U$. As we only deleted edge incident to $U$ in our construction of $T^{\prime}$, $\{X,X^{\prime}\}$ is an edge in $T^{\prime}$ which clearly implements the bipartition $\{X,Y\}$. Finally, every bipartition $\{X,Y\}\in\mathcal{B}$ is clearly implemented by only one edge as every inner node of the tree $T^{\prime}$ has degree at least three. Therefore, $T^{\prime}$ is the laminar tree of $(U,\mathcal{B})$.
Lemma 21.
Let $\Sigma$ be an extended vocabulary, including a unary set predicate name $\mathsf{BIPART}$ and not including the binary relational symbol $\mathsf{t}\text{-}\mathsf{edge}$. There exists a non-deterministic overlay $\operatorname{C}_{2}{\!\!}\operatorname{MSO}$-transduction $\tau$ such that, for each bipartition system $(U,\mathcal{B})$ represented as the $\{\mathsf{BIPART}\}$-structure $\mathbb{B}$ and inducing the laminar tree $T$ with L(T)=U{L(T)=U}, and for each $\Sigma$-structure $\mathbb{A}$ with $\mathbb{B}\sqsubseteq\mathbb{A}$, $\tau(\mathbb{A})$ is non-empty and every output in $\tau(\mathbb{A})$ is equal to $\mathbb{A}\sqcup\mathbb{T}$ for some $\{\mathsf{t}\text{-}\mathsf{edge}\}$-structure $\mathbb{T}$ representing $T$.
Proof 5.2.
Let $\mathcal{B}$ be a family of bipartitions of $U$ represented as the extended relational structure $\mathbb{B}$ and $\mathbb{A}$ a $\Sigma$-structure with $\mathbb{B}\sqsubseteq\mathbb{A}$. We first use the following two simple atomic $\operatorname{C}_{2}{\!\!}\operatorname{MSO}$-transductions to obtain a set system from $(U,\mathcal{B})$. Guess any colouring $A$ for which there is a single $a\in U$ with $A(a)$ being satisfied. Interpret using the following formula to obtain a $\mathsf{SET}$-predicate:
$\phi_{\mathsf{SET}}(X):=\forall a(A(a)\rightarrow(X=\{a\}\lor X=U\lor a\notin X))\land(\mathsf{BIPART}(X)\lor\mathsf{BIPART}(U\setminus X)).$
The resulting set system $(U,\mathcal{S}_{a})$ represented by the $\{\mathsf{SET}\}$-structure $\mathbb{S}$ is a set system by 20. Hence, we can use the transduction from 14 to obtain the $\Sigma\cup\{\mathsf{ancestor}\}$-structure $\mathbb{A}\sqcup\mathbb{T}$ where $\mathbb{T}$ is the $\{\mathsf{ancestor}\}$-structure representing the laminar tree $T$ of $(U,\mathcal{S}_{a})$. We now use the following two atomic transductions to obtain the weakly-bipartitive tree of $(U,\mathcal{B})$ according to 20. We first restrict the universe to all nodes which are not the root of $\mathbb{T}$. We then interpret using the following formula to obtain the $\mathsf{t}\text{-}\mathsf{edge}$-predicate which makes the graph undirected and adds an edge from $a$ to the node representing the set $U\setminus\{a\}$ which is the root of $T$ after removing the node corresponding to $U$:
$\phi_{\mathsf{t}}\text{-}\mathsf{edge}(x,y):=$ $\mathsf{parent}(x,y)\lor\mathsf{parent}(y,x)\lor$
$(x=a\land y\not=a\land\lnot\exists z\mathsf{parent}(z,y))\lor(x\not=a\land y=a\land\lnot\exists z\mathsf{parent}(z,x)).$
Finally, the transduction keeps only $\mathsf{t}\text{-}\mathsf{edge}$ and any $Q\in\Sigma$ as relations for its output. We do not need to filter the outputs as any guess for $A$ yields a valid laminar tree.
Similar to the set system case, for each weakly-bipartitive family $\mathcal{B}$ we can equip the laminar tree with a labelling of its inner nodes and a partial order which characterizes the family $\mathcal{B}$. We remind the reader that for a node $t$ of a tree $T$ with neighbour $s$ we denote by $T^{t}_{s}$ the connected component of TtT-t containing $s$.
{thmC}
[[dM03, Theorem 3]] Let $\mathcal{B}$ be a weakly-bipartitive family, $\mathcal{B}_{!}$ be its subfamily of strong bipartitions, and $T$ the laminar tree it induces. There exists a total labeling function $\lambda$ from the set $V(T)\setminus L(T)$ of the inner nodes of $T$ to the set $\{\mathsf{degenerate},\mathsf{prime},\mathsf{linear}\}$, and, for each inner node $t\in\lambda^{-1}(\mathsf{linear})$, a linear ordering $_{t}$ of its neighbours, such that every inner node of degree exactly three is labeled by $\mathsf{degenerate}$ and the following conditions are satisfied:
  • for each bipartition $\{X,Y\}\in\mathcal{B}\setminus\mathcal{B}_{!}$, there exists $t\in V(T)\setminus L(T)$ and a subset $\mathcal{C}$ of neighbours of $t$ such that either $X$ or $Y$ is equal to the set $\bigcup_{c\in\mathcal{C}}L(T^{t}_{c})$ and either $\lambda(t)=\mathsf{linear}$ and $\mathcal{C}$ is a $_{t}-$interval, or $\lambda(t)=\mathsf{degenerate}$;
  • conversely, for each inner node $t$ and each non-empty subset $\mathcal{C}$, $\mathcal{C}\not=U$ of neighbours of $t$, the bipartition $\big{\{}\bigcup_{c\in\mathcal{C}}L(T^{t}_{c}),\bigcup_{c\notin\mathcal{C}}L(T^{t}_{c})\big{\}}$ is a member of $\mathcal{B}$ if either $\lambda(t)=\mathsf{linear}$ and $\mathcal{C}$ is $_{t}-$interval or $\lambda(t)=\mathsf{degenerate}$.
The tree $T$ and its labelling function $\lambda$ are uniquely determined by $\mathcal{B}$. We note that the total orders $_{t}$ are uniquely determined by $\mathcal{B}$ up to inverting and cyclic shifting, i.e., for a linear order << of $X=\{x_{1},\dots,x_{\ell}\}$ with $x_{1}<\dots<x_{\ell}$ we call the order $^{\prime}$ of $X$ with $x_{\ell}<^{\prime}x_{1}<^{\prime}\dots<^{\prime}x_{\ell-1}$ a cyclic shift of <<. To see this, note that in the theorem above the bipartitions obtained from nodes with label $\mathsf{linear}$ can be equivalently obtained from a $_{t}-$interval or the complement of a $_{t}-$interval. Indeed, a cyclic shifting of a linear order << maintains the property of being either a <<-interval or the complement of a <<-interval. Neglecting that $_{t}$ are only unique up to inverting and cyclic shifting, we represent every weakly-bipartitive family $\mathcal{B}$ by a triple $(T,\lambda,(<_{t})_{t\in\lambda^{-1}(\mathsf{linear})})$ where $T$ is the laminar tree induced by $\mathcal{B}_{!}$, $\lambda:V(T)\setminus L(T)\to\{\mathsf{degenerate},\mathsf{prime},\mathsf{linear}\}$ is the labeling function, and $_{t}$ is the total order of the neighbours of $t$ described above. Note that in this case $\bigcup_{t\in\lambda^{-1}(\mathsf{linear})}<_{t}$ is not a partial order as it is not necessarily transitive. We call the triple $(T,\lambda,(<_{t})_{t\in\lambda^{-1}(\mathsf{linear})})$ from 5.2 the bipartitive tree of $\mathcal{B}$.
We model weakly-bipartitive trees $(T,\lambda,(<_{t})_{t\in\lambda^{-1}(\mathsf{linear})})$ of a weakly-bipartitive family $\mathcal{B}$ of bipartition of $U$ by the $\{\mathsf{t}\text{-}\mathsf{edge},\mathsf{degenerate},\mathsf{cross}\}$-structure $\mathbb{T}$ of universe $U_{\mathbb{T}}=V(T)$ such that $\langle V(T),\mathsf{t}\text{-}\mathsf{edge}_{\mathbb{T}}\rangle\sqsubseteq\mathbb{T}$ models $T$ with L(T)=U{L(T)=U}, $\mathsf{degenerate}_{\mathbb{T}}$ is a unary relation which selects all inner nodes of $T$ of label $\mathsf{degenerate}$, i.e., $\mathsf{degenerate}_{\mathbb{T}}=\lambda^{-1}(\mathsf{degenerate})$}, and $\mathsf{cross}_{\mathbb{T}}$ is a relation of arity five selecting all 5-tuples $(t,w,x,y,z)$ for which the chord $wy$ crosses the chord $xz$ in $_{t}$ understood as a cyclic order. Similarly to the weakly-partitive case, we can recover the inner nodes of $T$ that are labeled by $\mathsf{linear}$ through an $\operatorname{MSO}$-formula as those inner nodes whose neighbours are related by $\mathsf{cross}$. Then $\mathsf{prime}$ nodes are the ones which are labeled neither by $\mathsf{degenerate}$ nor by $\mathsf{linear}$ which can be obtained by an $\operatorname{MSO}$-transduction. Furthermore, 5.2, gives rise to an $\operatorname{MSO}$-transduction which takes as input a weakly-bipartitive tree and outputs the weakly-bipartitive set system which induced it.
We can now extend the transduction from 21 to obtain a $\operatorname{C}_{2}{\!\!}\operatorname{MSO}$-transduction producing the weakly-bipartitive tree of a weakly-bipartitive family.
Theorem 22.
There exists a non-deterministic $\operatorname{C}_{2}{\!\!}\operatorname{MSO}$-transduction $\tau$ such that, for every weakly-bipartitive bipartition system $(U,\mathcal{B})$ represented as the $\{\mathsf{BIPART}\}$-structure $\mathbb{B}$, $\tau(\mathbb{B})$ is non-empty and every output in $\tau(\mathbb{B})$ is equal to some $\{\mathsf{t}\text{-}\mathsf{edge},\mathsf{degenerate},\mathsf{cross}\}$-structure $\mathbb{T}$ representing the weakly-bipartitive tree $\left(T,\lambda,(<_{t})_{t\in\lambda^{-1}(\mathsf{linear})}\right)$ of $(U,\mathcal{B})$.
Proof 5.3.
21 already provides a $\operatorname{C}_{2}{\!\!}\operatorname{MSO}$-transduction which on input $\mathbb{B}$ outputs the $\{\mathsf{BIPART},\mathsf{t}\text{-}\mathsf{edge}\}$-structure $\mathbb{B}\sqcup\mathbb{T}^{\prime}$, where $\mathbb{T}^{\prime}$ represents the $\{\mathsf{t}\text{-}\mathsf{edge}\}$-structure modeling the laminar tree $T$ of $(U,\mathcal{B})$. To complete the proof of the theorem, we provide an $\operatorname{MSO}$-transduction which outputs $\mathbb{T}$ given input $\mathbb{B}\sqcup\mathbb{T}^{\prime}$. For the transduction we define an $\operatorname{MSO}$-formula $\phi_{\mathsf{degenerate}}$ which selects the inner nodes of $T$ which are labeled $\mathsf{degenerate}$, and a formula $\phi_{\mathsf{cross}}$ which selects 5-tuples of nodes $(t,w,x,y,z)$ such that either $w<_{t}x<_{t}y<_{t}z$ or $z<_{t}y<_{t}x<_{t}w$ or some cyclic shift of this. Recall, that the ordering $_{t}$ for each inner node $t$ of the neighbours of $t$ are only unique up to inverting the order and cyclic shift. We use a functional symbol $\mathsf{leafset}(t,s)$ which is interpreted as the set $L(T_{s}^{t})$ where $T_{s}^{t}$ is the connected component of TtT-t containing $s$. The functional predicate $\mathsf{leafset}$ clearly can be defined in $\operatorname{MSO}$ as an element $z$ belongs to $\mathsf{leafset}(t,s)$ if and only if $z$ is a leaf and the path from $z$ to $s$ does not contain $t$.
Observe that by 5.2 an inner node $t$ is labeled $\mathsf{degenerate}$ if and only if for every pair of neighbours $s_{1},s_{2}$ of $t$ the bipartition $\big{\{}L(T^{t}_{s_{1}})\cup L(T^{t}_{s_{2}}),U\setminus\big{(}L(T^{t}_{s_{1}})\cup L(T^{t}_{s_{2}})\big{)}\big{\}}$ is in $\mathcal{B}$. We can define this in $\operatorname{MSO}$ using the following formula:
$\phi_{\mathsf{degenerate}}$ $(x):=\forall y\forall z\Big{(}\big{(}\mathsf{t}\text{-}\mathsf{edge}(x,y)\land\mathsf{t}\text{-}\mathsf{edge}(x,z)\big{)}\rightarrow$
$\mathsf{BIPART}\big{(}\mathsf{leafset}(x,y)\cup\mathsf{leafset}(x,z),U\setminus(\mathsf{leafset}(x,y)\cup\mathsf{leafset}(x,z))\big{)}\Big{)}.$
By 5.2, a set $\mathcal{C}$ of neighbours of $t$ is consecutive in $_{t}$ (understood as a cyclic order) if and only if, the bipartition $\big{\{}\bigcup_{c\in\mathcal{C}}L(T^{t}_{c}),U\setminus\big{(}\bigcup_{c\in\mathcal{C}}L(T^{t}_{c})\big{)}\big{\}}$ is a member of $\mathcal{B}$. Furthermore, we can express that $w<_{t}x<_{t}y<_{t}z$ in $_{t}$ (considered as a cyclic order) if there is a $_{t}-$interval $\mathcal{C}$ which does not contain $z$, but contains two $_{t}-$intervals $\mathcal{C}_{1},\mathcal{C}_{2}$ such that $w,x\in\mathcal{C}_{1}$, $y\notin\mathcal{C}_{1}$ and $w\notin\mathcal{C}_{2}$ and $x,y\in\mathcal{C}_{2}$. Hence, we can define in $\operatorname{MSO}$ the predicate $\mathsf{cross}$ as follows:
$\phi_{\mathsf{cross}}(t,x_{1},x_{2},x_{3},x_{4}):=\Big{[}\lnot\mathsf{degenerate}(t)\land\bigwedge_{i<j}x_{i}\not=x_{j}\land\bigwedge_{i}\mathsf{t}\text{-}\mathsf{edge}(t,x_{i})\land\exists X,Y,Z$
$\Big{(}\mathsf{BIPART}(X)\land\mathsf{BIPART}(Y)\land\mathsf{BIPART}(Z)\land Y\subseteq X\land Z\subseteq X\land\mathsf{leafset}(t,x_{1})\subseteq Y\setminus Z\land$
$\mathsf{leafset}(t,x_{2})\subseteq Y\cap Z\land\mathsf{leafset}(t,x_{3})\subseteq Z\setminus Y\land\mathsf{leafset}(t,x_{4})\subseteq U\setminus X\Big{)}\Big{]}.$
Note that any node which is not labeled $\mathsf{degenerate}$ and for which there is a set $\mathcal{C}$ of at least two neighbours of $t$ for which $\big{\{}\bigcup_{c\in\mathcal{C}}L(T^{t}_{c}),U\setminus\big{(}\bigcup_{c\in\mathcal{C}}L(T^{t}_{c})\big{)}\big{\}}$ is a member of $\mathcal{B}$ must be labeled $\mathsf{linear}$. Additionally, note that any node with less than three children which is not labeled $\mathsf{prime}$ must be labeled $\mathsf{degenerate}$.
Finally, the transduction forgets the predicate $\mathsf{BIPART}$ to produce the output $\mathbb{T}$.
Indeed, considering a bipartitive family $\mathcal{B}$ makes the use of nodes labeled $\mathsf{linear}$ in the associated tree obsolete, as the following theorem states. {thmC}[[dM03, Theorem 4]] Let $\mathcal{B}$ be a weakly-biaprtitive family and let $(T,\lambda,(<_{t})_{t\in\lambda^{-1}(\mathsf{linear})})$ be the weakly-bipartitive tree it induces. If $\mathcal{B}$ is bipartitive, then $\lambda^{-1}(\mathsf{linear})=\emptyset$ and all $_{t}$ are empty. As a consequence, we can capture any bipartitive family $\mathcal{B}$ by the simpler structure $(T,\lambda)$, called the bipartitive tree induced by $\mathcal{B}$ (or the bipartitive tree of $\mathcal{B}$), where $\lambda$ is now a labeling function labeling each inner nodes of $T$ by either $\mathsf{degenerate}$ or $\mathsf{prime}$. Hence, we obtain the following as a corollary from 22, where naturally the bipartitive tree $(T,\lambda)$ of a bipartitive family $\mathcal{B}$ is modeled by a $\{\mathsf{t}\text{-}\mathsf{edge},\mathsf{degenerate}\}$-structure $\mathbb{B}$.
Corollary 23.
There exists a non-deterministic $\operatorname{C}_{2}{\!\!}\operatorname{MSO}$-transduction $\tau$ such that, for each bipartitive set system $(U,\mathcal{B})$ represented as the $\{\mathsf{BIPART}\}$-structure $\mathbb{B}$, $\tau(\mathbb{B})$ is non-empty and every output in $\tau(\mathbb{B})$ is equal to some $\{\mathsf{t}\text{-}\mathsf{edge},\mathsf{degenerate}\}$-structure $\mathbb{T}$ representing the bipartitive tree $(T,\lambda)$ of $(U,\mathcal{B})$.
5.2. Application to split decomposition
Let $G$ be a directed graph. A split in $G$ is a bipartition $\{X,Y\}$ of $V(G)$ such that there are subsets $X^{\operatorname{in}},X^{\operatorname{out}}\subseteq X$, $Y^{\operatorname{in}},Y^{\operatorname{out}}\subseteq Y$ such that for all pairs of vertices $x\in X$ and $y\in Y$ the edge $xy\in E(G)$ if and only if either $x\in X^{\operatorname{out}}$ and $y\in Y^{\operatorname{in}}$ or $x\in Y^{\operatorname{out}}$ and $y\in X^{\operatorname{in}}$. Note that if $G$ is connected then $X^{\operatorname{in}}\cup X^{\operatorname{out}}\not$$\emptyset$ and $Y^{\operatorname{in}}\cup Y^{\operatorname{out}}\not$$\emptyset$. We call splits $\{X,Y\}$ for which either $|X|=1$ or $|Y|=1$ trivial splits. A graph is considered prime with respect to splits if it has only trivial splits. In a family $\mathcal{B}$ of splits we use the notion of strong splits to coincide with the notion of strong bipartitions. See Figure 8 (left) for an example of a strong split in a directed graph.
One can define split decomopositions in greater generality, however, we only require the canonical split decomposition obtained by considering only strong splits in the following. Hence, consider a directed graph $G$ and let $\mathcal{B}_{\mathsf{split}}^{G}$ be the set of all splits of $G$. The following theorem shows that $\mathcal{B}_{\mathsf{split}}^{G}$ is weakly-bipartitive.
{thmC}
[[Cun82]] For every connected graph $G$ family of splits $\mathcal{B}_{\mathsf{split}}^{G}$ is weakly-bipartitive.
[…]
  • $H$ is a directed graph consisting of the disjoint union of the graphs $G_{u}$ for every inner node $u$ of $T$ and
  • $F$ is a symmetric edge relation containing edges $\vec{uv}\vec{vu}$ for every edge $uv$ of $T$.
For a clear distinction, we call the edges in $H$ $\mathsf{c}$-edges and the edges in $F$ $\mathsf{t}$-edges. Note that $\mathsf{t}$-edges correspond to the edges of $T$. See Figure 8 for an illustration.
Figure 8. A directed graph $G$ with 33 strong splits (left). The weakly-partitive tree of $G$ (middle) and the split decomposition of $G$ where $\mathsf{c}$-edges are the thin, black edges, $\mathsf{t}$-edges are the fat, dashed, blue edges and the (non-singleton) components of the decomposition are signified by grey circles.
The following (standard) technical lemma is needed to show that a graph can be recovered from an enriched split decomposition (by an $\operatorname{MSO}$ transduction).
Lemma 24 (See e.g. Lemma 2.10 in [AKK17]).
Let $G$ be a connected, directed graph and $(H,F)$ the enriched split decomposition of $G$. Then $uv\in E(G)$ if and only if there exists a directed path from $u$ to $v$ in $(H,F)$ on which c-edges and t-edges alternate.
To model enriched split decompositions as relational structures we use the relational vocabulary $\{\mathsf{t}\text{-}\mathsf{edge},\mathsf{c}\text{-}\mathsf{edge}\}$ where $\mathsf{t}\text{-}\mathsf{edge}$ and $\mathsf{c}\text{-}\mathsf{edge}$ are binary relational symbols. An enriched split decomposition $(H,F)$ of a graph $G$ is the $\{\mathsf{t}\text{-}\mathsf{edge},\mathsf{c}\text{-}\mathsf{edge}\}$-relation $\mathbb{H}$ with universe $U_{\mathbb{H}}=V(H)$, $\mathsf{t}\text{-}\mathsf{edge}_{\mathbb{H}}$ the set of t-edges in $(H,F)$ and $\mathsf{c}\text{-}\mathsf{edge}_{\mathbb{H}}$ the set of c-edges of $(H,F)$.
Since the nodes of $T$ correspond to the graphs $G_{u}$ which are the connected components of the enriched split decomposition $(H,F)$ after removing $F$ and the vertices of $G$ are the leaves of $T$, it is clear that the 24 allows to $\operatorname{MSO}$-transduce the graph $G$ from its enriched split decomposition.
We use 21 to transduce the canonical split decomposition of a graph.
Theorem 25.
There exists a non-deterministic $\operatorname{C}_{2}{\!\!}\operatorname{MSO}$-transduction $\tau$ such that for any connected, directed graph $G$ represented as the $\{\mathsf{edge}\}$-structure $\mathbb{G}$; $\tau(\mathbb{G})$ is non-empty and every output in $\tau(\mathbb{G})$ is equal to some $\{\mathsf{t}\text{-}\mathsf{edge},\mathsf{c}\text{-}\mathsf{edge}\}$-structure $\mathbb{H}$ representing the canonical split decomposition of $G$.
Proof 5.4.
Let $G$ be a connected, directed graph represented by the $\{\mathsf{edge}\}$-structure $\mathbb{G}$. We associate the following objects with $G$:
  • let $\mathcal{B}_{\mathsf{split}}^{G}$ be the family of splits of $G$ and $\mathbb{B}$ the $\{\mathsf{BIPART}\}$-structure modeling the weakly-bipartitive bipartition system $(V(G),\mathcal{B}_{\mathsf{split}}^{G})$ with $U_{\mathbb{B}}=V(G)$;
  • let $T$ be the laminar tree induced by the weakly-bipartitive family $\mathcal{B}_{\mathsf{split}}^{G}$ and let $\mathbb{T}$ be the $\{\mathsf{t}\text{-}\mathsf{edge}\}$-structure modeling it with $U_{\mathbb{T}}=V(T)$ and $L(T)=V(G)\subset U_{\mathbb{T}}$;
  • let $(H,F)$ be the enriched canonical split decomposition of $G$ represented by the $\{\mathsf{t}\text{-}\mathsf{edge},\mathsf{c}\text{-}\mathsf{edge}\}$-structure $\mathbb{H}$.
Our $\operatorname{C}_{2}{\!\!}\operatorname{MSO}$-transduction is obtained by composing the following transductions:
  • $\tau_{1}$: an $\operatorname{MSO}$-interpretation which outputs the $\{\mathsf{edge},\mathsf{BIPART}\}$-structure $\mathbb{G}\sqcup\mathbb{B}$ from $\mathbb{G}$;
  • $\tau_{2}$: the non-deterministic $\operatorname{C}_{2}{\!\!}\operatorname{MSO}$-transduction from 21 which outputs the $\{\mathsf{edge},\mathsf{BIPART},\mathsf{t}\text{-}\mathsf{edge}\}$-structure $\mathbb{G}\sqcup\mathbb{B}\sqcup\mathbb{T}$ on input $\mathbb{G}\sqcup\mathbb{B}$;
  • $\tau_{3}$: an $\operatorname{MSO}$-transduction which produces the $\{\mathsf{t}\text{-}\mathsf{edge},\mathsf{c}\text{-}\mathsf{edge}\}$-structure $\mathbb{H}$ from $\mathbb{G}\sqcup\mathbb{B}\sqcup\mathbb{T}$.
To define $\tau_{1}$ it is sufficient to observe that there is an $\operatorname{MSO}$-formula $\phi_{\mathsf{BIPART}}(X)$ with one monadic free-variable $X$ such that $\phi_{\mathsf{BIPART}}$ is satisfied exactly when $\{X,U\setminus X\}$ is a split in $G$. As $\tau_{2}$ is constructed in 21, we are left with describing how to obtain $\tau_{3}$. We first need to find unique representatives for the vertices in $V(H)$. Recall that $V(H)$ consists of two marker vertices $\vec{uv},\vec{vu}$ for every edge $uv$ in $T$. We use the following atomic transductions to ensure unique representatives for all vertices in $V(H)$. We first guess a colouring $R$ which identifies one non-leaf element of $\mathbb{G}\sqcup\mathbb{B}\sqcup\mathbb{T}$ as root. We use this root element to interpret an auxiliary binary relation $\mathsf{parent}$ in the usual way. Now we copy the structure $\mathbb{G}\sqcup\mathbb{B}\sqcup\mathbb{T}$ once introducing the binary relation $\mathsf{copy}_{1}$ in which $\mathsf{copy}_{1}(x,y)$ indicates that $x$ is the copy of $y$. Assume that $\{s,t\}$ is an edge in $T$ corresponding to a strong bipartition in $\mathcal{B}_{\mathsf{split}}^{G}$ and $\mathsf{parent}(s,t)$ is satisfied. We use the original node $t$ to represent the marker $\vec{st}$ and we use the copy of $t$ to represent the marker $\vec{ts}$. As every node has a unique parent in a rooted tree, for each marker the chosen representatives are unique. To express this we use the predicate $\mathsf{marker}(x,y,z)$ which expresses that $x$ is the representative of the marker $\vec{yz}$ defined as follows:
$\mathsf{marker}(x,y,z):=\Big{(}\mathsf{parent}(y,z)\land x=z\Big{)}\lor\Big{(}\mathsf{parent}(z,y)\land\mathsf{copy}_{1}(x,y)\Big{)}.$
We now filter the universe only keeping all representatives of markers (recall that every original vertex $u$ of $G$ has a marker $\vec{uv}$, where $v$ is the neighbour of $u$ in $T$, representing it). We are left with defining the two edge relations $\mathsf{t}\text{-}\mathsf{edge}$ and $\mathsf{c}\text{-}\mathsf{edge}$. To do so, we use the previously defined functional predicate $\mathsf{leafset}(t,s)$ which is interpreted by the set $L(T_{s}^{t})$.
First note that markers $\vec{st}$ and $\vec{s^{\prime}t^{\prime}}$ are in the same component of $H$ if and only if $s=s^{\prime}$. By definition $\vec{st}\vec{st^{\prime}}$ is a c-edge in $(H,F)$ if and only if there is an edge $uu^{\prime}\in E(G)$ with $u\in L(T^{s}_{t})$ and $u^{\prime}\in L(T^{s}_{t^{\prime}})$. Consequently, we can express $\mathsf{c}\text{-}\mathsf{edge}$ as follows:
$\phi_{\mathsf{c}\text{-}\mathsf{edge}}(x,x^{\prime}):=$ $\exists t\exists t^{\prime}\exists s\exists y\exists y^{\prime}\Big{(}\mathsf{t}\text{-}\mathsf{edge}(s,t)\land\mathsf{t}\text{-}\mathsf{edge}(s,t^{\prime})\land\mathsf{marker}(x,s,t)\land$
$\mathsf{marker}(x^{\prime},s,t^{\prime})\land y\in\mathsf{leafset}(s,t)\land y^{\prime}\in\mathsf{leafset}(s,t^{\prime})\land E(y,y^{\prime})\Big{)}.$
On the other hand, the $\mathsf{t}\text{-}\mathsf{edge}$-relation in $\mathbb{H}$ is the relation between pairs of markers and hence can be easily (re)-defined using the $\mathsf{marker}$ predicate by:
$\phi_{\mathsf{t}\text{-}\mathsf{edge}}(x,x^{\prime}):=\exists y\exists z\big{(}\mathsf{marker}(x,y,z)\land\mathsf{marker}(x^{\prime},z,y)\big{)}.$
Finally, the transduction only keeps the relation $\mathsf{t}\text{-}\mathsf{edge}$ and $\mathsf{c}\text{-}\mathsf{edge}$. We remark that any choice of $R$ yields a valid output and hence we do not need to apply filtering.
5.3. Application to bi-join decomposition
Bi-joins were introduced in [dMR05, Rao06] as a generalization of modules and splits in undirected graphs, and were used for instance in [LdMR07] to decide isomorphism in some graph classes of rank-width at most $2$. We follow definitions and notations from [Rao06]. A bi-join in an undirected graph $G$ is a bipartition $\{X,Y\}$ of $V(G)$ such that $\left|X\right|$, $\left|Y\right|\geq 1$ and there are (possibly empty) subsets $X^{\prime}\subseteq X$ and $Y^{\prime}\subseteq Y$ such that $\big{(}X^{\prime}\times Y^{\prime}\big{)}\cup\big{(}(X\setminus X^{\prime})\times(Y\setminus Y^{\prime})\big{)}\subseteq E(G)$, and there are no additional edges between $X$ and $Y$ in $G$. If $X$ is a module in a graph $G$, then $\{X,V(G)\setminus X\}$ is a bi-join with $X^{\prime}=\emptyset$, and if $V(G)\setminus X$ is complete to $X$, then $Y^{\prime}$ is also equal to the emptyset.
As for splits, a bi-join $\{X,Y\}$ is considered trivial if $\left|X\right|=1$ or $\left|Y\right|=1$, and a graph is considered prime with respect to bi-join if it has only trivial bi-joins. Let us denote by $\mathcal{B}_{\mathsf{join}}^{G}$ the set of bi-joins of a graph $G$. A proof of the bipartitiveness of bi-joins can be found in [Rao06].
{thmC}
[[Rao06]] For every undirected graph $G$, the system $\mathcal{B}_{\mathsf{join}}^{G}$ is bipartitive.
A corollary of Section 5.3 and of 23 is the existence of a $\operatorname{C}_{2}{\!\!}\operatorname{MSO}$-transduction taking as input an undirected graph $G$ represented as the $\{\mathsf{edge}\}$-structure $\mathbb{G}$ and outputting the bipartitive tree representing the set of bi-joins of $G$. We now propose a $\operatorname{C}_{2}{\!\!}\operatorname{MSO}$-transduction that takes as input the structure $\mathbb{G}\sqcup\mathbb{T}$ where $\mathbb{T}$ is the structure representing the bipartitive tree of $G$ and outputs a structure that is similar to the canonical split decomposition, from which one can reconstruct the original graph in $\operatorname{MSO}$. A corollary of this $\operatorname{C}_{2}{\!\!}\operatorname{MSO}$-transduction is that, for fixed positive $k$, we reduce the existence of a $\operatorname{C}_{2}{\!\!}\operatorname{MSO}$-transduction for computing rank-decompositions of cut-rank at most $f(k)$, for some function $f$, to the existence of a $\operatorname{C}_{2}{\!\!}\operatorname{MSO}$-transduction for computing rank-decompositions of cut-rank at most $f(k)$ on prime graphs with respect to bi-join.
Transducing the Skeleton graph [dM03]
Let $G$ be an undirected graph. If $X$ is a subset of $V(G)$, let $\equiv_{X}$ be the binary relation on $X\times X$ where $x\equiv_{X}y$ if $N(x)\cap(V(G)\setminus X)=N(y)\cap(V(G)\setminus X)$. One easily checks that $\equiv_{X}$ is an equivalence relation. Now if $\{X,Y\}$ is a bi-join, then $\equiv_{X}$ and $\equiv_{Y}$ both have at most two equivalence classes. We remind that if $T$ is a tree and $t$ is a node of $T$ and $s$ a neighbour of $t$, then $T_{s}^{t}$ is the connected component of TtT-t containing $s$, and $L(T)$ is the set of leaves of $T$.
Let $T$ be the bipartitive tree of the set $\mathcal{B}_{\mathsf{join}}^{G}$ of the bi-joins of $G$. By construction of $T$, for every node $u$ of $T$ and every neighbour $v$ of $u$, $\{L(T_{v}^{u}),L(T_{u}^{v})\}$ is a bi-join, and then $\equiv_{L(T_{v}^{u})}$ and $\equiv_{L(T_{u}^{v})}$ have both at most two equivalence classes. We denote by $\vec{uv}_{1}$ and $\vec{uv}_{2}$ the two equivalence classes of $\equiv_{L(T_{v}^{u})}$ in what follows (we omit $\vec{uv}_{2}$ if there is only one). Observe that if $v$ is a leaf, then $\equiv_{L(T_{v}^{u})}$, with $u$ its unique neighbour, has a single equivalence class reduced to $\{v\}$ and that we define as $\vec{uv}_{1}$. For every internal node $u$ of $T$, let’s denote by $G_{u}$ the graph whose vertex set is $\{\vec{uv}_{1},\vec{uv}_{2}\mid v$ is a neighbour of u}u\}, and there is an edge $z_{i}z_{j}$ in $G_{u}$ if there is $x\in z_{i}$ and $y\in z_{j}$ such that $xy\in E(G)$. Notice that there are $x\in z_{i}$ and $y$ in $z_{j}$ such that $xy\in E(G)$ if and only if every vertex in $z_{i}$ is adjacent to every vertex in $z_{j}$, and thus the graph $G_{u}$, for every node $u$, is unique up to isomorphism. The Skeleton graph associated with $T$ is a triple $(H,F_{1},F_{2})$ where
  • $H$ is a graph with vertex set $\bigcup\limits_{u\in V(T)\setminus L(T)}V(G_{u})$ and
  • edge set $\bigcup\limits_{u\in V(T)\setminus L(T)}E(G_{u})$,
  • $F_{1}$ is the set of edges $\{\vec{uv}_{i},\vec{vu}_{i}\}$, for $i\in[2]$ for which $uv\in E(G)$.
  • $F_{2}$ is the set of pairs $\{\vec{uv}_{1},\vec{uv}_{2}\}$ for every edge $uv$ of $E(T)$.
We call the edges of $H$ $\mathsf{c}$-edges, edges in $F_{1}$ $\mathsf{t}$-edges and edges in $F_{2}$ $\mathsf{r}$-edges and we naturally model the skeleton graph as $\{\mathsf{t}\text{-}\mathsf{edge},\mathsf{c}\text{-}\mathsf{edge},\mathsf{r}\text{-}\mathsf{edge}\}$-structures.
It is worth mentioning that the Skeleton graph is uniquely defined from $T$, and since $T$ is unique, up to isomorphism, one can conclude that the Skeleton graph is unique, up to isomorphism. The Skeleton graph defined in [dM03] does not include $\mathsf{r}$-edges, but we need such edges to be able to recover the graphs $G_{u}$ from the Skeleton graph because $G_{u}$ without the $\mathsf{r}$-edges is not necessarily connected. See Figure 9 for an example.
It is routine to show the following.
{lemC}
Let $G$ be a connected graph and let $(H,F_{1},F_{2})$ be its Skeleton graph. Then the following hold:
  1. (1)
    two vertices $x$ and $y$ of $G$ are adjacent if and only if there is a path between $x$ and $y$ in $(H,F_{1},F_{2})$ on which $\mathsf{c}$-edges and $\mathsf{t}$-edges alternate;
  2. (2)
    the vertices of $G$ are precisely the vertices of $(H,F_{1},F_{2})$ that are not incident to any $\mathsf{t}$-edges.
Figure 9. A graph $G$ with three non-trivial strong bi-joins (left), the partitive tree of $\mathcal{B}_{\mathsf{join}}^{G}$ (middle) and its Skeleton graph (right). Dashed edges are $\mathsf{t}$-edges, squiggly edges are $\mathsf{r}$-edges and the rest are $\mathsf{c}$-edges. Equivalence class vertices are represented by squares apart from the singleton equivalence classes corresponding to vertices of $G$. Note that the edges of the graph $G$ correspond one-to-one to paths in the skeleton graph that alternate between $\mathsf{t}$-edges and $\mathsf{c}$-edges. The graph depicted was taken from [Rao06, Figure 4.5].
It is routine to write an $\operatorname{MSO}$-formula checking that, between two vertices, there is a path that alternate $\mathsf{c}$-edges and $\mathsf{t}$-edges. Hence, we can conclude that there is an $\operatorname{MSO}$-transduction which on input of the $\{\mathsf{t}\text{-}\mathsf{edge},\mathsf{c}\text{-}\mathsf{edge},\mathsf{r}\text{-}\mathsf{edge}\}$-structures $\mathcal{H}$ modeling the skeleton graph of $G$ outputs the $\{\mathsf{edge}\}$-structure $\mathbb{G}$ modeling $G$.
We are now ready to construct the $\operatorname{C}_{2}{\!\!}\operatorname{MSO}$-transduction that computes the Skeleton graph of a graph.
Theorem 26.
There exists a non-deterministic $\operatorname{C}_{2}{\!\!}\operatorname{MSO}$-transduction $\tau$ such that for any connected graph $G$ represented as the $\{\mathsf{edge}\}$-structure $\mathbb{G}$, $\tau(\mathbb{G})$ is non-empty and every output in $\tau(\mathbb{G})$ is equal to some $\{\mathsf{t}\text{-}\mathsf{edge},\mathsf{c}\text{-}\mathsf{edge},\mathsf{r}\text{-}\mathsf{edge}\}$-structure $\mathbb{H}$ representing the Skeleton graph of $G$.
Proof 5.5.
Let $G$ be a connected graph represented by the $\{\mathsf{edge}\}$-structure $\mathbb{G}$, $(V(G),\mathcal{B}_{\mathsf{join}}^{G})$ the bipartitive set system consisting of all bi-joins of $G$ represented by the $\{\mathsf{BIPART}\}$-structure $\mathbb{B}$, $(H,F_{1},F_{2})$ its Skeleton graph represented by the $\{\mathsf{t}\text{-}\mathsf{edge},\mathsf{c}\text{-}\mathsf{edge}\}$-structure $\mathbb{H}$ and $T$ represented by the $\{\mathsf{t}\text{-}\mathsf{edge},\mathsf{degenerate}\}$-structure, the bipartitive tree of $(V(G),\mathcal{B}_{\mathsf{join}}^{G})$.
Our first step is to transduce $\mathbb{B}$ from $\mathbb{G}$. For this it is sufficient to observe that we can easily define an $\operatorname{MSO}$-formula $\phi_{\mathsf{BIPART}}(X)$ with one monadic free-variable $X$ which defines bi-joins $\{X,U\setminus X\}$ of $G$.
Since $\mathbb{B}$ is bipartitive, we can apply 21 to transduce $\mathbb{G}\sqcup\mathbb{B}\sqcup\mathbb{T}$ from $\mathbb{G}\sqcup\mathbb{B}$, where $\mathbb{T}$ is the structure representing the bipartitive tree of $G$. In the following we describe how to transduce $\mathbb{H}$ from $\mathbb{G}\sqcup\mathbb{B}\sqcup\mathbb{T}$.
  • The first transduction chooses non-deterministically a node $r$ of $T$ and root $T$ at $r$ and defines the laminar set system $\mathbb{S}$ where $\mathsf{SET}(X)$ holds if there is a node $u$ with $X$ the set of leaves of the subtree of $(T,r)$ rooted at $u$. Let’s denote by $L_{u}$ the set of leaves of the subtree rooted at $u$.
  • By 12 and 13 there are four bi-colourings $(A_{i},B_{i})_{i\in[4]}$ identifying $V(T)$ and for each $i\in[4]$, there are $\operatorname{C}_{2}{\!\!}\operatorname{MSO}$-formulas ${\mathsf{repr}}_{B_{i}}(a,X)$ and ${\mathsf{repr}}_{A_{i}}(b,X)$ that are satisfied exactly when $X$ is, respectively, $A_{i}$-represented by $a$ and $B_{i}$-represented by $b$. The second transduction guesses non-deterministically the four bi-colouring $(A_{i},B_{i})_{i\in[4]}$ that identifies nodes of $(T,r)$ as a laminar tree of the laminar set system $\mathbb{S}$. Notice we can filter by accepting only the guesses with unique request and such that, for each node, there is a exactly one $i\in[4]$ such that it is uniquely $A_{i}$-represented and uniquely $B_{i}$-represented.
  • The third transduction makes four copies of each internal node $u$, adds $\mathsf{r}$-edges between copies $1$ and $2$, and between copies $3$ and $4$.
  • It remains now to create the $G_{u}$ graphs, for each internal node $u$. Assume $v$ is the parent of $u$, and let $a_{v}$ be the leaf such that (${\mathsf{repr}}_{A_{i}}(a_{v},L_{v})$ or ${\mathsf{repr}}_{B_{i}}(a_{v},L_{v})$) and $a_{v}\notin L_{u}$, and let $a_{u}$ be a leaf such that ${\mathsf{repr}}_{A_{j}}(a_{u},L_{u})$. Notice that $a_{v}$ exists because $v$ is the least common ancestor of the two leaves that $A_{i}$ and $B_{i}$-represent $v$, and if both do not belong to $L_{u}$, we deterministically define $a_{v}$ as the one that $A_{i}$-represents $v$. Copy $1$ of $u$ will play the role of the equivalence class of $\equiv_{T_{v}^{u}}$ containing $a_{v}$. The copy $2$ will play the role of the equivalence class of $\equiv_{T_{v}^{u}}$ not containing $a_{v}$. The copy $3$ will play the role of the equivalence class of $\equiv_{L_{u}}$ containg $a_{u}$, and the copy $4$ will play the role of the equivalence class of $\equiv_{L_{u}}$ not containing $a_{u}$. The vertex set of $G_{u}$ will be copies $1$ and $2$ of $u$, and the copies $3$ and $4$ of each child of $u$. Notice that there is an edge between two equivalence classes $z$ and $z^{\prime}$ when there are vertices $x\in z$ and $y\in z^{\prime}$ such that $xy$ is an edge of $\mathbb{G}$. Since we have access to a vertex representing an equivalence class in $L_{w}$, for $w$ a child of $u$, and a vertex representing an equivalence class in $L(T_{v}^{u})$, one can write a $\operatorname{C}_{2}{\!\!}\operatorname{MSO}$-formula checking that the other equivalence class of $L_{w}$, $w$ a child of $u$, (resp. of $L(T_{v}^{u})$) is non-empty, and then whether there are edges between two equivalence classes. We can therefore write a $\operatorname{C}_{2}{\!\!}\operatorname{MSO}$-formula stating that two copies of vertices are related by a $\mathsf{c}$-edge.
  • The fourth transduction adds the $\mathsf{t}$-edges. We add a $\mathsf{t}$-edge between a copy $i\in\{1,2\}$ of $u$ and a copy $j\in\{3,4\}$ of $u$, if a vertex of the equivalence class the $i$th copy represents is adjacent to a vertex of the equivalence class the $j$th copy represents. Such edges can be $\operatorname{C}_{2}{\!\!}\operatorname{MSO}$-defined in the same way the $\mathsf{c}$-edges are defined, except they relate only particular copies of a same node.
The composition of the four transductions above computes the Skeleton graph of $G$. Since the non-determinism comes only from the construction of $\mathbb{T}$ and the guessing of the bi-colouring, and we can filter after each such step, we guarantee that an output always satisfies the desired outcome.
6. Conclusion
We provide transductions for obtaining tree-like graph decompositions such as modular decompositions, cotrees, split decompositions and bi-join decompositions from a graph using $\operatorname{C}{\!\!}\operatorname{MSO}$. This improves upon results of Courcelle [Cou96, Cou06] who gave such transductions for ordered graphs. It is worth mentioning that 22 can be also used to $\operatorname{C}{\!\!}\operatorname{MSO}$-transduce canonical decompositions of other structures such as Tutte’s decomposition of matroids or generally split-decompositions of submodular functions [Cun82] or modular decompositions of $2$-structures [EHR99] or of hypergraphs [HdMMZ22]. As shown by the application given in [Cou06] for transducing Whitney’s isomorphism class of a graph, a line of research is to further investigate which structures can be $\operatorname{C}{\!\!}\operatorname{MSO}$-transduced from a graph or a set system by using the transductions from 2. Also, naturally, the question arises whether counting is necessary or whether $\operatorname{MSO}$ is sufficient to transduce such decompositions.
Furthermore, it is known that if a family of set systems $\mathfrak{S}$ is obtained from a family $\mathfrak{F}$ of $\Sigma$-labeled trees, for some finite alphabet $\Sigma$, by a $\operatorname{C}{\!\!}\operatorname{MSO}$-transduction $\tau$, then for any $\operatorname{C}{\!\!}\operatorname{MSO}$ formula $\varphi$, the set $\{T\in\mathfrak{\mid}\tau(\mathbb{T})\models\varphi\}$ is a recognisable subfamily of $\mathfrak{F}$ [CE12, FMN22]. However, the converse, i.e., whenever a subfamily $\mathfrak{F^{\prime}}$ of $\mathfrak{F}$ is recognisable, then the set $\{\tau(T)\mid T\in\mathfrak{F^{\prime}}\}$ is $\operatorname{C}{\!\!}\operatorname{MSO}$-definable is widely open, and was originally conjectured in [Cou90] for graph classes of small tree-width. Even if Courcelle’s conjecture is now settled [BP16], the case of graphs that are images of $\operatorname{C}{\!\!}\operatorname{MSO}$-transductions, equivalently graph classes of bounded clique-width [CE12], is still open. The $\operatorname{C}{\!\!}\operatorname{MSO}$-definability results shown in this paper, combined with some known special cases, imply that we can push further the known graph classes where recognizability equals $\operatorname{C}{\!\!}\operatorname{MSO}$-definability. In the following subsection we introduce the necessary notation before providing the corollary.
6.1. Recognizability equals definability in restricted classes of graphs of small rank-width
If $G$ is a graph, we denote by $A_{G}$, the adjacency matrix of $G$, a matrix over the binary field whose rows and columns are indexed by $V(G)$ and such that $A_{G}[x,y]=1$ only when $xy$ is an edge. For a graph $G$, we denote by $\rho_{G}:2^{V(G)}\to\mathbb{N}$, the cut-rank function of $G$ where $\rho_{G}(X)$, for $X\subseteq V(G)$, is the rank, over the binary field, of the submatrix of $A_{G}$ whose rows are indexed by $X$ and its columns by $V(G)\setminus X$.
A rank-decomposition of a graph $G$ is a pair $(T,\delta)$ where $T$ is a tree and $\delta:V(G)\to L(T)$ is a bijection from the vertex set of $G$ to the leaves of $T$. The cut-rank of $(T,\delta)$ over $G$ is defined as
$\max\limits_{u\in V(T),X\subseteq N_{T}(u)}\left\{\rho_{G}\left(\bigcup_{v\in X}\delta^{-1}(L(T_{v}^{u}))\right)\right\}.$
The rank-width of a graph $G$ is the maximum cut-rank over all rank-decompositions of $G$. It is known that a graph class $\mathcal{C}$ is included in the image of a $\operatorname{C}{\!\!}\operatorname{MSO}$-transduction taking as inputs labeled trees if and only if the rank-width of graphs in $\mathcal{C}$ is bounded by a constant, see for instance [CE12, CGK+25b]. It is also proved in [CGK+25b] that, for every $k$, there are two $\operatorname{C}{\!\!}\operatorname{MSO}$-transductions $\tau_{k}$ and $\mathsf{val}_{k}$ such that $\tau_{k}$ takes as input $\mathbb{G}\sqcup\mathbb{T}$, where $\mathbb{T}$ is a structure describing a rank-decomposition of $G$ of cut-rank at most $k$ and whose leaves are copies of vertices of $G$, and outputs a finitely labeled tree $t$ with $G$ included in $\mathsf{val}_{k}(t)$. Combining this with the following implies that $\operatorname{C}{\!\!}\operatorname{MSO}$-definibility equals recognizability for the following graph classes : distance-hereditary graphs and more generally graphs whose prime graphs are bounded by a constant or have small linear clique-width or have small tree-width or belong to some families of $H$-free graphs such as the once studied in [BDLM05, Rao07].
Theorem 27.
Let $\mathcal{C}$ be a class of graphs for which there is, for every $k$, a $\operatorname{C}{\!\!}\operatorname{MSO}$-transduction $\varphi_{\mathcal{C},k}$, that takes as input a graph $G$ in $\mathcal{C}$ of rank-width at most $k$ and outputs a rank-decomposition of $G$ of width at most $f(k)$, for some computable function $f$. Then, for every positive integer $k$ and for every class of graphs $\mathcal{D}$ whose prime graphs with respect to split decomposition (resp. bi-join decomposition) belong to $\mathcal{C}$, there is a $\operatorname{C}{\!\!}\operatorname{MSO}$-transduction $\psi_{\mathcal{D},k}$ that takes as input a graph $G$ in $\mathcal{D}$ of rank-width at most $k$ and outputs a rank-decomposition of $G$ of width at most f(k)f(k)(resp. $f(k)+2$).
Let’s recall the Parallel Application Lemma before proceeding to its proof.
[…]
  • the universe of $\bigsqcup_{0<i\leq n}\mathbb{A}_{i}$ is the union of the universes of the $\mathbb{A}_{i}$’s (which are required to be disjoint);
  • for each relation or predicate name from $\Sigma$, its interpretation in $\bigsqcup_{0<i\leq n}\mathbb{A}_{i}$ is the union of its interpretations in each of the $\mathbb{A}_{i}$’s;
[…]
{lemC}
[(Parallel Application Lemma [BGP21])] Let $\tau$ be a $\Sigma$-to-$\Gamma$ $\operatorname{C}{\!\!}\operatorname{MSO}$-transduction. Then there is a $(\Sigma\cup\{\thicksim\})$-to-$(\Gamma\cup\{\thicksim\})$ $\operatorname{C}{\!\!}\operatorname{MSO}$-transduction $\hat{\tau}$ such that, for every sequence $\mathbb{I}_{1},\ldots,\mathbb{I}_{n}$ of $\Sigma$-structures and every sequence $\mathbb{O}_{1},\ldots,\mathbb{O}_{n}$ of $\Gamma$-structures, we have $(\bigsqcup_{0<i\leq n}\mathbb{I}_{i},\bigsqcup_{0<i\leq n}\mathbb{O}_{i})\in\hat{\tau}$ if and only if there exists a permutation $\pi$ of $[n]$ such that $(\mathbb{I}_{i},\mathbb{O}_{\pi(i)})\in\tau$ for all $i\in[n]$.
Proof 6.1 (Proof of 27).
Let $G$ be a connected graph represented by the $\{\mathsf{edge}\}$-structure $\mathbb{G}$. Assume first that we are interested in prime graphs with respect to split decomposition. The transduction is the composition of the following.
  1. (1)
    The first transduction takes as input $\mathbb{G}$ and outputs, thanks to 25, a $\{\mathsf{c}\text{-}\mathsf{edge},\mathsf{t}\text{-}\mathsf{edge}\}$-structure $\mathbb{H}$ which is the canonical split decomposition of $G$.
  2. (2)
[…]
  3. (3)
[…]
  4. (4)
[…]
  5. (5)
    The last transduction takes all rank-decompositions outputted by $\hat{\varphi^{\prime}}$ and fuses any two leaves that are related by a $\mathsf{t}$-edge in $\mathbb{H}$.
It is routine to check that the composition, in the same order, of the above transductions output a rank-decomposition of $G$ as the $\mathsf{t}$-edges form a tree and because each of them yields a correct output when given a correct input. Because the rank-decompositions outputted by $\hat{\varphi^{\prime}}$ have cut-rank at most $f(k)$ on prime subgraphs and of cut-rank $1$ otherwise, it is easy to check that the cut-rank of the outputted rank-decomposition of $G$ has cut-rank at most $f(k)$. If $G$ is not a connected graph, then we apply the above $\operatorname{C}{\!\!}\operatorname{MSO}$-transduction on each connected component thanks again to the Parallel Application Lemma, and then a second non-deterministic transduction adds a new node whose neighbour in the rank-decomposition of each connected component is an arbitrary internal node. The proof for bi-join decompositions works very similar.